### v2

# Sketch-Plan-Generalize: Continual Few-Shot Learning of Inductively Generalizable Spatial Concepts

Abstract
Our goal is to enable embodied agents to learn inductively generalize-able spatial concepts, e.g., learning staircase as an inductive composition of towers of increasing height. Given a human demonstration, we seek a learning architecture that infers a succinct program representation that explains the observed instance. Additionally, the approach should generalize inductively to novel structures of different size or complex structures expressed as a hierarchical composition of previously learned concepts. Existing approaches that use code generation capabilities of pre-trained large (visual) language models as well as purely neural models show poor generalization to a-priori unseen complex concepts. Our key insight is to factor inductive concept learning as: (i) Sketch: detecting and inferring a coarse signature of a new concept (ii) Plan: performing MCTS search over grounded action sequences (iii) Generalize: abstracting out grounded plans as inductive programs. Our pipeline facilitates generalization and modular re-use enabling continual concept learning. Our approach combines the benefits of code generation ability of large language models (LLM) along with grounded neural representations, resulting in neuro-symbolic programs that show stronger inductive generalization on the task of constructing complex structures in relation to LLM-only and neural only approaches. Furthermore, we demonstrate reasoning and planning capabilities with learned concepts for embodied instruction following.
1 Introduction
The ability to learn novel concepts which generalize inductively is one of the hallmarks of human intelligence [mind]. Humans are highly data efficient – observing at a few instances of towers of a certain heights, we can generalize to constructing towers of any height. Further, we interpret increasingly complex concepts as hierarchical composition over simpler ones, e.g., a tower as a sequence of blocks placed on top of each other, or a staircase composed of towers of increasing height.
This paper considers the problem of learning a program representation, from few demonstration, that models the inductive realization of grounded spatial concepts. Learning of such concepts is challenging task due to an expansive space of programs and the need to reason about their physical plausibility. Further, the representation must support inductive generalization over learned concepts as well as express complex hierarchical concepts via modular re-use of concepts learnt previously.
Figure 1: Learning inductive spatial concepts. (Left) A human demonstrates the construction of a row and tower of size three. (Right) The agent learns program representation that enables inductive generalization to novel structures (varied sizes and visual attributes) and expresses complex concepts as hierarchical composition of previously acquired ones. E.g., learning a tower as a sequence of blocks placed one on top of another and a pyramid as rows of decreasing size.
Prior efforts such as [code-as-policies] uses a LLM (Large Language Model) to generate control program for a given task specification but fail to generalize to complex spatial concepts which are difficult to tokenize. Extension of this to VLMs (Vision Language Models), [achiam2023gpt], also fail to generalize when presented with linguistically novel concepts, due to over reliance on prior knowledge and their inability to effectively learn novel concepts from given demonstrations. On the other hand, neural approaches such as  [struct-diffusion], learn from given demonstrations but generalize poorly due to their inability to (a) explicitly model symbolic concept of induction and (b) modularize as well as re-use previously acquired concepts. Approaches such as [li2019practical] train RL-based policies to attain spatial assembly by encoding an inductive spatial prior using GNNs within the policy architecture. However, generalization is still limited and assumes an elicitation of the goal as per object positions, resulting in lack of ability to take in goal description such as “construct a tower of size three”. In summary, we attribute poor generalization of such approaches to an implicit entangling of the following objectives: (i) postulating a high-level program for new concepts, (ii) evaluating plausibility of grounded plans to align with human demonstration of concepts to be learned and (iii) abstracting out a grounded program to facilitate inductive generalization and modular re-use in a continual manner.
In response, this paper introduces an approach that factorizes the concept learning task as: (a) Sketch: Given a language-annotated demonstration of a novel concept, an LLM is used to postulate a function signature. (b) Plan: Refinement of program sketches via MCTS search, rapidly evaluating actions sequences guided by a reward associated with constructing a concept. The search is accelerated by training a neural action predictor that uses the given demonstrations. (c) Generalize: Leveraging the code generation capability of an LLM to distill grounded plans into a program that is inductively generalize-able. This results in a continually evolving library of concepts which can be used to hierarchically learn complex concepts in future. The modular architecture enables continual learning by providing the ability to decide whether the new concept encountered either as a symbolic composition of existing concepts, or, a neural embedding trained via gradient update.
Our experimental evaluation demonstrates accurate learning of simple and complex concepts from few demonstrations for a range of spatial structures. Further, our approach also shows inductive generalization in out-of-distribution settings, significantly over the baselines. We also present deeper insights related to efficiency gains obtained by combining symbolic MCTS with neural action predictor. Furthermore, we show how learned concepts can be grounded in the visual input, enabling a robot to follow natural language instructions referring to a-priori unseen spatial configurations.
2 Related Works
The ability to ground symbolic knowledge is central to building intelligent embodied agents. The symbol grounding problem is often cast as learning an association between linguistic description of a concept and aspects of the state-action space such as objects [paul2018efficient, ramp], spatial relations  [tellex2011approaching, kim2024lingo], reward functions  [boularias2015grounding], or motion constraints  [howard2014natural]. These works assume the presence of grounded representation for symbolic concepts and only learn associations between language and concepts. In contrast our work jointly learns higher-order concepts composed of simpler concepts along with their grounding in the robot’s state and action space.
Another line of work  [singh2022progprompt, wang2023demo2code, ahn2022can, code-as-policies] leverages prior-knowledge embodied in LLMs to determine grounded robot control programs for a high-level tasks, given a tokenized description of the scene and robot capabilities. Our experiments (reported subsequently) demonstrate that complex spatial relations are challenging for textual descriptions to encode and are difficult to reason over. Hence, we adopt a factored approach that uses a search procedure guided by visual reward, acting as a medium to connect the task sketch generation and program abstraction stages (performed by an LLM).
Works such as [wang2023voyager, jarvis-1] shows lifelong learning of skills by learning to plan high-level tasks through composition of simple skills for simulated agents. Efforts such as  [lotus],  [lifelong-halp] detect when control primitives for a robot are insufficient to perform a task and initiate skill learning, building a library over time. However, such efforts do not model deep inductive use of learned concepts.
Finally, this work draws inspiration from concept learning approaches in the broader AI community. Related works infer latent generative programs to express hand-written digits [lake2015human], object arrangements [ellis2018learning], motion plans [nscl] or goal-directed policies [lpp]. These approaches do not explicitly model inductive concept of modular re-use of previously acquired concepts. Other efforts such as [li2019practical] learn RL-based policies to construct object stacks. Here, the neural policy models locality of object influence via GNNs. However, the model suffers from poor generalization to unseen examples, (eg. tower of larger size) and do not possess a mechanism to re-use previously acquired concepts.
3 Background and Problem Formulation
We consider an embodied agent that uses a visual and depth sensor to observe its environment and can grasp and release objects at specified poses. We represent the robot’s domain as a goal-conditioned MDP $<\mathcal{S},\mathcal{A},\mathcal{T},g,\mathcal{R},\gamma>$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ is the transition function, $g$ is the goal, $R$ is the reward model and $\gamma$ is the discount factor. The agent’s objective is to learn a policy that generates a sequence of actions from an initial state $s_{0}$ to achieve the goal $g$ in response to an instruction $\Lambda$ specifying the intended goal from a human. We assume that the agent possesses a model of semantic relations (e.g., left(), right() etc.) as well as semantic actions such as moving an object by grasping and releasing at a target location. Such modular and composable notions can be acquired from demonstrations via approaches outlined in [nsrm-Kalithasan2022LearningNP, nscl, mao2022pdsketch]. Such notions populate a library of concepts $\mathcal{L}$ available as grounded executable function calls. Following recent efforts [code-as-policies, Huang2022InnerME, ahn2022can, singh2022progprompt] in representing robot control directly as executable programs, we represent action sequence corresponding to a plan as a program consisting of function calls to executable actions and grounded spatial reasoning.
Our goal is to enable a robot to interpret instructions such as “construct a tower with red blocks of height five". Specifically, we aim to learn spatial constructs like a tower that requires sequential actions that repeatedly places a block on top of a previously constructed assembly; a process akin to induction. Given a few demonstrations of a constructing a spatial assembly, $\mathcal{D}$, each consisting of natural language description $\Lambda$ (“construct a tower of red blocks of size five") and a sequence of key frame states $\{S_{1},\cdots,S_{g}\}$ associated with the construction process, we seek to learn a program that models the inductive nature of the concept of tower. Online, the agent must generalize inductively to following an instruction as “construct a tower of blue blocks of size ten". Once the concept of tower is acquired, it is added to the concept library. This aids in learning the notion of a staircase as sequence of towers of increasing height.
We formalize the notion of inductive spatial concepts and formulate the learning objective. Let $C_{1},\cdots,C_{|\mathcal{L}|}$ represent the topological order of concepts present in the concept library $\mathcal{L}$ as per their structural complexity. An inductive spatial concept $C_{k}$ of size $n$ at position $p$ can be expressed as:
$h(C_{k},n,p)=\underbrace{h^{\lambda}(C_{k},n-1,pos(.))}_{\text{Induction (I)}}\circ\underbrace{\prod\limits_{l=1}^{L(C_{k})}h\big{(}C_{k^{\prime}_{l}},\mathtt{size}(.),pos(.)\big{)}}_{\text{Composition (C)}}\circ\underbrace{\prod\limits_{l=1}^{L^{\prime}(C_{k})}\eta_{\theta}^{l}\big{(}pos(.)\big{)}}_{\text{Base (B)}},$ (1)
where, $\lambda\in\{0,1\}$, $k^{\prime}<k$, $0\leq L(C_{k}),L^{\prime}(C_{k})\leq o(|\mathcal{L}|)$ and $\eta_{\theta}^{l}$ is the set of primitive actions resulting in placement of object at a desired spatial position. Here, $pos(.)=pos(C_{k},l,n,p)$ and $\mathtt{size}(.)=\mathtt{size}(C_{k},l,n)$ are functions that predict the size and position of the structure to be constructed. For example, constructing a tower of size $n$ at position $p$ i.e., $h(Tower,n,p)=h^{\lambda=1}(Tower,n-1,p)\,\circ\,OnTop_{\theta}(p+(n-1)d)$. Here, $\lambda=1$, $L(C_{k})=0$, $L^{\prime}(C_{k})=1$, $\eta^{1}=OnTop$, $pos(Tower,l=1,n,p)=p+(n-1)\cdot d$ and $d$ is the dimension of the object. Thus, a tower can be described using induction and a single primitive base function with zero compositions. On the other hand, enclosures can be described as a composition of two rows and two columns without induction or base components. The functional space of inductive concepts ($h$) leads to a hypothesis space $\mathcal{H}$ of associated neuro-symbolic programs. Each goal-reaching demonstration corresponds to a particular instantiation of a given inductive concept, i.e. $H(C_{k},n,p)$, where the $p$ comes from the sequence of frames, and $n$, $C_{k}$ comes from $\Lambda$.
Given demonstration, we aim to learn a generic representation for the given concept, which is general for all n, p i.e. $h(C_{k},\cdot,\cdot)$. We model this as learning a particular $H=h(C_{k},\cdot,\cdot)$ where $H\in\mathcal{H}$. Given (few) demonstrations of a human constructing a spatial structure, concept learning can be formulated as the Bayesian posterior [lake2015human, shah2018bayesian, lpp], $\mathtt{P}_{\mathcal{H}}(H|\Lambda,S_{1}..S_{g})\propto\mathtt{P}(S_{1}..S_{g}|\Lambda,H)\cdot\mathtt{P}(H|\Lambda)$. Here, the likelihood term associates a candidate program, and the prior term regularizes the program space. The maximum a-posteriori estimate, representing the learnt program, is obtained by optimizing the following objective:
$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\left(\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))\right)-\log\left(\mathtt{P}(H|\Lambda)\right)\right]$ (2)
Since exact inference is intractable, approximate inference is performed via search in the program space. Note that learning inductive spatial concepts given demonstration considers programs that represent plans that attain physically grounded/feasible structures, an object we model during the search. Additionally, we seek strong generalization from a few instances of an inductive structure to structures with arbitrary sizes, in effect favouring programs with iterative looping constructs.
4 Technical Approach: Learning Inductive Spatial Concepts
Figure 2: Approach Overview. We learn a neuro-symbolic program for inductive spatial concepts factored as (a) Sketch (b) Plan (c) Generalize. The example above shows the progressive realization of a program for the concept of a tower acquired by observing a single demonstration of building a tower of size three and its corresponding natural language instruction, “construct a tower of height three using green die."
We address the problem of estimating a succinct generalized program, Eq.  2, modeling an inductive spatial concept modeling structures whose construction is observed in a human demonstration. Direct symbolic search in the space of programs is intractable (particularly due to looping constructs needed for modeling induction) but can explicitly reason over previously acquired concepts. Alternatively, neural methods attempting to predict action sequence to attain the assembly are challenged by continual setting where concepts can increase over time building on previously learnt ones but are resilient to noise. Our approach blends both approaches and factors the concept learning task as:
  • Sketch: From the natural language instruction ($\Lambda$), we extract a task sketch ($H_{S}^{*}$) using an LLM that provides the signature (concept name and instantiated arguments) of the concept to be learned. When grounded in the initial scene of the demonstration, the task sketch provides a particular instance of the concept demonstrated in the given demonstration.
  • Plan: MCTS-based search using the already learnt concepts that outputs the sequence of grounded actions, best explaining the given demonstration.
  • Generalize: The grounded plan $H_{P}^{*}$ and task sketch $H_{S}^{*}$ are provided to an LLM to obtain a general Python program whose execution on the given scene matches the searched plan.
The progress exploration of the program space can be summarised as follows:
$H_{S}^{*}\leftarrow Sketch(\Lambda\,;\,{\theta_{S}})\,;\,H_{P}^{*}\leftarrow Plan(S_{1}..S_{g},H_{S}^{*};\,{\theta_{P}})\,;\,H_{G}^{*}\leftarrow Generalize(H_{P}^{*},\,H_{S}^{*};\,{\theta_{G}})$ (3)
The concept library $\mathcal{L}$ is initialized with primitive visual and action concepts. Upon acquiring a new inductive concept $H_{G}^{*}=H^{*}$, we update our library accordingly: $\mathcal{L}\leftarrow\mathcal{L}\cup{H^{*}}$. Next, we detail each of the three steps mentioned above. Fig. 2 illustrates an example of progressive program estimation.
4.1 (Sketch) Grounded Task Sketch Generation
An LLM driven by in-context learning is used to get a program signature (a sketch) for a concept from the natural language instruction. The task sketch is a tree of nested function calls that outlines the function header (name and the parameters) of the inductive concept/program to be learned. A detailed exposition on prompting appears in the Appendix  C.1. The task sketch thus obtained is then grounded on the input scene using a quasi-symbolic visual grounding module akin to [nscl, nsrm-Kalithasan2022LearningNP, wang2023progport]. This module has three key components: (1) a visual extractor (ResNet-34 based) that extracts the features of all objects in the scene, (2) a concept embedding module that learns disentangled representations for visual concepts like green and dice, and (3) a quasi-symbolic executor equipped with pre-defined behaviours such as “filter” to select/ground the objects of interest. For example, grounding the task sketch “Tower (height =3, objects = filter(green, dice))” results in an instantiated function call “Tower(height = 3, objects = [1, 2, 3])” where [1, 2, 3] are the green coloured dice.
4.2 (Plan) Physical Reward Guided Plan Search
The plan search involves finding a generalizable plan that effectively explains the demonstration $S_{1},\cdots,S_{g}$. Specifically, this involves determining the concepts, their respective grounded parameters, and the order of composition as specified in the equation 1.
Identifying the position $pos_{\theta}(.)$ of the object placement/structure construction is achieved through a notion of head. Essentially, the head represents a cuboidal enclosure in 3D space. The movement of any object involves two key stages: (1) relocating the abstract head to the desired position and (2) picking the target object and placing it at the current position of the head. Conceptually, the movement of the head can be likened to the robot’s cognitive exploration of potential object placements to attain the desired spatial configuration. We define a set of four primitive functions, $\mathcal{A}_{p}$, that manipulate the head. (1) store_head(): remembers the current position of the head in a stack. (2) reset_head(): sets the head to the last remembered position (i.e., to the top of the stack) and pops this position from the stack (3) keep_at_head(objects): places the object corresponding to the index with highest value in objects (a list of values) at the current location of head and (4) move_head(direction): A neural operator that takes the head and predicts the new location of the head with respect to the given direction. The move_head function is trained using a corpus of pick and place instructions, such as “move the green object to the right of the red cube” similar to the approach in [nsrm-Kalithasan2022LearningNP]. When an object is passed as the parameter, the move_head operator is overloaded to assign the head to the location of the object (also referred to as assign_head(object)).
The state is defined by bounding boxes (including the depth of the center) and visual attributes of all the objects that are present on the table. We assume that the head is initially assigned to the position where the structure needs to be constructed. For each learned inductive concept <cpt>, we define a macro-action “Make_<cpt>(size)" that executes the corresponding program with the given size argument, resulting in the construction of the desired inductive concept. Thus, the action space $\mathcal{A}$ is the union of primitive actions $\mathcal{A}_{p}$ and compound/macro-actions $\mathcal{A}_{c}$. All actions except keep_at_head(objects) and macro actions yield a reward of zero, while for the latter, the Intersection over Union (IoU) between the attained state and the expected state in the demonstration is provided as a reward. An MCTS procedure is performed to find a plan that maximizes the reward. The search outputs a sequence of grounded actions for an instantiation of the given inductive concept by the task arguments.
MCTS that searches for a plan only in terms of primitive actions may not be generalizable due to lack of modularity  C.3. The use of macro-actions in the search ensures that the plan $H_{p}^{*}$ for a given demonstration is concise, modular, and easily generalizable. This can be seen as a form of regularization in terms of the length of concept description by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2) $H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{n}\},H(\Lambda))+\alpha\log|H|\right]$. However, as the action space expands with the learning of more concepts, the search becomes slower, necessitating the pruning of the search space. To avoid searching over the size parameter in macro-actions, we greedily select the smallest size that achieves the maximum average reward from the current state. Additionally, to prune primitive actions, we train a reactive policy $\pi_{neural}$ which, given the current state $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (from the demonstration), outputs one of the primitive actions $a_{t}^{*}\in\mathcal{A}_{p}$. Consequently, the effective branching factor of the search is reduced to $|\mathcal{A}_{c}|+1$. Thus, our MCTS algorithm is modular through the hierarchical composition of learned concepts and efficient through action space pruning, and is referred to as MCTS+L+P. Further details regarding modifications in MCTS are given in the Appendix  A.4.
4.3 (Generalize) Plan to Program Abstraction
Leveraging the code generation and pattern matching abilities of LLMs [pattern] we use GPT-4 to distil out a general Python program from the sequence of grounded actions as determined by MCTS+L+P. The learnt program is incorporated in the concept library, $\mathcal{L}$, for modular reuse in subsequent learning tasks. Additional details, prompting mechanism and use of learnt programs in the search step of future learning tasks are described in Appendix  C.2,  A.4. We take a curriculum learning approach beginning from learning of primitive actions and visual attributes, followed by structures of increasing complexity. Additional details about learning of inductive concepts via curriculum, architecture details, and learning from multiple demonstrations are in Appendix  C.2, Fig.  7,  A.5.
5 Evaluation Setup
Data sets. Evaluation data set was created with a simulated Franka Emika Robot Manipulator constructing arrangement of blocks of uniform size with varied type/appearance (e.g., cubes, dice, lego etc.). The scene is observed with a visual and depth sensor with object masks extracted from a standard segmentation procedure. A total of $15$ structure types are considered including towers, staircases, X (cross), arch-bridge. The state and action sequences for attaining these structures forms the demonstration data fro training, details in Appendix  10.
Figure 3: Two illustrative examples of complex spatial structures from our dataset with a hierarchical composition of inductively defined structures.
The training data for each concept consists of $3$ demonstration per structure, of size (num. of inductive steps) $3-5$ with up to $20$ objects. We form three evaluation data sets, each consisting of simple structures (different size instances of a concept) and complex structures composed of simpler concepts (e.g., staircase consists of towers as substructure). Dataset I includes contains structures ranging in sizes $3-5$. Dataset II reverses the linguistic labels (e.g., the “tower" becomes “rewot"). Dataset III includes concepts of larger sizes ($4-7$).
Alternative approaches. We compare our model with four baselines spanning two approaches:
(1) Purely-Neural: An end-to-end neural model inspired by StructDiffusion, [struct-diffusion] that treats structure construction as a rearrangement problem. We consider two variations of the model 1.1) Struct-Diff (SD): End-to-end approach without any additional supervision regarding which objects need to be moved. 1.2) With-Grounder (SD+G): Similar to 1.1) except that we assume a perfect object selector/grounder that identifies the relevant set of objects which are to be moved.
(2) Pre-trained models that directly output symbolic programs: 2.1) LLM for Scene-Graph Reasoning: pre-trained language models (GPT-4) that generate Python programs from instructions which would describe the given demonstration. To help LLM understand the underlying structure, symbolic spatial relationships (e.g., left(a,b)) between objects in the demonstration are also provided to the LLM. For this baseline we further assume absence of distractor objects in the scene. 2.2) Vision Language Model (GPT-4-V): Similar to 2.1 but has the ability to take input demonstration as images. For learning program of a new inductive concept we give the demonstration to the VLM in the form of $\Lambda,(S_{1}..S_{g})$. Additional details on prompting method in Appendix C.6111Experiments were also conducted with open source code-generation LLM such as CodeLlama (70Bq), Due to significantly poorer performance w.r.t. GPT-4, it was removed as the primary LLM baseline..
Model variants. We implement three variants of the MCTS search to perform a grounded plan search over the action space $\mathcal{A}$: (i) MCTS+P+L: Our approach as described in section  4.2, that uses the learnt concepts from $\mathcal{L}$ as macro actions (e.g., Make_Tower(3, objects) $\in\mathcal{A}_{c}$) (L). Further it performs pruning of $\mathcal{A}_{p}$ using $\pi_{\text{neural}}$ (P), (ii) MCTS-P+L: Our approach without neural pruning and (iii) MCTS+P-L: No access to library of concepts during continual learning and thereby lacks ability to use macro actions. This method greedily selects the action from $\mathcal{A}_{p}$ as given by $\pi_{\text{neural}}$. We provide more details about the 3 methods in Appendix  D.4.
Metrics. We adopt the following metrics to evaluate our models: (i) Program Accuracy: A binary score obtained through human evaluation. (ii) Target Construction IoU: Intersection over Union (2D-IoU) between the bounding boxes (excluding depth). (iii) Target Construction Loss: Mean Squared Error (MSE) loss between bounding boxes (including depth). Note that both IoU and MSE are computed invariant to order between objects with same set of attributes.
6 Results
Our experiments evaluate the following (i) How does our model perform as compared to baselines in terms of concept learning and execution ability (In-Distribution) ? (ii) How does our model perform in comparison to baselines in ability to generalize to concept instances larger than those seen during training (Out-of-Distribution) ? (iii) How robust and efficient is our learning pipeline ? (iv) How can the acquired concepts be used in complex reasoning and planning tasks ?
6.1 Effectiveness of Concept Learning
Tables  2 and  2 compares the the program accuracy [^2] Purely neural approach is marked as NA due to absence of program representations. and IoU/MSE metrics on attaining the final state of the proposed and the baseline approaches in the in-distribution setting. The proposed SPG approach significantly improves over other approaches. We make the following observations. First, for complex compositional structures, the accuracy of the pre-trained models is poor (zero), indicating their inability to reason over numerous and complex spatial relations inherent in structures (e.g., staircases, bridge, x-shaped etc.). Additionally, we note that program inference via an LLM (GPT4) shows improved performance learning simple structures (e.g., rows or columns) using the tokenized language description of the scene compared to a VLM (GPT-4V) provided with a visual input (reflecting in better IoU/MSE scores). However, we observe that for complex structures, the performance of VLM (GPT-4v) is better than an LLM (GPT) indicating the difficulty in complex spatial reasoning inherent in interpreting complex concepts. Purely neural approach that is completely grounded in the demonstration performs much better on complex structures as compared to pre-trained foundation models. However our approach still outperforms it and is more data efficient.
Table 1: Program Accuracy (PA)
Model Simple Complex
SPG(Ours) 1.00 0.83
GPT-4V 0.33 0.00
GPT-4 0.78 0.00
SD+G NA NA
SD NA NA
Table 2: In-distribution Performance (Mean $\pm$ Std-error)
Model Simple Complex
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.96 $\pm$ 0.00 0.01 $\pm$ 0.00 0.85 $\pm$ 0.02 2.06 $\pm$ 1.02
GPT-4V 0.75 $\pm$ 0.01 4.33 $\pm$ 0.41 0.50 $\pm$ 0.02 7.29 $\pm$ 1.10
GPT-4 0.89 $\pm$ 0.01 1.36 $\pm$ 0.26 0.28 $\pm$ 0.02 13.5 $\pm$ 1.65
SD+G 0.74 $\pm$ 0.01 1.42 $\pm$ 0.29 0.61 $\pm$ 0.02 2.43 $\pm$ 0.48
SD 0.49 $\pm$ 0.01 1.48 $\pm$ 0.24 0.46 $\pm$ 0.02 3.71 $\pm$ 1.53
6.2 Generalization Performance
Table  4 compares the generalization performance on Dataset III [^3] Full table with standard-errors reported in Appendix,  8. The proposed model outperforms other approaches. Further, we observe a relative decrease in 2D-IoU being 7.27% for simple and 5.74% for complex structures. In contrast Struct-Diffusion with grounder shows a relative decrease of 63.25% on simple structures and 74.72% on complex structures; highlighting their inability to generalize inductively. Pre-trained models on the other hand have much higher relative decrease in performance for complex structures (53.87% and 41.64%), which is attributed to their inability generate the correct program that can generalize inductively to unseen data. Table 9 in Appendix quantifies relative performance decrease values as compared to Dataset I.
Table 3: Out-of-Distribution Performance
Model Simple Complex
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.89 0.43 0.80 1.49
GPT-4V 0.58 13.2 0.29 10.9
GPT-4 0.78 5.51 0.13 19.1
SD+G 0.27 6.21 0.15 14.2
SD 0.24 6.86 0.15 11.6
Table 4: Perf. on Dataset II: Names Reversed
Model Simple Complex
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.86 1.74 0.78 3.93
GPT-4 0.78 3.16 0.00 22.73
GPT-4V 0.71 3.92 0.09 21.29
6.3 Robustness and Efficiency Analysis
6.3.1 Reliance on pre-trained Knowledge vs. Demonstration.
Next, we evaluate the degree to which concept learning relies on background knowledge vs. observed action sequences for its construction. We compare pre-trained models against our approach by learning programs on another training dataset differing only in terms of the names of the structures (names being reversed). Table  10 in the Appendix gives the corresponding program accuracies. Table  4 indicates the corresponding IoU and MSE, with our approach again showing the best performance [^4] For the IoU/MSE values along with standard errors refer to table  11 in Appendix.. The relative decrease in performance for our approach (12%) is lower than GPT-4(14%) and much lower than GPT-4-V(30%) (in terms of program accuracy). The poorer generalization of pre-trained models can be attributed to over-reliance prior knowledge and failing to effectively use demonstrations.
6.3.2 Analysis of MCTS Variants for Concept Learning
Figure 4: MCTS Variants. Num. of expansion steps in search (log scale) (X-axis) vs Program accuracy (Y-axis).
Figure  4 compares the program accuracy for the three methods of plan search. For the third method the program accuracy is expected to be independent of expansion steps as it greedily chooses the action for which $\pi_{\text{neural}}$ gives the highest probability. For the MCTS+L based methods the accuracy increases beyond 0.6 with time, which demonstrates that having a composable library of concepts allows us to learn a much richer class of inductive concepts. MCTS+P+L saturates to a program accuracy of 0.933 in 4000 expansion steps, whereas MCTS-P+L takes 512000 expansion steps, which demonstrates a significant increase in learning efficiency via use of a neural prunner. For very low number of expansions steps (<40) accuracy of MCTS+L based methods is lower than MCTS-L as they spend some expansion steps on exploration based on UCB bound (rather than taking greedy choice).
6.3.3 Effectiveness of Continual Learning of Visual Concepts
Figure 5: Acquisition of new visual concepts. Likelihood of correct grounding of the referred object as training proceeds.
Having a disentangled representations allows us to 1) intersperse learning of new visual attributes with learning of new inductive concepts 2) avoid catastrophic forgetting of already learnt attributes. For example, the model can learn the chocolate from an instruction “construct a tower using chocolate blocks of size 4”, even if it has not seen the color in the pretraining phase. Because of our modular architecture, we can learn the color as a new embedding in the space of visual attributes. Plot  5 demonstrates the benefit of having such disentangled representations. As the training proceeds the probability of selecting the chocolate block increases with time. Further, the likelihood of selecting a magenta colored object (in a different scene) is unchanged and the problem of catastrophic forgetting is avoided. Additional details for continual learning of visual concepts appear in Appendix  D.2.
6.4 Application of Learnt Concepts for Embodied Instruction Following
Complex instruction execution via LLM. We demonstrate our ability to build over the learnt program representations and perform complex guided robot manipulation tasks, in a simulated Franka Panda environment. E.g., consider instructing a robot to : “Construct a tower of green die having the same height as the existing tower of white die.” or “Construct a tower of total 6 blocks using alternating blue and red blocks”. For both the above tasks we prompt GPT-4 by providing it with the set of inductive concepts that are already learnt, the set of primitive actions, and some pre-defined helper functions by using Python import statements in a manner similar to [code-as-policies]. GPT-4 generates an executable Python code in terms of these functions which on running generates the resultant trajectory. Top portion of figure  6 shows the execution for both these tasks. We provide details in Appendix  D.3.
Grounding learnt concepts into visual input for plan synthesis. We further demonstrate that the concepts we have acquired can help us to perform goal conditioned planning. Bottom portion of figure  6 demonstrates the results of our approach for the task of constructing a staircase beginning from an adversarial and assistive initial state. Note, the planner that we learn is a grounded neuro-symbolic planner, as a PDDL based planner cannot be hand-coded easily ( D.5), and LLMs/VLMs are unable to perform such complex reasoning tasks. Relevant details are provided in the Appendix  D.5
Figure 6: Application of learnt concepts. Top: Using LLM to generate the executable code for a novel tasks, given the concept definitions. Bottom: Integrating a neuro-symbolic planner over the concepts. In the bottom-left figure the planner optimally outputs the sequence of actions to replace the green cube. In the bottom-right, the planner is able to complete a staircase from an initially constructed row.
7 Limitations
Currently our methodology and evaluation setup assumes that each object is visible in the initial scene and does not handle partial observability. We assume all the objects are cubes of similar size in order to ensure dynamic stability of structures while stacking. During training we assume tracking and the ability to associate object with corresponding instances. We assume that the demonstration data given for learning is free of errors.
8 Conclusions
This paper introduces a novel approach for learning inductive spatial concepts as neuro-symbolic programs from language-annotated demonstrations. We observe that direct generation of a program via an LLM/VLM does not generalize to new concepts and purely neural approaches fail to learn the inherent inductive prior inherent in such concepts. Our approach factors program learning as: Sketch: generating the high-level program signature via an LLM, Plan: searching for a grounded plan that maximises the total discounted reward with the respect to the demonstration, and Generalize: abstracting the grounded plan into an inductively generalize-able abstract plan via an LLM. Continual learning is achieved via learning of modular programs by giving preference to shorter programs (through composition of learnt ones) and allowing for learning of primitive neural concepts via gradient updates. Extensive evaluation demonstrates accurate program learning and stronger generalization in relation to baselines. Further, we show the grounding of learned concepts in visual data to facilitate reasoning and planning for embodied instruction following.
[…]
Appendix A Additional Details on Technical Approach
Figure  7 illustrates the pipeline for learning inductive concepts.
Figure 7: SPG: Concept Learning Learning of grounded inductive programs is factored into three phases (a) Sketch: generating a sketch for the task intended in the instruction in the context of the environment (b) Plan: searching for the likely plan (guided by physical rewards) to construct the given assembly and (c) Generalize: generating a program representation to explain a class of demonstrations. Finally, our architecture allows learning of increasingly complex symbolic abstractions composed of simpler ones.
Figure  8 illustrates the pipeline for online inference to realize to realize construction of novel structures.
Figure 8: SPG: Inference First the library of concept $\mathcal{L}$ is loaded with the corresponding set of learnt programs. Then the given instruction is converted into task-sketch $H_{s}$, which is grounded in the initial scene. The required program is fetched from the library, and the grounded task-sketch is executed based on the semantics of the learnt program.
A.1 Symbolic Constructs and their Semantics used in Programs
Table 5 defines the types of the signature and semantics of all the operators. Table 6, includes the type definition of various symbols. Standard Python constructs such as for loops, if else $\cdots$) as assumed in addition to the constructs defined here.
A.2 Ground-Truth Inductive Concepts
1######
2# row
3def row(length, objects):
4 for i in range(length):
5 keep_at_head(obj = objects)
6 move_head(dir = "right")
7
8######
9# tower
10def tower(height, objects):
11 for i in range(height):
12 keep_at_head(obj = objects)
13 move_head(dir = ’top’)
14
15######
16# column
17def column(size, objects):
18 for _ in range(size):
19 keep_at_head(obj = objects)
20 move_head(dir = ’front’)
21
22######
23# staircase
24def staircase(steps, objects):
25 for step in range(1, steps+1):
26 tower(height = step, objects = objects)
27 move_head(dir = ’right’)
28
29######
30# inverted_row
31def inverted_row(num, objects):
32 for i in range(num):
33 keep_at_head(obj=objects)
34 move_head(dir=’left’)
35
36######
37# inverted_column
38def inverted_column(size, objects):
39 for _ in range(size):
40 keep_at_head(obj = objects)
41 move_head(dir = ’back’)
42 return None
43
44######
45# inverted_staircase
46def inverted_staircase(steps, objects):
47 for step in range(1, steps+1):
48 tower(height = step, objects = objects)
49 move_head(dir = "left")
50
51######
52# diagonal_135
53def diagonal_135(length, objects):
54 for i in range(length):
55 keep_at_head(obj = objects)
56 move_head(dir = ’front’)
57 move_head(dir = ’left’)
58 return
59
60######
61# diagonal_315
62def diagonal_315(length, objects):
63 for i in range(length):
64 keep_at_head(obj = objects)
65 move_head(dir = ’back’)
66 move_head(dir = ’right’)
67 return
68
69######
70# diagonal_225
71def diagonal_225(length, objects):
72 for _ in range(length):
73 keep_at_head(obj = objects)
74 move_head(dir = ’back’)
75 move_head(dir = ’left’)
76
77######
78# diagonal_45
79def diagonal_45(length, objects):
80 for _ in range(length):
81 keep_at_head(obj = objects)
82 move_head(dir = ’front’)
83 move_head(dir = ’right’)
84
85######
86# boundary
87def boundary(size, objects):
88 row(length=size-1, objects=objects)
89 for _ in range(size-1):
90 move_head(dir = ’right’)
91 move_head(dir = ’front’)
92
93 column(length=size-1, objects=objects)
94 for _ in range(size-1):
95 move_head(dir = ’front’)
96 move_head(dir = ’left’)
97
98 inverted_row(length=size-1, objects=objects)
99 for _ in range(size-1):
100 move_head(dir = ’left’)
101 move_head(dir = ’back’)
102
103 inverted_column(length=size-1, objects=objects)
104 for _ in range(size-1):
105 move_head(dir = ’back’)
106 move_head(dir = ’right’)
107
108######
109# arch_bridge
110def arch_bridge(height, objects):
111 staircase(steps = height, objects = objects)
112 move_head(dir = ’left’)
113 inverted_staircase(steps = height, objects = objects)
114 return
115
116######
117# x-shaped structure
118def x(size, objects):
119 diagonal_45(length = size, objects = objects)
120 move_head(dir = ’back’)
121 diagonal_315(length = size, objects = objects)
122 move_head(dir = ’left’)
123 diagonal_225(length = size, objects = objects)
124 move_head(dir = ’front’)
125 diagonal_135(length = size, objects = objects)
126
127######
128# pyramid
129def pyramid(height, objects):
130 for i in range(height):
131 row_length = (height * 2) - (i * 2) - 1
132 row(length = row_length, objects = objects)
133 if i != height - 1:
134 move_head(dir = ’top’)
135 move_head(dir = ’right’)
136######
Listing 1: Definition of inductive concepts
Table 5: Symbols and Semantics Signature and semantics for the primitive concepts and operations that are used in the construction of the programs used to express inductive spatial concepts.
Function Signature Semantics
filter (VisualConcept, ObjSet) $\rightarrow$ ObjSet Returns the objects that contain the VisualConcept
move_head (Head, Dir) $\rightarrow$ Head Moves the head to the given direction (May or may not take input/return the head, based on a flag)
assign_head a.k.a move_head(overloaded) (Head, ObjIdx) $\rightarrow$ Head Given the index/one-hot representation for an object, it moves the head to the position corresponding to that object
keep_at_head (ObjSet, Head) $\rightarrow$ None Keeps the argmax of ObjSet at the head
reset_head None $\rightarrow$ Head Sets the head to the top position of stack and pops this position from the stack as well
store_head Head $\rightarrow$ None Pushes the current position of head into the stack
Table 6: Symbolic representation. The table lists the type definitions used in the implementation of SPG programs.
Defined Types Python Type Usage
IntArg int Argument for the structures that defines the size (height, length etc)
Obj torch.Tensor One-hot vector whose non-zero index represents the selected objects
ObjSet torch.tensor Probability mask over the selected objects
Dir string Primitive directions like left, right, front, top, etc
ConceptName string name of the visual, action or inductive concept
Head torch.Tensor Bounding box with depth. 3D cuboidal space.
A.3 Curriculum Learning
We follow a curriculum approach where the visual concepts are trained first from simpler linguistically-described demonstrations ($t_{0}$ in figure  9). This is followed by learning of action concepts through sequentially composed pick and place tasks. Its essential to use such long range sequential instructions in order to ensure that the semantics of action concepts are learnt for placement of objects at a height much above the tabletop ($t_{1}$ in figure  9). After the pre-training phase, the agent can continually learn new inductive concepts and visual attributes. ($t_{2},t_{3}$ in figure  9)
Figure 9: Continual learning through curriculum: Using simple pick and place demonstrations we learn visual attributes such as blue cube, yellow lego ($t_{0}$). Using long range instructions which are sequentially concatenated descriptions of pick and place tasks we train our action concepts such as left, right, top ($t_{1}$). After pre-training the agent can perform continual learning of concepts such as learning generalized representation for tower ($t_{2}$). Because of disentangled representation of neural and symbolic concepts, interspersed learning of new visual attributes such as chocolate color are also possible through few demonstrations of structure creation($t_{3}$).
A.4 Details for Plan-Search and Generalization
Plan-Search: We present additional details on the MCTS procedure for searching for a plan conditioned on a program signature and guided by the demonstration.
In order to incorporate the objective of searching for physically realizeable plans and facilitating generation via re-use of concepts, the following conceptual changes are incorporated in the standard MCTS procedure [sutton2018reinforcement].
Modularity (MCTS + L): We want to allow learning of novel inductive concepts in terms of existing ones. This would ensure that the plan $H_{P}^{*}$ corresponding to a given demonstration is concise and can be easily generalized to the $H_{G}^{*}$.  C.3 in appendix give example of two plans for the structure Pyramid one which is modular and can be successfully generalized by GPT-4, other for which GPT-4 fails in generalization due to lack of modularity. This can be seen as a form of regularization in terms of the length of concept description, by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2)
$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))+\alpha\log|H|\right]$ (4)
In order to allow modular learning of programs, for every inductive concept already stored in the library we define corresponding action instantiations which can be a potential candidate actions during our search. As an example, for the concept Tower we have one of the action instantiation as Make_Tower(3, objects) which would be the action of constructing desired tower. This can be visualized as a compound/macro-action which is composed of primitive actions keep_at_head(objects), move_head(‘top’). We define $\mathcal{A}_{c}$ as the space of such compound actions, and $\mathcal{A}_{p}$ as the space of primitive action/function consisting reset_head(), move_head(direction), keep_at_head(objects), store_head(). Realization of equation (6) (increased preference of macro-actions over primitive ones) is done through discounted IoU rewards during our search (Make_Tower(3) would have a reward of 1+1+1, as compared to 1 + $\gamma^{2}$(1) + $\gamma^{4}$(1) for a sequence of 3 (keep_at_head(objects), move_head(’top’)0)). We refer to MCTS using concept instantiations from $\mathcal{L}$ as macro actions in search as MCTS+L.
Scalablility (MCTS+P): as more concepts are added to the library $\mathcal{L}$ the action space of our search $\mathcal{A}$ (specifically $\mathcal{A}_{c}$, the space of compound concepts) increases, therefore we want to prune the search space effectively. For this during the pre-training phase we train a reactive policy $\pi_{neural}$ which given the current state, $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (part of the demonstration) would output one of the primitive action, $a_{t}^{*}\in\mathcal{A}_{p}$ where $\mathcal{A}_{p}$ is the primitive action space, i.e. $a_{t}^{*}=\pi_{neural}(a_{t}|\tilde{s_{t}},{s_{t+1}}),\;\;a_{t}\in\mathcal{A}_{p}$ Note that while expanding our search tree we only search among the space of compound actions $\mathcal{A}_{c}$ and the action $a_{t}^{*}$, thereby reducing the branching factor of search from $|\mathcal{A}_{c}\cup\mathcal{A}_{p}|$ to $|\mathcal{A}_{c}|+1$. We refer to MCTS using neural pruning as MCTS+P. Therefore our MCTS algorithm is modular through hierarchical composition of learnt concepts and efficient through pruning of action space and is referred to as MCTS+L+P.
Modifications to the Simulation and Reward Back propagation Steps: Next, we outline the modifications in the simulation and the reward back propagation steps of the standard MCTS algorithm for our setting. During program search we assume access of intermediate scenes in the demonstration. This allows us to provide intermediate rewards that can guide the search well. We observed that making the following changes in simulation and back propagation step increased the efficiency of our search procedure.
  • Simulation: Rather than performing Monte Carlo simulations at each newly expanded leaf node (to estimate its value) we completely avoid these simulation steps. This was motivated by the fact that our reward is not completely sparse and the intermediate IoU rewards for each object we place allow us to guide the search effectively.
  • Back propagation: We perform off policy Q-learning updates during back propagation similar to one indicated by [complexmctsbackup] :
    $V(s_{t})=max_{a\in\mathcal{A}}Q(s_{t},a).$ (5)
    $\tau(s_{t},a)=s_{t+1}$ (6)
    $Q(s_{t},a)=r_{t}+\gamma V(s_{t+1})$ (7)
Generalization: GIven multiple equal length plans for a given demonstration, we seek to recover a plan one that can be easily generalized by the LLM.  6 shows a plan which could be correctly abstracted out into a generic program by GPT-4. Whereas  7 shows another plan with similar semantics, for which GPT-4 is unable to correctly find the generalized program (Note that row or column of size 1 is equivalent to keep_at_head). We tackle this problem in the following manner.
  1. 1.
    Rather than getting a single plan from the plan search we get the top k plans $\{H_{P,i}\}_{i=1}^{i=k}$. In order to get these top k plans we expand the complete tree (based on UCB criteria) starting from the root node corresponding to the initial state, till a predefined budget of expansions. Then we select the top k paths(potential plans) from the root node to all the leaf nodes (where the top k ones are those that give the highest accumulated IoU reward with respect to the given demonstration).
  2. 2.
    Later we abstract out each of these k plans into corresponding generalized programs, $\{H_{G,i}\}_{i=1}^{i=k}$ using GPT-4. We again run each of these programs on the given demonstration and then choose the one which gives the highest IoU reward (resolving ties based on predefined order). Note that some program $H_{G,i}$ upon execution may result in a plan $\tilde{H}_{P,i}$ different from the original plan ${H}_{P,i}$ using which it was generalized. This can be attributed to potential errors in GPT-4s program generalization process.
A.5 Additional Details: Learning with Increasing Number of Demonstrations
Given k demonstrations for a novel inductive concept, we independently find k task sketch $\{H_{S,i}^{*}\}_{i=1}^{i=k}$ and grounded plans $\{H_{P,i}^{*}\}_{i=1}^{i=k}$. During the generalization phase we give these k pair of task-sketch and corresponding plans to GPT-4 and ask into infer a single abstraction over them.  C.5 in appendix gives a concrete example. Equation for generalize step (getting $H_{G}^{*}$ from $H_{P}^{*},H_{S}^{*}$) can be modified as follows.
$H_{G}^{*}\leftarrow Generalize(H_{G}\,|\,{\{H_{P,i}^{*},\,H_{S,i}^{*}\}}_{i=1}^{i=k};\,{\theta_{G}}),\;\;H_{G}\in\mathcal{H}^{\mathcal{G}}$ (8)
Appendix B Additional details regarding datasets
Figure  10 demonstrates the kind of inductive concepts for which we want to learn generic (i.e instance agnostic) representations.
Figure 10: Illustration of the inductive concepts.
Dataset for Pre-training: We use 5k examples of constructing twin-towers i.e. 2 towers adjacent to each other, for learning semantics of move_head(dir), a basic set of visual attributes, reactive policy $\pi_{\text{neural}}$, and neural modules required for grounded planning. The twin towers allow us to learn various action semantics for all possible configurations of blocks in 3D-space (and not being limited to blocks placed directly on table top surface). Since we are not aware of the underlying semantics of tower during pre-training phase the corresponding natural language instruction consists of step by step pick and place actions.  11 gives example demonstrations from this dataset.
Figure 11: Example from pre-training dataset.
Dataset for Inductive Structures: We learn a variety of structures which we have divided into Simple and Complex structures. A structure is considered complex if it can be expressed as an inductive composition of simpler structures. As an example, we can express a staircase to be a composition of towers of increasing height. The structures are listed in the Table 7. Figure  12 shows the hierarchical relationship among these structures in the form of a DAG (directed acyclic graph).  A.2 gives the ground truth program representations for each structure.
Table 7: Structure Types. Examples of simple and complex structures considered in this work for the robot to construct.
Simple Structures Complex Structures
Row, Column, Tower X (cross-shape), Staircase
Inverted-Row, Inverted-Column Inverted-Staircase, Pyramid
Diagonal-45, Diagonal-135 Arch-Bridge, Boundary
Diagonal-225, Diagonal-315
Figure 12: Hierarchy of the structures/programs. This diagram shows the hiearchical nature of the structures in our dataset. MH is abbrevation for move_head and KH is abbreviation for keep_at_head
Appendix C Prompting Strategy and Examples
This section gives various prompting examples for our approach and baselines, along with examples motivating particular design decisions in our approach.
C.1 Prompt Example for Task Sketch Generation Stage (Sketch)
In order to get a program representation (high level task sketch) of the given natural language instruction, we prompt GPT-4 with few shot examples in a manner similar to [code-as-policies]. Code segment  2 gives an example of getting the task sketch given the demonstration for constructing a staircase. We first import the available primitive operators and functions and also give examples in order to demonstrate the signature of the available primitives(line 1-8). Then we give incontext examples of how to parse various natural language instructions in a program representation(line 10-14). We append to this prompt the instruction for current task(line 16-17).
1# importing the available functions
2from visual_operators import filter
3from inductive_operators import get_parameters, find_structure
4
5# function signature of the imported functions
6filter(color, cube) # filter the objects that are cubes and color ..
7get_parameters(structure) # parameters of the structure ...
8find_structure(type, description) # finding structure of given type, description
9
10# examples:
11# instruction: Find the tower with green cubes
12find_structure(type = tower, description = filter(green, cube))
13# instruction: Construct a tower of height 3 with yellow cubes
14Tower(height = 3, objects = filter(yellow, cubes))
15
16# current task: Construct a staircase of 4 steps using cyan legos
17
18# (GPT-4s output)
19Staircase(steps = 4, filter(cyan, legos))
Listing 2: Task Sketch Generation Using GPT-4 (Sketch)
C.2 Prompt Example for Generalizing a sequence of actions/plan to a general program (Generalize)
Code segment  3 give an example of getting the general Python program from the plan found using MCTS. We first provide a base prompt giving details to GPT-4 about the desired task (line 1-2). Then we give the input arguments and the corresponding output/plan for a given demonstration (line 3-5). We expect the GPT-4 to output the final Python program (line 7-10)staircase (line 6-8).
1# Write a general python code which on the given input produces the desired output, do not output anything other than the function description.
2
3# input: n = 4, objects = ObjSet
4# output: tower(1, ObjSet), move_head(‘right’), tower(2, ObjSet), move_head(‘right’), tower(3, ObjSet), move_head(‘right’), tower(4, ObjSet)
5
6# Program (GPT-4s output)
7def staircase(n, objects):
8 for i in range(n):
9 tower(i+1, objects)
10 move_head(’right’)
Listing 3: Plan to Program using GPT-4 (Generalize)
C.3 Benefit of Estimating Modular/Smaller Plans
The below examples demonstrate the benefit of learning new inductive concepts in terms of already acquired inductive concepts (more modular representation). The program  4 is obtained through generalization of plan $H_{p}$ that represents pyramid in terms of rows of decreasing size. The generated program generalizes to pyramid of different height.
1# input: n = 3, objects = ObjSet
2# output: row(5, ObjSet), move_head(’right’), move_head(’top’), row(3, ObjSet), move_head(’right’), move_head(’top’), row(1, ObjSet)
3
4# Program (GPT-4s output)
5def pyramid(n, objects)
6 for i in range(n, 0, -1):
7 # Calculate the number of objects in the current row
8 row_count = 2 * i - 1
9
10 # Perform actions for the current layer
11 row(row_count, objects)
12 if i > 1: # No need to move head after the last row
13 move_head(’right’)
14 move_head(’top’)
Listing 4: Pyramid in terms of rows
On the other hand program  5 which is obtained in an attempt to generalize the plan expressed in terms of primitives is incorrect.
1# input: n = 3, objects = ObjSet
2# output: store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’top’), move_head(’right’), store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’right’), move_head(’top), keep_at_head(ObjSet)
3
4# Program (GPT-4s output)
5def pyramid(n, objects)
6# Store initial position
7 store_head()
8
9 # Initial sequence of placing and moving right
10 for _ in range(n):
11 keep_at_head(objects)
12 move_head(’right’)
13
14 # Reset and move to top right position
15 reset_head()
16 move_head(’top’)
17 move_head(’right’)
18
19 # Store again and repeat placing and moving right
20 store_head()
21 for _ in range(n):
22 keep_at_head(objects)
23 move_head(’right’)
24
25 # Reset and adjust final position
26 reset_head()
27 move_head(’right’)
28 move_head(’top’)
29 keep_at_head(objects)
Listing 5: Pyramid in terms of primitive actions
C.4 Comparing Same-length Plans in terms of Generalizability
Below we illustrate plan to program generalization for 2 different plans which are correct and have same length, but LLMs find it difficult to generalize the second one as compared to first. (Note that row and column of size 1 are equivalent to keep_at_head)
1# input: n = 3, objects = ObjSet
2# output: keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet)
3
4
5# Program
6def tower(n, objects):
7 for _ in range(n):
8 keep_at_head(objects)
9 move_head(’top’)
Listing 6: Plan for tower that can be easily generalized (correct generalization)
1# input: n = 3, objects = ObjSet
2# output: row(1, ObjSet), move_head(’top’), keep_at_head(1, ObjSet), move_head(’top’), column(1, ObjSet)
3
4
5# Program
6def tower(n, objects):
7 for i in range(1, n + 1):
8 row(i, objects)
9 move_head(’top’)
10 keep_at_head(objects)
11 move_head(’top’)
12 if i < n:
13 column(i, objects)
14 move_head(’top’)
Listing 7: Plan for tower that is difficult to generalize (Incorrect generalization)
C.5 Generalizing via Multiple Demonstrations
Given multiple demonstrations we independently find task sketch and corresponding grounded plans for each demonstration. These are further given to GPT-4 for generalization. Code segment  8 gives an example of getting a single Python program from multiple demonstrations. Note that we explicitly prompt the LLM that some of the grounded plans might be incorrect (which may lead to more robust generalization in case of noisy demonstrations).
1
2# Function Call: wor(height = 3, objects = ObjSet_1)
3# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
4# Function Call: wor(height = 3, objects = ObjSet_1)
5# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
6# Function Call: wor(height = 3, objects = ObjSet_1)
7# Execution: column(size=1, obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
8
9#Write the function definition, which generalizes the above executions. Note that some of the executions can be partially wrong.
10‘‘‘python
11def wor(height, objects):
12‘‘‘
13
14GPT-4s Output .....
Listing 8: Generalizing through multiple plans
C.6 Prompt Examples for Learning Programs using LLM/VLM Models
Below we describe the prompting methodologies for learning programs through LLM/VLM models. Note that although the prompt examples described below are for the case of learning novel structure from 1 demonstration, we use 3 demonstration per novel structure in our main results (for both our approach and LLM/VLM baseline).
LLM/GPT-4 Code segment 9 depicts our prompting methodology given a demonstration for a new concept tower. For this baseline we aim to check demonstration following and spatial reasoning abilities of LLMs (GPT-4). We provide supervision of the intermediate scenes by using tokenized spatial relations between objects in the scene (Given in the form of Scene = [right(1, 0) …]). We further assume that only those objects that are required to perform the task are present in the scene (no distractor objects). For every structure (that needs to be learned at time t) we give LLM a prompt providing in-context example on how to generalize (line 19-35), the set of primitive operators (line 4) available and the set of structures learnt/present in library (till time t-1) (line 5-18). Finally we append to this prompt the expected declaration (arguments and keywords arguments) of the inductive concept that is to be learnt along with the spatial relations for each scene of the given demonstration (36-51). Note that we assume absence of distractor objects for this baseline.
1# Consider a block world domain ...... Given a structure creation task along with intermediate scnes complete a general Python function for it. The function should be in terms of primitive operators and already learnt structures that are present in the program library. Enclose the function within backtick (‘‘‘)
2
3primitive_operators = [keep_at_head, move_head ..]
4# this would be our program library
5learnt_structures = {
6 "row": {
7 "program_tree":
8 ’’
9 def row(size, objects):
10 for i in range(size):
11 keep_at_head(obj = objects)
12 move_head(dir = right’)
13 ’’,
14 },
15
16 ......
17}
18# the example task
19Example task:- Place all the objects to the right of each other.
20Final state :- [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
21Intermediate scenes :-
22Scene 0 = []
23Scene 1 = []
24Scene 2 = [right(1, 0)]
25Scene 3 = [right(1, 0), right(2, 1)]
26Scene 4 = [right(1, 0), right(2, 1), right(3, 2)]
27Scene 5 = [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
28Python function :-
29‘‘‘python
30def placing_all_right(objects):
31 for i in range(len(objects)):
32 keep_at_head(objects) # select one object from the objects set and keep the head at this location
33 move_head(dir = ’right’) # move the head to the right of the previous position
34‘‘‘
35# The current task for which program needs to be found
36Current task:- Construct a tower of size 6.
37Final state :- [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
38Intermediate scenes :-
39Scene 0 = []
40Scene 1 = []
41Scene 2 = [top(1, 0)]
42Scene 3 = [top(1, 0), top(2, 1)]
43Scene 4 = [top(1, 0), top(2, 1), top(3, 2)]
44Scene 5 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3)]
45Scene 6 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
46Python function :-
47‘‘‘python
48def tower(size, objects):
49??
50‘‘‘
Listing 9: Prompting Strategy for LLM baselines (GPT-4)
VLM/GPT-4-V Unlike LLM, VLMs have the abilities to process the demonstration as a sequence of visual frames. Therefore rather than providing the symbolic spatial relations between every scene we instead directly provide all the intermediate scenes for the given demonstration. Further we also relax the assumption that there are no distractor objects. As shown in figure 13 We first give information about the set of primitive operators and the structures that we have already learnt (library of concepts). In order to visually ground the semantics of our primitive actions we give 3 example tasks (natural language instruction and intermediate scenes) that do not directly refer to any structure, along with corresponding sequence of actions taken (# Demonstration for visual grounding). We further provide another example (without scenes) demonstrating how to write generalizable Python function for a given task using our operators (# Example for generalization). Finally we give the natural language instruction and corresponding scenes for the current task along with signature of the program to be learnt (# Current task description).
Figure 13: Prompt example for VLM-baseline. Figure shows the prompting strategy for GPT-4-V that includes the primitive actions, the example tasks for visual grounding, example of writing generalizable Python functions and the demonstration frames.
Appendix D Supplementary Results
Out-of-Distribution Performance:
Table 8: Out-of-Distribution Performance (mean $\pm$ std-error)
Model Simple Complex
IoU MSE IoU MSE
SPG(Ours) 0.892 $\pm$ 0.065 4.386e-4 $\pm$ 2.387e-4 0.804 $\pm$ 0.025 0.001 $\pm$ 5.391e-4
GPT-4 0.776 $\pm$ 0.023 0.006 $\pm$ 0.001 0.131 $\pm$ 0.019 0.019 $\pm$ 1.498e-3
GPT-4V 0.575 $\pm$ 0.026 0.013 $\pm$ 0.001 0.290 $\pm$ 0.016 0.011 $\pm$ 1.316e-3
SD 0.236 $\pm$ 0.005 0.006 $\pm$ 7.495e-4 0.150 $\pm$ 0.011 0.011 $\pm$ 2.860e-3
SD+G 0.273 $\pm$ 0.004 0.006 $\pm$ 6.180e-4 0.154 $\pm$ 0.010 0.014 $\pm$ 2.958e-3
Relative decrease in Out-of-Distribution performance:
Table 9: Relative Decrease in IoU
Model Simple (%) Complex (%)
SPG(Ours) 7.27 5.74
GPT-4 12.61 53.87
GPT-4V 23.33 41.64
SD 51.84 67.67
SD+G 63.25 74.72
Performance Dataset II (i.e. name reversed evaluation): Table  10 gives the corresponding program accuracies, while Table  11 give the corresponding IoU/MSE metrics along with standard errors.
Table 10: PA (Reversed Names)
Model Simple Complex
SPG(Ours) 0.88 0.78
GPT-4 0.67 0.00
GPT-4V 0.23 0.00
Table 11: Performance on Dataset II(Names Reversed)
Model Simple Complex
IoU MSE (1e-3) IoU MSE (1e-3)
SPG(Ours) 0.86 $\pm$ 0.03 1.74 $\pm$ 0.45 0.78 $\pm$ 0.02 3.93 $\pm$ 1.09
GPT-4 0.78 $\pm$ 0.03 3.16 $\pm$ 0.51 0.00 $\pm$ 0.00 22.73 $\pm$ 1.48
GPT-4V 0.71 $\pm$ 0.01 3.92 $\pm$ 0.45 0.09 $\pm$ 0.02 21.29 $\pm$ 1.59
D.1 Qualitative Comparison between Purely-neural (Struct-Diff+Grounder) vs. Ours(SPG)
Figure  14 gives compares the qualitative results for our approach against Struct-Diffusion with grounder on both in-distribution, Dataset I and out-of-distribution (larger size), Dataset III. In in-distribution setting the our method performs slightly better in terms of structure creation for both simple and complex structures, but the difference is not significant. However for out-of-distribution setting structures created by our approach are much better than those created by Struct-Diffusion+Grounder. Further for this setting structure creation by Struct-Diffusion seems to be much worse for complex structures than simple ones.
Figure 14: Structure creation comparison between SPG(Ours), and Struct-Diff+Grounder
D.2 Continual Learning of Neural Concepts
Given demonstration for the task “Construct a tower of height 4 using chocolate cubes”, we would like to learn the neural embedding for the unknown color chocolate (where we assume that tower has already been learnt and stored in the library $\mathcal{L}$). First the instruction is converted into corresponding plan sketch $\mathcal{H}_{S}$ = tower(4, Filter(chocolate, cubes)), which is passed to the visual grounder. The grounder detects the presence of an unknown attribute chocolate as an argument to filter, and randomly initializes a new neural embedding for it. Using this new embedding along with the already present embedding of cube and the visual features found through ResNet-34, the quasi-symbolic executor outputs a grounded task-sketch. The executor executes the grounded task-sketch by getting the semantics of the underlying function i.e. tower from the library $\mathcal{L}$. MSE+IoU loss computed over the final scene obtained and the expected final scene is backpropogated through the network. Note that during backpropogation all the neural modules (action semantics, visual attributes, ResNet-34) are frozen, except for the newly initialized embedding for chocolate. For the purpose of differentiable sampling during tower construction we use gumbel-softmax [gumbel] with masking. Figure  15 illustrates our approach.
Figure 15: Continual learning of visual primitives
D.3 Details for Inference on Novel Tasks using an LLM
Below we show the [code-as-policies] inspired prompting methodology that we use to get the executable code corresponding to a language specified manipulation task. We initially begin by importing the helper functions, spatial direction, primitive functions, and learnt inductive concepts/structures (line 5-11). Then we give few examples for how to use and compose the various functions for different tasks (line 16-83). Finally we give the instruction of current task, and expect GPT-4 to output the corresponding executable code.
1# Given a task you have to provide Python code for executing the task
2
3# importing available functions
4
5from spatial_directions import top, front, back, left, right
6
7from primitives import assign_head, move_head, keep_at_head
8# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
9
10from helpers import find_size, filter
11from structures import row, column, tower
12
13
14#function signature of the imported functions
15# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
16filter(color, shape)
17
18# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
19find_size(struct_name = str_name, objects = ObjSet)
20
21# assigns the head to the location of the object
22assign_head(at_obj_loc)
23
24# moves the head in the given dir
25move_head(dir)
26
27# keeps the object obj at the head
28keep_at_head(obj)
29
30
31#Examples:
32#Instruction: Move the green block to the left of the red dice
33assign_head(at_obj_loc = filter(red, dice))
34move_head(left)
35keep_at_head(obj = filter(green, cube))
36
37# Instruction: Find the size of the tower made of yellow legos
38find_size(struct_name = tower, objects = filter(yellow, lego))
39
40#Instruction: Find the size of the row made of orange cubes
41find_size(struct_name = row, objects = filter(orange, cube))
42
43# Instruction: Find the size of the column made of cyan cubes
44find_size(struct_name = column, objects = filter(cyan, cube))
45
46# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
47assign_head(at_obj_loc = filter(red, dice))
48move_head(left)
49keep_at_head(obj = filter(green, cube))
50assign_head(at_obj_loc = filter(green, cube))
51move_head(top)
52keep_at_head(obj = filter(yellow, cube))
53
54# Instruction: Construct a row of green legos of length 3 to the right of the blue block
55assign_head(at_obj_loc = filter(blue, block))
56move_head(right)
57row(length = 3, objects = filter(green, legos))
58
59
60# Instruction: Construct a tower of size 3 using red cubes
61tower(height = 3, objects = filter(red, cube))
62
63# Instruction: Construct a row of size 5 using blue legos
64row(length = 5, objects = filter(blue, lego))
65
66# Instruction: Construct a column of size 6 using green die
67column(length = 6, objects = filter(green, dice))
68
69# Instruction: Place 3 green blocks so that one block is to the right of the other
70green_blocks = filter(green, block)
71keep_at_head(green_blocks)
72move_head(right)
73keep_at_head(green_blocks)
74move_head(right)
75keep_at_head(green_blocks)
76
77# Instruction: Place 3 red legos on top of one another
78red_legos = filter(red, lego)
79keep_at_head(red_legos)
80move_head(top)
81keep_at_head(red_legos)
82move_head(top)
83keep_at_head(red_legos)
84
85
86# CURRENT TASK
87# Instruction: Construct tower of white cubes to the same height as
88# existing tower of green die
89
90# GPT-4s output
91#First, we have to find the height of the tower of green dice,
92#then construct a tower of white cubes of the same size
93
94tower_size = find_size(struct_name = tower, objects = filter(green, die))
95tower(height = tower_size , objects = filter(white, cube))
Listing 10: Prompting method for the task of constructing tower of white cubes to the same height as existing tower of green die
1# Given a task you have to provide Python code for executing the task
2
3# importing available functions
4from spatial_directions import top, front, back, left, right
5from primitives import assign_head, move_head, keep_at_head
6# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
7from helpers import find_size, filter
8from structures import row, column, tower
9#function signature of the imported functions
10# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
11filter(color, shape)
12# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
13find_size(struct_name = str_name, objects = ObjSet)
14# assigns the head to the location of the object
15assign_head(at_obj_loc)
16# moves the head in the given dir
17move_head(dir)
18# keeps the object obj at the head
19keep_at_head(obj)
20
21#Examples:
22#Instruction: Move the green block to the left of the red dice
23assign_head(at_obj_loc = filter(red, dice))
24move_head(left)
25keep_at_head(obj = filter(green, cube))
26
27# Instruction: Find the size of the tower made of yellow legos
28find_size(struct_name = tower, objects = filter(yellow, lego))
29
30#Instruction: Find the size of the row made of orange cubes
31find_size(struct_name = row, objects = filter(orange, cube))
32
33# Instruction: Find the size of the column made of cyan cubes
34find_size(struct_name = column, objects = filter(cyan, cube))
35
36# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
37assign_head(at_obj_loc = filter(red, dice))
38move_head(left)
39keep_at_head(obj = filter(green, cube))
40assign_head(at_obj_loc = filter(green, cube))
41move_head(top)
42keep_at_head(obj = filter(yellow, cube))
43
44# Instruction: Construct a row of green legos of length 3 to the right of the blue block
45assign_head(at_obj_loc = filter(blue, block))
46move_head(right)
47row(length = 3, objects = filter(green, legos))
48
49
50# Instruction: Construct a tower of size 3 using red cubes
51tower(height = 3, objects = filter(red, cube))
52
53# Instruction: Construct a row of size 5 using blue legos
54row(length = 5, objects = filter(blue, lego))
55
56# Instruction: Construct a column of size 6 using green die
57column(length = 6, objects = filter(green, dice))
58
59# Instruction: Place 3 green blocks so that one block is to the right of the other
60green_blocks = filter(green, block)
61keep_at_head(green_blocks)
62move_head(right)
63keep_at_head(green_blocks)
64move_head(right)
65keep_at_head(green_blocks)
66
67# Instruction: Place 3 red legos on top of one another
68red_legos = filter(red, lego)
69keep_at_head(red_legos)
70move_head(top)
71keep_at_head(red_legos)
72move_head(top)
73keep_at_head(red_legos)
74
75
76# CURRENT TASK
77# Instruction: Construct a tower of height 6 using red and blue blocks that are alternating
78
79# GPT-4s output
80# Python code:
81
82# Define the red and blue blocks
83red_blocks = filter(’red’, ’block’)
84blue_blocks = filter(’blue’, ’block’)
85
86# Start at the bottom and alternate building the tower
87for i in range(6):
88 if i%2 == 0: # if the stack position is even
89 keep_at_head(red_blocks)
90 else: # if the stack position is odd
91 keep_at_head(blue_blocks)
92 if i != 5: # if not at the top of the tower
93 move_head(top)
Listing 11: Prompting method for the task of constructing tower of alternating red and blue cubes
To find the size of a given structure in the given scene we define the function find_size, which takes the name of structure, all the objects in the initial scene, mask of the objects (a distribution over the objects based on the attributes), and the initial state. (we assume that this function has access to the semantics of all the concepts learnt so far, through a transition function). Algorithm  1 gives the pseudocode for the function find_structure. Below we provide a brief explanation for it.
  1. 1.
    First we assign our head to every block in the available blocks (line 6)
  2. 2.
    Then we begin constructing/visualizing the corresponding structure from that block beginning with a size of 1. (line 7)
  3. 3.
    For each structure created/visualized we compare the blocks moved for the structure creation with corresponding blocks originally present in the scene, and perform a matching between these blocks and a subset of the blocks originally present (line 11-26).
  4. 4.
    If we are able to find a mapping for each moved block, such that each mapped pair has an IoU greater than a threshold, we increase the next potential size to test by 1 (line 26-27).
  5. 5.
    The final size is the size corresponding to 2nd last iteration, before termination (line 28).
  6. 6.
    We return the maximum of all the possible structures that are found (line 34)
Algorithm 1 Find Size
1:$name$: structure name, $objs$: object list, $state$: initial state, $mask$: object mask
2:Size of the maximum sized structure found
3:$found\leftarrow[]$
4:for each $cand$ in $objs$ do
5:     $size\leftarrow 1$
6:     $curr\leftarrow state.copy()$
7:     while true do
8:         $new\_state\leftarrow assign\_head(cand)$
9:         $vis\_state,num\_mov,rew\leftarrow transition($
10:          $new\_state,name,[(^{\prime}size^{\prime},size),(^{\prime}objects^{\prime},mask)])$
11:         $topk\leftarrow torch.topk(mask,num\_mov)$
12:         $possible\leftarrow true$
13:         $matches\leftarrow[]$
14:         for each $idx$ in $topk$ do
15:              $possible\leftarrow$ (idx   in  objs)
16:              $match\_ok\leftarrow false$
17:              for $m\_idx,obj$ in enumerate(state.state) do
18:                  $iou\leftarrow iou2d(vis\_state.state[idx],$
19:                   $state.state[m\_idx])$
20:                  $match\_ok\leftarrow(iou>0.75)and$
21:                   $(m\_idx\,in\,objs)$
22:                  if $match\_ok$ then
23:                       $matches.append(m\_idx)$
24:                       break
25:                  end if
26:              end for
27:              if not $match\_ok$ then
28:                  break
29:              end if
30:         end for
31:         if $possible$ then
32:              $size\leftarrow size+1$
33:         else
34:              $found.append((size-1,matches))$
35:              break
36:         end if
37:     end while
38:end for
39:return max(size for size, match in $found$)
D.4 Details on MCTS Variants for Plan Search
Here we provide the details for 3 different plan search methods, that search over the space $\mathcal{A}=\mathcal{A}_{c}\cup\mathcal{A}_{p}$
  • MCTS+L+P: This is the approach that we describe in section 4.2. For every concept say Tower $\in\mathcal{L}$ we have a corresponding set of macro action say Make_Tower(3, objects) (L). Further we use a neural pruner $\pi_{\text{neural}}$ that outputs a primitive action $a_{p}^{*}$ (given current state and next expected state). We only consider the actions $\mathcal{A}_{c}\cup\{a_{p}^{*}\}$ during our search from the given state. This helps to reduce the effective branching factor and allows to search longer length plans within the same computational budget.
  • MCTS+L-P: Here we do not use the reactive policy, therefore the branching factor for every node becomes $\mathcal{A}$ = $\mathcal{A}_{c}\cup\mathcal{A}_{p}$.
  • MCTS-L+P: We only search among the space of primitive actions i.e $\mathcal{A}_{p}$. Given a state $\tilde{s}_{t}$ and corresponding next expected state $s_{t+1}$ we greedily pick the action $a_{p}^{*}=\pi_{\text{neural}}(\tilde{s}_{t},s_{t+1})$. This method is much more faster than the previous 2 methods as there is no explicit search. However the corresponding policy is trained only to output an action $a\in\mathcal{A}_{p}$ and lacks the ability to output modular plans composed of macro actions such as Make_Tower(3, objects) $\in\mathcal{A}_{c}$ (the action space $\mathcal{A}_{c}$ is increasing with time and the architecture of network needs to be changed accordingly). As a result the plans found are not modular and difficult to generalize. Further, the reactive policy is not trained to output reset_head(), store_head() as additional annotated data is required in order to train a classifier over them. This further decreases the space of grounded plans (and therefore corresponding generic programs) such a policy can represent. Training a reactive policy that can handle actions such as reset_head() and an action space $\mathcal{A}_{c}$ that grows with time is part of future work.
D.5 Goal-conditioned Planning with Learnt Concepts
Why is it difficult to hand encode a PDDL for our domain? Most of the PDDL description of blocks world assume actions involving only the spatial relation onTop, which limits their applicability to describing different structures like row that need spatial relations like onRight. Further a single action might lead to varied effects/post-conditions based on the initial state. 16 gives 2 example of the same action moveOnTop(A, B) which would end up giving adding different number of spatial relations.
Figure 16: Difficult to encode post-conditions. Illustration of a domain where encoding a PDDL for direct planning is challenging.
Approach Overview. Given an instruction $\Lambda$ = “Construct a staircase of magenta die having 3 steps”, we first convert it into corresponding grounded task sketch $H_{S}^{*}$ = staircase(3, [3, 2, 1, $\cdots$]). Executing the corresponding program of staircase (by getting the semantics from the library $\mathcal{L}$) on the desired objects we get the expected final scene $S_{f}^{\prime}$ in bounding box space. The initial scene $S_{i}$ and expected final scene $S_{f}^{\prime}$ are converted into scene graph $SG_{i}^{\prime}$ and $SG_{f}^{\prime}$ (described in  16). The relations between the task relevant objects in $SG_{f}^{\prime}$ act as propositions/relations for goal check and the initial scene graph act as the initial state. Then a neuro-symbolic planner is used to obtain the optimal plan from the start state to a state that satisfies the goal. Below we also detail different aspects of the approach.
Scene-graph Extraction.  2 gives the algorithm used for generating scene graph from a given scene (set of bounding boxes). Suppose we need to check whether there exists a relation of the form $(i,j,direction)$ i.e. block $i$ is in the direction $direction$ of block $j$, in a given scene. We first initialize the head at the position/bounding-box of block $j$ (line 7). Then we move the head in direction $direction$ (line 9). We claim that the relationship would exists if bounding box for block $i$ has IoU $>$ 0.75 with the predicted_head (line 10-12).
Algorithm 2 Get Scene Graph
1:procedure get_scene_graph($bboxes$) // bboxes describe the corresponding scene for which scene graph is to be found
2:     $spatial\_relations\leftarrow[]$
[…]
4:     for $i,bbox$ in $enumerate(bboxes)$ do
5:         for $j,other\_bbox$ in $enumerate(bboxes)$ do
6:              $final\_head\leftarrow bboxes[i]$
7:              $init\_head\leftarrow bboxes[j]$
8:              for $direction$ in $directions$ do
9:                  $pred\_head\leftarrow\text{move\_head}(direction,init\_head)$
10:                  $iou\_score\leftarrow IoU(final\_head,pred\_head)$
11:                  if $iou\_score>0.75$ then
12:                       $spatial\_relations.append((i,j,direction))$
13:                  end if
14:              end for
15:         end for
16:     end for
17:     return $spatial\_relations$
18:end procedure
Pre-conditions. We define the following 2 preconditions (and learn their grounding) in order to ensure that the generated plans are physically possible.
  1. 1.
    is-clear(blk, dir): We need to check whether a block blk has some free space in direction dir. For this we simply move the head in the direction dir with respect to the block blk. If the resulting position of head has 0 overlap with bounding boxes of all the other objects the predicate is True otherwise False.
  2. 2.
    will-not-be-floating(pred_loc): We need to check whether the resultant/predicted location of an object on taking an action would be dynamically stable or not. The location would be stable if either it is on top of some already placed object or it is on the table surface. The former can be checked through the resultant scene graph itself (that is obtained by applying the algorithm  2), while for the later we train an on-table classifier. This would take as input a bounding box and predict whether the box is on the table or not. For training this classifier we use the dataset of pretraining phase. The blockwise positive and negative sample annotation can be done automatically by giving GPT-4 the corresponding scene graph and then querying which objects are on the table and which aren’t.  17 gives an example. (Though we have not taken this approach for the complete dataset of 5k samples due to high cost).
Figure 17: Method to get annotation for training on-table classifier.
Actions. We define the following two types actions
  1. 1.
    place-random(blk): To place a block at a random free position on the table. For this we train a generative model (VAE) which learns the underlying distribution of bounding boxes for all the blocks that lie on the table. Given a scene we would sample position (bounding box) from this until we get a position that is not overlapping with the existing blocks in the scene. For training the VAE we assume that for every demonstration in the pretraining data, the first scene has all the objects randomly placed on the table (we could have also used the positive examples used for training on-table classifier).
  2. 2.
    move(rel, blk1, blk2): This action corresponds to moving the blk1 in the direction rel of blk2, resulting in the addition of a relation (rel, blk1, blk2) in the set of spatial relations. This action is defined as a sequential composition of the actions assign_head(blk2), move_head(rel), keep_at_head(blk1) (blk1 is a one hot tensor for the corresponding block).
Techniques and heuristic for efficient planning. Since the action space for the planner could be $o(n^{2})$, where n is the number of objects we adopt the following techniques to make planning scalable/efficient:
  1. 1.
    Heuristic - We define the heuristic value h(s) for a state s, as the number of relations that are present in the goal but are absent in the scene graph corresponding to the state s. Even though this heuristic is not admissible (as it may over overestimate the cost to goal), it was found to work optimally in most of the cases.
  2. 2.
    Greedy-pruning - We assume that all the actions resulting in states with higher or same heuristic value would be of the form place-random(blk). This means among the actions of the form move(rel, blk1, blk2) we only select those that lead to states with decreased heuristic value.
  3. 3.
    Relevant-object-set - Suppose O is the set of objects that are part of atleast one of the predicate in goal. We define O’ as the transitive closure of O with respect to the relation $Related$ in the initial state $s_{i}$, where $SG(s_{i})$ is scene graph for the initial state
    $Related(a,b,s_{i})\iff\exists dir((dir,a,b)\in SG(s_{i})\vee(dir,b,a)\in SG(s_{i}))\;$ (9)
    We assume O’ is the relevant set of object for completing the task and actions that move any other object should not be taken.
Appendix E Broader Impact
This work creates foundational knowledge in understanding human-like spatial abstractions. This work contributes towards the development of explainable and interpret-able learning architectures that may eventually contribute towards the development of embodied agents collaborating with and assisting humans in performing tasks. No negative impact of this work is envisioned.
Appendix F Details regardign Compute Requirements
All our experiments were run on a server with the following machine specifications.
CPU Specification:
Specification Value
Architecture x86_64
CPU op-mode(s) 32-bit, 64-bit
Address sizes 46 bits physical, 57 bits virtual
Byte Order Little Endian
CPU(s) 112
On-line CPU(s) list 0-111
Vendor ID GenuineIntel
Model name Intel(R) Xeon(R) Gold 6330 CPU @ 2.00GHz
CPU family 6
Model 106
Thread(s) per core 2
Core(s) per socket 28
Socket(s) 2
Stepping 6
CPU max MHz 3100.0000
CPU min MHz 800.0000
BogoMIPS 4000.00
GPU Specification:
Specification Value
GPU 1
Description VGA compatible controller
Product Integrated Matrox G200eW3 Graphics Controller
Vendor Matrox Electronics Systems Ltd.
Physical ID 0
Bus Info pci@$0$:03:00.0
Logical Name /dev/fb0
Version 04
Width 32 bits
Clock 66MHz
Capabilities pm vga_controller bus_master cap_list rom fb
Configuration depth=32 driver=mgag200 mingnt=16
Resources irq:16 memory:91000000-91ffffff memory:92808000-9280bfff
memory:92000000-927fffff memory:c0000-dffff
GPU 2
Description 3D controller
Product GA102GL [A40]
Vendor NVIDIA Corporation
Physical ID 0
Bus Info pci@$0$:17:00.0
Version a1
Width 64 bits
Clock 33MHz
Capabilities pm bus_master cap_list
Configuration driver=nvidia latency=0
Resources iomemory:21000-20fff iomemory:21200-211ff irq:18
memory:9c000000-9cffffff memory:210000000000-210fffffffff
memory:212000000000-212001ffffff memory:9d000000-9d7fffff
memory:211000000000-211fffffffff memory:212002000000-212041ffffff
GPU 3
Description 3D controller
Product GA102GL [A40]
Vendor NVIDIA Corporation
Physical ID 0
Bus Info pci@$0$:ca:00.0
Version a1
Width 64 bits
Clock 33MHz
Capabilities pm bus_master cap_list
Configuration driver=nvidia latency=0
Resources iomemory:28000-27fff iomemory:28200-281ff irq:18
memory:e7000000-e7ffffff memory:280000000000-280fffffffff
memory:282000000000-282001ffffff memory:e8000000-e87fffff
memory:281000000000-281fffffffff memory:282002000000-282041ffffff
Time Required: The time required for pretraining phase of all the neural modules is around 36 hours. For learning of inductive concepts the time taken varies from 5 minutes to 1 day depending on the search method used and the specific set of hyperparameters. However for our best approach we get the maximum performance in approx 12 minutes. Time taken for our approach during inference is less than 2 minutes per dataset.
Appendix G Hyperparameters and architecture details
G.1 Architecture for neural modules
Action Simulator:
1import torch.nn as nn
2
3class ActionSimulatorNetwork(nn.Module):
4 def __init__(self, bbox_mode, hidden_size = 256):
5 super(ActionSimulatorNetwork, self).__init__()
6 self.bbox_mode = bbox_mode
7 self.hidden_size = hidden_size
8
9 self.action_semantics_encoder = nn.Sequential(
10 nn.Linear(5, hidden_size),
11 nn.ReLU(),
12 nn.Linear(hidden_size, hidden_size),
13 nn.ReLU()
14 )
15 self.argument_encoder = nn.Sequential(
16 nn.Linear(5, hidden_size),
17 nn.ReLU(),
18 nn.Linear(hidden_size, hidden_size),
19 nn.ReLU()
20 )
21 self.decoder = nn.Sequential(
22 nn.Linear(hidden_size, hidden_size),
23 nn.ReLU(),
24 nn.Linear(hidden_size, 5),
25 nn.Tanh()
26 )
Listing 12: Action Simulator Network in PyTorch
Reactive Policy($\pi_{\text{neural}}$):
1import torch.nn as nn
2
3class NeuralSearch(nn.Module):
4 def __init__(self, action_space=6):
5 super(NeuralSearch, self).__init__()
6 self.action_space = action_space
7 self.fc1 = nn.Linear(10, 256)
8 # self.bn1 = nn.BatchNorm1d(256)
9 self.bn1 = nn.Identity()
10 self.fc2 = nn.Linear(256, 256)
11 # self.bn2 = nn.BatchNorm1d(256)
12 self.bn2 = nn.Identity()
13 self.fc3 = nn.Linear(256, 256)
14 # self.bn3 = nn.BatchNorm1d(256)
15 self.bn3 = nn.Identity()
16 self.fc4 = nn.Linear(256, action_space)
Listing 13: Neural Search in PyTorch
Random Position predictor (for grounding of place-random(blk)):
1import torch.nn as nn
2
3class VAE(nn.Module):
4 def __init__(self, input_dim, latent_dim):
5 super(VAE, self).__init__()
6 self.input_dim = input_dim
7 self.latent_dim = latent_dim
8 # Encoder
9 self.fc1 = nn.Linear(input_dim, 512)
10 self.bn1 = nn.BatchNorm1d(512)
11 self.fc2 = nn.Linear(512, 512)
12 self.bn2 = nn.BatchNorm1d(512)
13 self.fc3 = nn.Linear(512, 512)
14 self.bn3 = nn.BatchNorm1d(512)
15 self.fc4 = nn.Linear(512, 512)
16 self.bn4 = nn.BatchNorm1d(512)
17 self.fc51 = nn.Linear(512, latent_dim) # Mean of the latent space
18 self.fc52 = nn.Linear(512, latent_dim) # Log-variance of the latent space (log-var for numerical stability)
19
20 # Decoder
21 self.fc5 = nn.Linear(latent_dim, 512)
22 self.bn5 = nn.BatchNorm1d(512)
23 self.fc6 = nn.Linear(512, 512)
24 self.bn6 = nn.BatchNorm1d(512)
25 self.fc7 = nn.Linear(512, 512)
26 self.bn7 = nn.BatchNorm1d(512)
27 self.fc8 = nn.Linear(512, 512)
28 self.bn8 = nn.BatchNorm1d(512)
29 self.fc9 = nn.Linear(512, input_dim)
Listing 14: VAE in PyTorch
On-table classifier (for grounding of will-not-be-floating(pred_loc):
1import torch.nn as nn
2
3class TableClassifier(nn.Module):
4 def __init__(self):
5 super(TableClassifier, self).__init__()
6 self.fc1 = nn.Linear(5, 16)
7 self.bn1 = nn.BatchNorm1d(16)
8 self.fc2 = nn.Linear(16, 16)
9 self.bn2 = nn.BatchNorm1d(16)
10 self.fc3 = nn.Linear(16, 16)
11 self.bn3 = nn.BatchNorm1d(16)
12 self.fc4 = nn.Linear(16, 1)
13 self.bn4 = nn.BatchNorm1d(1)
14 self.sigmoid = nn.Sigmoid()
Listing 15: Table Classifier in PyTorch
G.2 Hyperparameters used in experiment
As indicated in  A.4 for the purpose of generalization through multiple candidate plans (from 1 demonstration) we chose the top-k plans (as measured by overall IoU achieved). The k chosen for all our experiments involving MCTS was 5. (The performance of our best approach was found to be the same for k=5 to 20). For every plan we obtain 3 programs from GPT-4 by re-prompting it 3 times with the same input prompt (with temperature > 0). From the pool of these 3*k programs we chose the one with highest IoU reward by running each of them on the given demonstration. The discount factor kept for our search is $\gamma$ = 0.95, and unless explicitly specified the number of expansions steps used = 5000.

### v3

@@
-Sketch-Plan-Generalize: Continual Few-Shot Learning of Inductively Generalizable Spatial Concepts
+Sketch-Plan-Generalize : Learning and Planning with Neuro-Symbolic Programmatic Representations for Inductive Spatial Concepts
 Abstract
-Our goal is to enable embodied agents to learn inductively generalize-able spatial concepts, e.g., learning staircase as an inductive composition of towers of increasing height.
-Given a human demonstration, we seek a learning architecture that infers a succinct program representation that explains the observed instance.
-Additionally, the approach should generalize inductively to novel structures of different size or complex structures expressed as a hierarchical composition of previously learned concepts.
-Existing approaches that use code generation capabilities of pre-trained large (visual) language models as well as purely neural models show poor generalization to a-priori unseen complex concepts.
-Our key insight is to factor inductive concept learning as: (i) Sketch: detecting and inferring a coarse signature of a new concept (ii) Plan: performing MCTS search over grounded action sequences (iii) Generalize: abstracting out grounded plans as inductive programs.
-Our pipeline facilitates generalization and modular re-use enabling continual concept learning.
-Our approach combines the benefits of code generation ability of large language models (LLM) along with grounded neural representations, resulting in neuro-symbolic programs that show stronger inductive generalization on the task of constructing complex structures in relation to LLM-only and neural only approaches.
-Furthermore, we demonstrate reasoning and planning capabilities with learned concepts for embodied instruction following.
+Effective human-robot collaboration requires the ability to learn personalized concepts from a limited number of demonstrations, while exhibiting inductive generalization, hierarchical composition, and adaptability to novel constraints.
+Existing approaches that use code generation capabilities of pre-trained large (vision) language models as well as purely neural models show poor generalization to a-priori unseen complex concepts.
+Neuro-symbolic methods (Grand et al., 2023) offer a promising alternative by searching in program space, but face challenges in large program spaces due to the inability to effectively guide the search using demonstrations.
+Our key insight is to factor inductive concept learning as: (i) Sketch: detecting and inferring a coarse signature of a new concept (ii) Plan: performing an MCTS search over grounded action sequences guided by human demonstrations (iii) Generalize: abstracting out grounded plans as inductive programs.
+Our pipeline facilitates generalization and modular re-use, enabling continual concept learning.
+Our approach combines the benefits of code generation ability of large language models (LLMs) along with grounded neural representations, resulting in neuro-symbolic programs that show stronger inductive generalization on the task of constructing complex structures vis-á-vis LLM-only and purely neural approaches.
+Further, we demonstrate reasoning and planning capabilities with learned concepts for embodied instruction following.
+Machine Learning, ICML
+[columns=3, title=Alphabetical Index, intoc]
 1 Introduction
-The ability to learn novel concepts which generalize inductively is one of the hallmarks of human intelligence [mind].
-Humans are highly data efficient – observing at a few instances of towers of a certain heights, we can generalize to constructing towers of any height.
-Further, we interpret increasingly complex concepts as hierarchical composition over simpler ones, e.g., a tower as a sequence of blocks placed on top of each other, or a staircase composed of towers of increasing height.
-This paper considers the problem of learning a program representation, from few demonstration, that models the inductive realization of grounded spatial concepts.
-Learning of such concepts is challenging task due to an expansive space of programs and the need to reason about their physical plausibility.
-Further, the representation must support inductive generalization over learned concepts as well as express complex hierarchical concepts via modular re-use of concepts learnt previously.
-Figure 1: Learning inductive spatial concepts.
+For robots to collaborate effectively with humans, they must quickly learn personalised concepts from just a few demonstrations.
+This ability to form inductive representations of novel grounded concepts and apply them for tasks beyond training is a hallmark of human intelligence (Tenenbaum et al., 2011; Chollet, 2019).
+In the context of robot learning, such representations are characterized by four fundamental properties: (a) it should be grounded directly in human demonstrations, reflecting the human’s intent rather than relying sorely on prior world knowledge, (b) it must generalize beyond observed examples; e.g., from a few towers of certain heights, infer how to build towers of arbitrary heights, (c) it should allow for the hierarchical interpretation of increasingly complex concepts as compositions of simpler ones; e.g., a staircase as a sequence of towers of increasing height, (d) it should be easily modifiable to incorporate additional constraints in new instructions, such as ensuring no green block is placed near a blue or yellow one in a staircase.
+Existing approaches fall short of fully satisfying these criteria.
+Recent advances in large language models (LLMs) have enabled general-purpose planners.
+These planners use world knowledge to identify relevant concepts and generate plans from language task specifications (Singh et al., 2022; Liang et al., 2023; Ahn et al., 2022; Achiam et al., 2023; Huang et al., 2022a).
+However, they often struggle to learn new concepts from demonstrations, especially when the concepts are linguistically novel.
+This is because they rely heavily on prior world knowledge, which limits their faithfulness to demonstrations and their ability to generalize inductively.
+Purely neural approaches, such as (Liu et al., 2023), can learn directly from demonstrations but typically fail to generalize beyond training data because they (a) do not explicitly model symbolic induction and (b) lack modularity, restricting their ability to reuse learned concepts.
+In contrast, neuro-symbolic approaches (Grand et al., 2023; Ellis et al., 2021) offer a more promising alternative.
+They leverage LLMs or enumerative program search to discover generalizable programs directly in the program (string) space.
+While these models outperform purely LLM-based or neural methods in terms of generalization, they still face significant challenges: enumerative searches become computationally infeasible in complex program spaces (e.g., Python programs), and such search cannot be easily guided by human demonstrations.
+Figure 1: Problem Overview.
+Our goal is to enable an embodied agent to learn grounded and generalizeable representations for spatial abstractions possessing a notion of induction (e.g., constructing a tower, row or their combinations such as staircases, boundary etc.).
 (Left) A human demonstrates the construction of a row and tower of size three.
 (Right) The agent learns program representation that enables inductive generalization to novel structures (varied sizes and visual attributes) and expresses complex concepts as hierarchical composition of previously acquired ones.
-E.g., learning a tower as a sequence of blocks placed one on top of another and a pyramid as rows of decreasing size.
-Prior efforts such as [code-as-policies] uses a LLM (Large Language Model) to generate control program for a given task specification but fail to generalize to complex spatial concepts which are difficult to tokenize.
-Extension of this to VLMs (Vision Language Models), [achiam2023gpt], also fail to generalize when presented with linguistically novel concepts, due to over reliance on prior knowledge and their inability to effectively learn novel concepts from given demonstrations.
-On the other hand, neural approaches such as  [struct-diffusion], learn from given demonstrations but generalize poorly due to their inability to (a) explicitly model symbolic concept of induction and (b) modularize as well as re-use previously acquired concepts.
-Approaches such as [li2019practical] train RL-based policies to attain spatial assembly by encoding an inductive spatial prior using GNNs within the policy architecture.
-However, generalization is still limited and assumes an elicitation of the goal as per object positions, resulting in lack of ability to take in goal description such as “construct a tower of size three”.
-In summary, we attribute poor generalization of such approaches to an implicit entangling of the following objectives: (i) postulating a high-level program for new concepts, (ii) evaluating plausibility of grounded plans to align with human demonstration of concepts to be learned and (iii) abstracting out a grounded program to facilitate inductive generalization and modular re-use in a continual manner.
-In response, this paper introduces an approach that factorizes the concept learning task as: (a) Sketch: Given a language-annotated demonstration of a novel concept, an LLM is used to postulate a function signature.
-(b) Plan: Refinement of program sketches via MCTS search, rapidly evaluating actions sequences guided by a reward associated with constructing a concept.
-The search is accelerated by training a neural action predictor that uses the given demonstrations.
-(c) Generalize: Leveraging the code generation capability of an LLM to distill grounded plans into a program that is inductively generalize-able.
-This results in a continually evolving library of concepts which can be used to hierarchically learn complex concepts in future.
-The modular architecture enables continual learning by providing the ability to decide whether the new concept encountered either as a symbolic composition of existing concepts, or, a neural embedding trained via gradient update.
-Our experimental evaluation demonstrates accurate learning of simple and complex concepts from few demonstrations for a range of spatial structures.
-Further, our approach also shows inductive generalization in out-of-distribution settings, significantly over the baselines.
-We also present deeper insights related to efficiency gains obtained by combining symbolic MCTS with neural action predictor.
-Furthermore, we show how learned concepts can be grounded in the visual input, enabling a robot to follow natural language instructions referring to a-priori unseen spatial configurations.
+E.g., tower as blocks placed one on top of another and a pyramid as rows of decreasing size.
+In response, we present a neuro-symbolic agentic framework that (i) identifies when learning a new concept is required, (ii) requests human demonstrations to learn an inductively generalizable representation of the unknown concept, and (iii) uses the learned concepts to generate control policies for novel instruction-following tasks with discrete constraints.
+Unlike prior methods, our concept learning approach, called SPG, factorizes the concept learning task into three stages: (a) Sketch: Given a language-annotated demonstration of a novel concept, an LLM is used to postulate a function signature.
+(b) Plan: Using MCTS, grounded action sequences are discovered by maximizing a reward that measures how closely the constructed structure matches the human demonstration.
+(c) Generalize: The code-generation capability of an LLM is leveraged to distill grounded plans from multiple demonstrations into an inductively generalizable program.
+This enables learning of inductively generalizable representations while restricting the search to the grounded actions, which can be effectively guided through dense rewards from human demonstration and is less expensive than searching directly in the program space.
+We also show how the learned concepts can be used for goal-conditioned planning in novel input scenes and instructions, even when visual or spatial constraints are present.
+Our experiments demonstrate that our framework can accurately learn both simple and complex concepts from just a few demonstrations across various spatial structures.
+Moreover, it exhibits strong inductive generalization in out-of-distribution settings, significantly outperforming baseline methods.
 2 Related Works
-The ability to ground symbolic knowledge is central to building intelligent embodied agents.
-The symbol grounding problem is often cast as learning an association between linguistic description of a concept and aspects of the state-action space such as objects [paul2018efficient, ramp], spatial relations  [tellex2011approaching, kim2024lingo], reward functions  [boularias2015grounding], or motion constraints  [howard2014natural].
+Concept Learning: The problem of acquiring higher-order programmatic constructs is often modelled as Bayesian inference over a latent symbol space given observed instances.
+Seminal works have demonstrated efficient inference over latent generative programs to express handwritten digits (Lake et al., 2015), 2D drawings (Ellis et al., 2018), grid-world puzzles (Wang et al., 2024a), visual question answering (Mao et al., 2019), goal-directed policies (Silver et al., 2019), and compressed or refactored code (Grand et al., 2023; Ellis et al., 2021).
+These approaches typically either (i) search program space using a probabilistic grammar, which is infeasible for large program spaces like Python, or (ii) use LLMs to directly infer programs, which may produce physically ungrounded outputs.
+Furthermore, because partially generated programs cannot be evaluated, direct search in program space does not effectively leverage demonstrations to guide the search.
+In contrast, our approach first searches for grounded action sequences that can be efficiently guided by dense rewards from human demonstrations.
+We then use an LLM to abstract these sequences into a program, making program inference more tractable and efficient.
+While prior works focus on 2D problems without considering physical plausibility (for example, towers can be drawn from top to bottom but cannot be constructed that way in reality), we expose the search to assessing the physical construction plausibility, thereby learning physically grounded concepts.
+Learning-to-plan Methods: Our work is complementary to efforts that learn symbolic constructs for efficient planning.
+Works such as (Silver et al., 2023, 2024; Liu et al., 2024), infer state-action abstractions for planning by querying large pre-trained models or by optimizing a goal attainability objective.
+This paper, instead, focuses on learning a representation for complex spatial assemblies as inductive programs leading to the ability to infer complex goal specifications which can then be combined with aforementioned works for synthesize efficient plans to realize complex assemblies.
+Works such as (Li et al., 2020) learn to construct structures by encoding relational knowledge via graph neural network.
+However, this effort suffers from poor generalization to unseen examples, (e.g., tower of larger size) and do not possess a mechanism to re-use previously acquired concepts.
+Works such as (Wang et al., 2023a, 2024b) shows lifelong learning of skills by learning to plan high-level tasks through composition of simple skills for simulated agents.
+Others (Wan et al., 2023; Parakh et al., 2023) initiate new skill acquisition upon detecting task failure, building a library of skills over time.
+However, they do not model inductive use of learned concepts and initiate skill acquisition only upon failure as opposed to learning continually even from goal-reaching demonstrations.
+Robot Instruction Following: Instruction following involves grounding symbolic constructs expressed in language with aspects of the state-action space such as object assemblies (Paul et al., 2018; Collins et al., 2024; Lachmy et al., 2022), spatial relations  (Tellex et al., 2011; Kim et al., 2024), reward functions  (Boularias et al., 2015), or motion constraints  (Howard et al., 2014).
 These works assume the presence of grounded representation for symbolic concepts and only learn associations between language and concepts.
 In contrast our work jointly learns higher-order concepts composed of simpler concepts along with their grounding in the robot’s state and action space.
-Another line of work  [singh2022progprompt, wang2023demo2code, ahn2022can, code-as-policies] leverages prior-knowledge embodied in LLMs to determine grounded robot control programs for a high-level tasks, given a tokenized description of the scene and robot capabilities.
-Our experiments (reported subsequently) demonstrate that complex spatial relations are challenging for textual descriptions to encode and are difficult to reason over.
-Hence, we adopt a factored approach that uses a search procedure guided by visual reward, acting as a medium to connect the task sketch generation and program abstraction stages (performed by an LLM).
-Works such as [wang2023voyager, jarvis-1] shows lifelong learning of skills by learning to plan high-level tasks through composition of simple skills for simulated agents.
-Efforts such as  [lotus],  [lifelong-halp] detect when control primitives for a robot are insufficient to perform a task and initiate skill learning, building a library over time.
-However, such efforts do not model deep inductive use of learned concepts.
-Finally, this work draws inspiration from concept learning approaches in the broader AI community.
-Related works infer latent generative programs to express hand-written digits [lake2015human], object arrangements [ellis2018learning], motion plans [nscl] or goal-directed policies [lpp].
-These approaches do not explicitly model inductive concept of modular re-use of previously acquired concepts.
-Other efforts such as [li2019practical] learn RL-based policies to construct object stacks.
-Here, the neural policy models locality of object influence via GNNs.
-However, the model suffers from poor generalization to unseen examples, (eg.
-tower of larger size) and do not possess a mechanism to re-use previously acquired concepts.
-3 Background and Problem Formulation
+Others (Singh et al., 2022; Wang et al., 2023b; Ahn et al., 2022; Liang et al., 2023) leverage prior-knowledge embodied in large vision-language models to directly translate high-level tasks to robot control programs.
+Our experiments (reported subsequently) demonstrate their limitation in outputting programs for structure assembly-type tasks that require long-range (inductive) spatial reasoning and consideration of physical plausability of construction.
+Our approach addresses this problem by coupling abstract task knowledge from pre-trained models with physical reasoning in the space of executable plans.
+3 Preliminaries and Problem Setting
 We consider an embodied agent that uses a visual and depth sensor to observe its environment and can grasp and release objects at specified poses.
@@
 We assume that the agent possesses a model of semantic relations (e.g., left(), right() etc.) as well as semantic actions such as moving an object by grasping and releasing at a target location.
-Such modular and composable notions can be acquired from demonstrations via approaches outlined in [nsrm-Kalithasan2022LearningNP, nscl, mao2022pdsketch].
+Such modular and composable notions can be acquired from demonstrations via approaches outlined in (Kalithasan et al., 2023; Mao et al., 2019, 2022).
 Such notions populate a library of concepts $\mathcal{L}$ available as grounded executable function calls.
-Following recent efforts [code-as-policies, Huang2022InnerME, ahn2022can, singh2022progprompt] in representing robot control directly as executable programs, we represent action sequence corresponding to a plan as a program consisting of function calls to executable actions and grounded spatial reasoning.
-Our goal is to enable a robot to interpret instructions such as “construct a tower with red blocks of height five".
-Specifically, we aim to learn spatial constructs like a tower that requires sequential actions that repeatedly places a block on top of a previously constructed assembly; a process akin to induction.
-Given a few demonstrations of a constructing a spatial assembly, $\mathcal{D}$, each consisting of natural language description $\Lambda$ (“construct a tower of red blocks of size five") and a sequence of key frame states $\{S_{1},\cdots,S_{g}\}$ associated with the construction process, we seek to learn a program that models the inductive nature of the concept of tower.
-Online, the agent must generalize inductively to following an instruction as “construct a tower of blue blocks of size ten".
-Once the concept of tower is acquired, it is added to the concept library.
-This aids in learning the notion of a staircase as sequence of towers of increasing height.
+Following recent efforts (Liang et al., 2023; Huang et al., 2022b; Ahn et al., 2022; Singh et al., 2022) in representing robot control directly as executable programs, we represent action sequence corresponding to a plan as a program consisting of function calls to executable actions and grounded spatial reasoning.
+Our goal is to enable a robot to interpret and learn the concepts in instructions such as “construct a tower with red blocks of height five".
+Specifically, we aim to learn spatial constructs like a tower that requires sequential actions that repeatedly place a block on top of a previously constructed assembly, a process akin to induction.
+Given a few demonstrations of constructing a spatial assembly, $\mathcal{D}$, each consisting of natural language description $\Lambda$ (“construct a tower of red blocks of size five") and a sequence of key frame states $\{S_{1},\cdots,S_{g}\}$ associated with the construction process, we seek to learn a program that models the inductive nature of the concept of tower.
+This learned representation should enable the agent to generalize inductively to new instructions, such as "construct a tower of blue blocks of height ten." Moreover, the learned concepts should facilitate the learning of more complex structures, which are challenging to represent using primitive actions alone.
+For example, the concept of a "tower" should assist in learning a "staircase," which can be represented as a sequence of towers of increasing heights.
+4 Representing Inductive Spatial Concepts
 We formalize the notion of inductive spatial concepts and formulate the learning objective.
-Let $C_{1},\cdots,C_{|\mathcal{L}|}$ represent the topological order of concepts present in the concept library $\mathcal{L}$ as per their structural complexity.
-An inductive spatial concept $C_{k}$ of size $n$ at position $p$ can be expressed as:
-$h(C_{k},n,p)=\underbrace{h^{\lambda}(C_{k},n-1,pos(.))}_{\text{Induction (I)}}\circ\underbrace{\prod\limits_{l=1}^{L(C_{k})}h\big{(}C_{k^{\prime}_{l}},\mathtt{size}(.),pos(.)\big{)}}_{\text{Composition (C)}}\circ\underbrace{\prod\limits_{l=1}^{L^{\prime}(C_{k})}\eta_{\theta}^{l}\big{(}pos(.)\big{)}}_{\text{Base (B)}},$ (1)
-where, $\lambda\in\{0,1\}$, $k^{\prime}<k$, $0\leq L(C_{k}),L^{\prime}(C_{k})\leq o(|\mathcal{L}|)$ and $\eta_{\theta}^{l}$ is the set of primitive actions resulting in placement of object at a desired spatial position.
-Here, $pos(.)=pos(C_{k},l,n,p)$ and $\mathtt{size}(.)=\mathtt{size}(C_{k},l,n)$ are functions that predict the size and position of the structure to be constructed.
-For example, constructing a tower of size $n$ at position $p$ i.e., $h(Tower,n,p)=h^{\lambda=1}(Tower,n-1,p)\,\circ\,OnTop_{\theta}(p+(n-1)d)$.
-Here, $\lambda=1$, $L(C_{k})=0$, $L^{\prime}(C_{k})=1$, $\eta^{1}=OnTop$, $pos(Tower,l=1,n,p)=p+(n-1)\cdot d$ and $d$ is the dimension of the object.
-Thus, a tower can be described using induction and a single primitive base function with zero compositions.
-On the other hand, enclosures can be described as a composition of two rows and two columns without induction or base components.
-The functional space of inductive concepts ($h$) leads to a hypothesis space $\mathcal{H}$ of associated neuro-symbolic programs.
+Inductive Spatial Concepts: A spatial structure is an inductive concept if its construction can be described recursively using a similar structure of smaller size or as a composition of other simpler structures.
+Formally, let $C_{1},\cdots,C_{|\mathcal{L}|}$ represent the concepts in the concept library $\mathcal{L}$.
+We define a partial order on $\mathcal{L}$ where a concept $C$ is "dependent on" $\widetilde{C}$ if $\widetilde{C}$ is a substructure of $C$.
+For example, a staircase is dependent on a tower, and $X$ (cross) is dependent on diagonals, and so on.
+This partial order is referred to as structural complexity, where a concept $C$ is more structurally complex than $\widetilde{C}$ if $C$ is dependent on $\widetilde{C}$.
+Without loss of generality, assume that $C_{1},\cdots,C_{|\mathcal{L}|}$ are written in topological order as per their structural complexity.
+Now, the construction of an inductive spatial concept $C_{k}$ of size $n$ at position $p$, denoted by $h(C_{k},n,p)$, is defined recursively as:
+$h(C_{k},n,p)=\underbrace{h^{\lambda}(C_{k},n-1,pos(.))}_{\text{Induction (I)}}\circ\\ \underbrace{\prod\limits_{l=1}^{L(C_{k})}h\big{(}C_{k^{\prime}_{l}},\mathtt{size}(.),pos(.)\big{)}}_{\text{Composition (C)}}\circ\underbrace{\prod\limits_{l=1}^{L^{\prime}(C_{k})}\eta_{\theta}^{l}\big{(}pos(.)\big{)}}_{\text{Base (B)}}$ (1)
+where, $\lambda\in\{0,1\}$, $k^{\prime}<k$, $0\leq L(C_{k}),L^{\prime}(C_{k})\leq o(|\mathcal{L}|),pos(.)=pos(C_{k},l,n,p)$ and $\mathtt{size}(.)=\mathtt{size}(C_{k},l,n)$ are functions that predict the size and position of the structure to be constructed.
+  1. 1.
+    Induction term: The first expression (I) represents induction vis-á-vis the possibility of constructing $C_{k}$ of size $n$ using $C_{k}$ of size $n-1$.
+Here, $\lambda$ depicts the absence($\lambda=0$) or presence($\lambda=1$) of the (I) term.
+  2. 2.
+    Composition term: The second term (C) expresses the construction of $C_{k}$ as a composition of prev.
+known concepts in the library ($\mathcal{L}$).
+The number of required compositions ($L(C_{k})$) depends on $C_{k}$ and $\mathtt{len}(\mathcal{L})$.
+  3. 3.
+    Base term: The third term (B) defines the base case where the construction of concept $C_{k}$ may include $L^{\prime}$ number of primitive actions.
+Here, $\eta_{\theta}$ are pre-trained primitive actions resulting in the repositioning of an object to the desired spatial position.
+For ex., a tower of size $n$ can be constructed by a (primitive) move-top action after constructing a tower of size $n-1$.
+Learning Objective: The functional space of inductive concepts ($h$) leads to a hypothesis space $\mathcal{H}$ of associated neuro-symbolic programs.
 Each goal-reaching demonstration corresponds to a particular instantiation of a given inductive concept, i.e.
 $h(C_{k},n,p)$, where the $p$ comes from the sequence of frames, and $n$, $C_{k}$ comes from $\Lambda$.
-Given demonstration, we aim to learn a generic representation for the given concept, which is general for all n, p i.e.
-$h(C_{k},\cdot,\cdot)$.
-We model this as learning a particular $H=h(C_{k},\cdot,\cdot)$ where $H\in\mathcal{H}$.
-Given (few) demonstrations of a human constructing a spatial structure, concept learning can be formulated as the Bayesian posterior [lake2015human, shah2018bayesian, lpp], $\mathtt{P}_{\mathcal{H}}(H|\Lambda,S_{1}..S_{g})\propto\mathtt{P}(S_{1}..S_{g}|\Lambda,H)\cdot\mathtt{P}(H|\Lambda)$.
+We aim to learn a generic representation $H=h(C_{k},\cdot,\cdot)\in\mathcal{H}$ for the given concept, which is general for all $n$ and $p$.
+Given (few) demonstrations of a human constructing a spatial structure, concept learning can be formulated as the Bayesian posterior (Lake et al., 2015; Shah et al., 2018; Silver et al., 2019), $\mathtt{P}_{\mathcal{H}}(H|\Lambda,S_{1}..S_{g})\propto\mathtt{P}(S_{1}..S_{g}|\Lambda,H)\cdot\mathtt{P}(H|\Lambda)$.
 Here, the likelihood term associates a candidate program, and the prior term regularizes the program space.
 The maximum a-posteriori estimate, representing the learnt program, is obtained by optimizing the following objective:
-$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\left(\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))\right)-\log\left(\mathtt{P}(H|\Lambda)\right)\right]$ (2)
+$H^{*}=\arg\underset{H\in\mathcal{H}}{\mathrm{min}}[\mathtt{Loss}(\{S_{1}..S_{g}\},Exec(H,\Lambda,S_{1}))\\ -\log\mathtt{P}(H|\Lambda)]$ (2)
 Since exact inference is intractable, approximate inference is performed via search in the program space.
@@
 Additionally, we seek strong generalization from a few instances of an inductive structure to structures with arbitrary sizes, in effect favouring programs with iterative looping constructs.
-4 Technical Approach: Learning Inductive Spatial Concepts
-Figure 2: Approach Overview.
+5 Learning Inductive Concepts from Demonstrations
+Figure 2: Method Overview.
 We learn a neuro-symbolic program for inductive spatial concepts factored as (a) Sketch (b) Plan (c) Generalize.
-The example above shows the progressive realization of a program for the concept of a tower acquired by observing a single demonstration of building a tower of size three and its corresponding natural language instruction, “construct a tower of height three using green die."
-We address the problem of estimating a succinct generalized program, Eq.
-2, modeling an inductive spatial concept modeling structures whose construction is observed in a human demonstration.
-Direct symbolic search in the space of programs is intractable (particularly due to looping constructs needed for modeling induction) but can explicitly reason over previously acquired concepts.
-Alternatively, neural methods attempting to predict action sequence to attain the assembly are challenged by continual setting where concepts can increase over time building on previously learnt ones but are resilient to noise.
+The example above shows the progressive realization of a program for the concept of a staircase acquired by observing a single human demonstration of building a staircase of size four, and its corresponding language instruction
+For estimating a succinct generalized program as per Eq.
+2, a direct symbolic search in the space of programs can explicitly reason over previously acquired concepts, it is intractable, particularly due to looping constructs needed for modeling induction.
+Alternatively, neural methods that attempt to directly predict action sequence to explain the demonstrations are resilient to noise, but are challenged in the continual setting as the number of concepts can increase prohibitively over time.
 Our approach blends both approaches and factors the concept learning task as:
   • Sketch: From the natural language instruction ($\Lambda$), we extract a task sketch ($H_{S}^{*}$) using an LLM that provides the signature (concept name and instantiated arguments) of the concept to be learned.
-When grounded in the initial scene of the demonstration, the task sketch provides a particular instance of the concept demonstrated in the given demonstration.
-  • Plan: MCTS-based search using the already learnt concepts that outputs the sequence of grounded actions, best explaining the given demonstration.
+  • Plan: An MCTS-based search is performed using the library concepts (learnt previously) which outputs a sequence of grounded actions that best explain the given demonstration.
   • Generalize: The grounded plan $H_{P}^{*}$ and task sketch $H_{S}^{*}$ are provided to an LLM to obtain a general Python program whose execution on the given scene matches the searched plan.
-The progress exploration of the program space can be summarised as follows:
-$H_{S}^{*}\leftarrow Sketch(\Lambda\,;\,{\theta_{S}})\,;\,H_{P}^{*}\leftarrow Plan(S_{1}..S_{g},H_{S}^{*};\,{\theta_{P}})\,;\,H_{G}^{*}\leftarrow Generalize(H_{P}^{*},\,H_{S}^{*};\,{\theta_{G}})$ (3)
+Formally, the factored exploration of the program space for a demonstration is performed as:
+$H_{S}^{*}\leftarrow Sketch(\Lambda\,;\,{\theta_{S}})\,;\,H_{P}^{*}\leftarrow Plan(S_{1}..S_{g},H_{S}^{*};\,{\theta_{P}})\,;\\ \,H_{G}^{*}\leftarrow Generalize(H_{P}^{*},\,H_{S}^{*};\,{\theta_{G}})$ (3)
+Here, $\theta_{S}$, $\theta_{P}$ and $\theta_{G}$ are the learnable parameters (including hyperparameters) of the Sketch, Plan and Generalize functions, respectively.
 The concept library $\mathcal{L}$ is initialized with primitive visual and action concepts.
 Upon acquiring a new inductive concept $H_{G}^{*}=H^{*}$, we update our library accordingly: $\mathcal{L}\leftarrow\mathcal{L}\cup{H^{*}}$.
-Next, we detail each of the three steps mentioned above.
+An example is provided in Appendix Sec.
+A.5.
 Fig.
-2 illustrates an example of progressive program estimation.
+2 illustrates an example of progressive prog.
+estimation.
+Next, we detail each of these three steps.
 5.1 (Sketch) Grounded Task Sketch Generation
@@
 The task sketch is a tree of nested function calls that outlines the function header (name and the parameters) of the inductive concept/program to be learned.
-A detailed exposition on prompting appears in the Appendix  C.1.
-The task sketch thus obtained is then grounded on the input scene using a quasi-symbolic visual grounding module akin to [nscl, nsrm-Kalithasan2022LearningNP, wang2023progport].
-This module has three key components: (1) a visual extractor (ResNet-34 based) that extracts the features of all objects in the scene, (2) a concept embedding module that learns disentangled representations for visual concepts like green and dice, and (3) a quasi-symbolic executor equipped with pre-defined behaviours such as “filter” to select/ground the objects of interest.
-For example, grounding the task sketch “Tower (height =3, objects = filter(green, dice))” results in an instantiated function call “Tower(height = 3, objects = [1, 2, 3])” where [1, 2, 3] are the green coloured dice.
+Detailed exposition on prompting appears in the app.
+C.1.
+The task sketch thus obtained is then grounded on the input scene using a visual grounding module akin to (Mao et al., 2019; Kalithasan et al., 2023; Wang et al., 2023c).
+This module has three key components: (1) a visual extractor (ResNet-34 based) that extracts the features of all objects in the scene, (2) a concept embedding module that learns disentangled representations for visual concepts like green and dice, and (3) an executor equipped with pre-defined behaviours such as “filter” to select/ground the objects of interest.
+For example, grounding the task sketch “Tower (height =3, objects = filter(green, dice))” results in an instantiated function call “Tower(height = 3, ObjectSet)” where ObjectList is an ordered list of the green coloured dice.
+Since it is now grounded in the initial scene of the demonstration, the task sketch corresponds to a particular instance of the concept demonstrated in the given demonstration.
 5.2 (Plan) Physical Reward Guided Plan Search
@@
 Specifically, this involves determining the concepts, their respective grounded parameters, and the order of composition as specified in the Equation 1.
-Identifying the position $pos_{\theta}(.)$ of the object placement/structure construction is achieved through a notion of head.
-Essentially, the head represents a cuboidal enclosure in 3D space.
-The movement of any object involves two key stages: (1) relocating the abstract head to the desired position and (2) picking the target object and placing it at the current position of the head.
-Conceptually, the movement of the head can be likened to the robot’s cognitive exploration of potential object placements to attain the desired spatial configuration.
-We define a set of four primitive functions, $\mathcal{A}_{p}$, that manipulate the head.
-(1) store_head(): remembers the current position of the head in a stack.
-(2) reset_head(): sets the head to the last remembered position (i.e., to the top of the stack) and pops this position from the stack (3) keep_at_head(objects): places the object corresponding to the index with highest value in objects (a list of values) at the current location of head and (4) move_head(direction): A neural operator that takes the head and predicts the new location of the head with respect to the given direction.
-The move_head function is trained using a corpus of pick and place instructions, such as “move the green object to the right of the red cube” similar to the approach in [nsrm-Kalithasan2022LearningNP].
-When an object is passed as the parameter, the move_head operator is overloaded to assign the head to the location of the object (also referred to as assign_head(object)).
-The state is defined by bounding boxes (including the depth of the center) and visual attributes of all the objects that are present on the table.
-We assume that the head is initially assigned to the position where the structure needs to be constructed.
-For each learned inductive concept <cpt>, we define a macro-action “Make_<cpt>(size)" that executes the corresponding program with the given size argument, resulting in the construction of the desired inductive concept.
-Thus, the action space $\mathcal{A}$ is the union of primitive actions $\mathcal{A}_{p}$ and compound/macro-actions $\mathcal{A}_{c}$.
-All actions except keep_at_head(objects) and macro actions yield a reward of zero, while for the latter, the Intersection over Union (IoU) between the attained state and the expected state in the demonstration is provided as a reward.
-An MCTS procedure is performed to find a plan that maximizes the reward.
-The search outputs a sequence of grounded actions for an instantiation of the given inductive concept by the task arguments.
-MCTS that searches for a plan only in terms of primitive actions may not be generalizable due to lack of modularity  C.3.
+Primitive Actions.
+Constructing complex structures involves two steps: (1) identifying or imagining the placement location of an object/structure and (2) picking and placing the object at the imagined location.
+The position $pos_{\theta}(.)$ for placement is determined using a head, which represents a cuboidal enclosure in 3D space.
+Conceptually, moving the head is akin to the robot’s cognitive exploration of potential placements to achieve the desired spatial configuration.
+We define a set of primitive functions, $\mathcal{A}_{p}$, to guide the movement and placement of objects in two ways: (1) move_head(direction): This primitive is a neural operator that moves the abstract head to a desired relative position.
+It is trained on a corpus of pick-and-place instructions, similar to the approach in (Kalithasan et al., 2023); and (2) keep_at_head(objects): This primitive places the target object from the list objects at the current location of the head.
+MCTS Search.
+We use an object-centric state representation defined by bounding boxes (including the depth of the center) and visual attributes of all the objects that are present on the table.
+Each previously learned inductive concept <cpt> has an associated macro-action Make_<cpt>(size), that executes the corresponding program for the given size argument, resulting in the construction of the desired concept.
+Thus, the set of actions $\mathcal{A}$ is the union of primitive actions $\mathcal{A}_{p}$ and compound/macro-actions $\mathcal{A}_{c}$.
+Intersection over Union (IoU) between the attained state and the expected state in the demonstration is provided as a reward for all macro actions and keep_at_head(objects); all other actions yield zero reward.
+An MCTS procedure similar to (Khandelwal et al., 2016) (detailed in Appendix A.3) is performed to find a plan (sequence of grounded actions) that maximizes the reward.
+Modularity and Scalability.
+MCTS that searches for a plan only in terms of primitive actions may not be generalizable due to lack of modularity( C.3).
 The use of macro-actions in the search ensures that the plan $H_{p}^{*}$ for a given demonstration is concise, modular, and easily generalizable.
-This can be seen as a form of regularization in terms of the length of concept description by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2) $H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{n}\},H(\Lambda))+\alpha\log|H|\right]$.
+This can be seen as a form of regularization in terms of the length of concept description by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2)
+$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{n}\},H(\Lambda))+\alpha\log|H|\right]$
 However, as the action space expands with the learning of more concepts, the search becomes slower, necessitating the pruning of the search space.
@@
 Additionally, to prune primitive actions, we train a reactive policy $\pi_{neural}$ which, given the current state $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (from the demonstration), outputs one of the primitive actions $a_{t}^{*}\in\mathcal{A}_{p}$.
-Consequently, the effective branching factor of the search is reduced to $|\mathcal{A}_{c}|+1$.
+Consequently, the effective branching factor of the search is reduced from $N*|\mathcal{A}_{c}|+|\mathcal{A}_{p}|$ to $|\mathcal{A}_{c}|+1$, where $N$ is the average number of size-parameter variants.
 Thus, our MCTS algorithm is modular through the hierarchical composition of learned concepts and efficient through action space pruning, and is referred to as MCTS+L+P.
-Further details regarding modifications in MCTS are given in the Appendix  A.4.
+(Detailed in app.
+A.3)
 5.3 (Generalize) Plan to Program Abstraction
-Leveraging the code generation and pattern matching abilities of LLMs [pattern] we use GPT-4 to distil out a general Python program from the sequence of grounded actions as determined by MCTS+L+P.
+Leveraging the code generation and pattern matching abilities of LLMs (Mirchandani et al., 2023), we use GPT-4 to distil out a general Python program from the sequence of grounded actions as determined by MCTS+L+P.
 The learnt program is incorporated in the concept library, $\mathcal{L}$, for modular reuse in subsequent learning tasks.
-Additional details, prompting mechanism and use of learnt programs in the search step of future learning tasks are described in Appendix  C.2,  A.4.
-We take a curriculum learning approach beginning from learning of primitive actions and visual attributes, followed by structures of increasing complexity.
-Additional details about learning of inductive concepts via curriculum, architecture details, and learning from multiple demonstrations are in Appendix  C.2, Fig.
-7,  A.5.
+We take a curriculum learning approach (App.
+D.2).
+Further details & discussion can be found in app.
+items C.2A.4 and  A.3.
+6 Adapting Learned Concepts for Novel Instructions and Constraints
+Goal Grounding: A complex Natural Language (NL) instruction is passed to the Parser, which extracts two components: the relevant concepts and the associated constraints.
+If some of the concept(s) are not present in the concept library $\mathcal{L}$, the system queries for a Human Demonstration, which is then processed via the SPG framework, and the abstracted & generalised concept(s) are appended to $\mathcal{L}$, as in Figure: 3.
+Figure 3: Pipeline for using Solvers and LLMs to solve constraints in Instructions over concepts in Library $\mathcal{L}$ and learning additional concepts via SPG if necessary.
+Following this, an LLM identifies whether an explicit constraint resolution is required.
+If yes, the Language Parser isolates the constraint components from the NL Instruction.
+, and generates a program of constraints that can be solved independently.
+This bifurcation allows the system to resolve increasingly complex constraints without entangling them with the core structural learning.
+This is an improvement over directly relying on LLMs to simultaneously learn structures and resolve complex constraints, since LLMs often fail to reason about spatial grounding, visibility, occlusion, and physical stability (Appendix: D.5).
+Our pipeline modularizes the constraint resolution stage by invoking neural solver modules, external constraint solvers, or LLMs, based on downstream tasks, thereby improving the pure-LLM baseline.
+We aim to improve the robustness of our pipeline by integrating advanced reasoning capabilities.
+Plan Synthesis: We qualitatively demonstrate that the concepts we have acquired can help us to perform goal conditioned planning.
+These goals may either be direct instances of concepts in $\mathcal{L}$, or constraint-satisfying goals obtained from sec.
+6 above.
+We perform a forward search using the abstract state and action representations to the inferred goal akin to  (Liu et al., 2024; Kalithasan et al., 2024) (detailed in Appendix D.5).
 7 Evaluation Setup
-Data sets.
-Evaluation data set was created with a simulated Franka Emika Robot Manipulator constructing arrangement of blocks of uniform size with varied type/appearance (e.g., cubes, dice, lego etc.).
-The scene is observed with a visual and depth sensor with object masks extracted from a standard segmentation procedure.
-A total of $15$ structure types are considered including towers, staircases, X (cross), arch-bridge.
-The state and action sequences for attaining these structures forms the demonstration data fro training, details in Appendix  10.
-Figure 3: Two illustrative examples of complex spatial structures from our dataset with a hierarchical composition of inductively defined structures.
-The training data for each concept consists of $3$ demonstration per structure, of size (num.
-of inductive steps) $3-5$ with up to $20$ objects.
-We form three evaluation data sets, each consisting of simple structures (different size instances of a concept) and complex structures composed of simpler concepts (e.g., staircase consists of towers as substructure).
-Dataset I includes contains structures ranging in sizes $3-5$.
-Dataset II reverses the linguistic labels (e.g., the “tower" becomes “rewot").
-Dataset III includes concepts of larger sizes ($4-7$).
-Alternative approaches. We compare our model with four baselines spanning two approaches:
-(1) Purely-Neural: An end-to-end neural model inspired by StructDiffusion, [struct-diffusion] that treats structure construction as a rearrangement problem.
-We consider two variations of the model 1.1) Struct-Diff (SD): End-to-end approach without any additional supervision regarding which objects need to be moved.
-1.2) With-Grounder (SD+G): Similar to 1.1) except that we assume a perfect object selector/grounder that identifies the relevant set of objects which are to be moved.
-(2) Pre-trained models that directly output symbolic programs: 2.1) LLM for Scene-Graph Reasoning: pre-trained language models (GPT-4) that generate Python programs from instructions which would describe the given demonstration.
-To help LLM understand the underlying structure, symbolic spatial relationships (e.g., left(a,b)) between objects in the demonstration are also provided to the LLM.
-For this baseline we further assume absence of distractor objects in the scene.
-2.2) Vision Language Model (GPT-4-V): Similar to 2.1 but has the ability to take input demonstration as images.
-For learning program of a new inductive concept we give the demonstration to the VLM in the form of $\Lambda,(S_{1}..S_{g})$.
-Additional details on prompting method in Appendix C.6111Experiments were also conducted with open source code-generation LLM such as CodeLlama (70Bq), Due to significantly poorer performance w.r.t.
-GPT-4, it was removed as the primary LLM baseline..
+Corpus.
+A corpus is created using a simulated Robot Manipulator assembling spatial structures on a table-top, viewed by a visual-depth sensor.
+Demonstration data ($3$ demonstrations per structure, up to $20$ objects per scene) includes RGBD observations of the action sequences (picking and placing of blocks) resulting in the construction of the final assembly using varied block instances and types (cubes, dice, lego etc.).
+Figure 4: Illustrative examples of spatial structures from the corpus, showing inductive composition over simpler structures. Details and visualizations in Appendix  12
+The scope of concepts and associated evaluation tasks are adapted from closely related works.
+The staircase and enclosure construction tasks are inspired from from (Silver et al., 2019), adapted to 3D from the original 2D grid world setting.
+Structures such as boundaries involving repetitive use of columns and rows (w/o explicit joint fastening) are inspired by a robotic assembly data set (Collins et al., 2024).
+Finally, the arch-bridge and x-shaped patterns are inspired from concept learning works as  (Lake et al., 2015).
+A total of $15$ structures types are incorporated and are additionally modulated in size/spatial arrangement for generalization evaluation.
+Three evaluation data sets are formed each with simple structures and complex structures composed of simpler concepts (e.g., staircase consists of towers as substructure).
+Dataset I and II contain demonstrations constructing structures with $\mathtt{size}(.)\in[3,5]$, where $\mathtt{size}$ is defined in  4.
+Dataset II reverses the linguistic labels used (e.g., the “tower" in I becomes “rewot" in II) to assess model reliance on pre-training knowledge in presence of new labels for concepts.
+Dataset III includes concepts of larger size than those in training to test generalization.
+Baselines. Four baselines are formed from two alternative approaches as follows.
+(1) Purely-Neural: An end-to-end neural model inspired by StructDiffusion (Liu et al., 2023) that treats structure construction as a rearrangement problem.
+We consider two variations: (1.1) Struct-Diff+Grounder(SD+G): End-to-end approach that assumes a perfect object selector/grounder which identifies the relevant set of objects to be moved.
+(1.2) Struct-Diff(SD): (1.1) without assuming a grounder.
+(2) Pre-trained models that directly output symbolic programs: (2.1) LLMs for Scene-Graph Reasoning: This approach uses a Pre-trained Language Model (GPT-4) to generate Python programs from instructions which describe the given demonstration.
+Input is provided to the LLM via textual symbolic spatial relationships (e.g., left(a,b)) between objects in the demonstration.
+This baseline further assumes the absence of distractor objects in the scene.
+(2.2) Vision Language Model (GPT-4V): Similar to (2.1) but has the ability to take input demonstration as images.
+For learning the program of a new inductive concept, we give the demonstration to the VLM in the form of $\Lambda,(S_{1}..S_{g})$.
+Additional details on prompting method in Appendix C.6.
+Experiments were also conducted with open-source LLMs such as CodeLlama (70Bq), Due to significantly poorer performances w.r.t.
+GPT-4, GPT-4 was retained as the primary LLM baseline.
 Model variants.
-We implement three variants of the MCTS search to perform a grounded plan search over the action space $\mathcal{A}$: (i) MCTS+P+L: Our approach as described in section  4.2, that uses the learnt concepts from $\mathcal{L}$ as macro actions (e.g., Make_Tower(3, objects) $\in\mathcal{A}_{c}$) (L).
-Further it performs pruning of $\mathcal{A}_{p}$ using $\pi_{\text{neural}}$ (P), (ii) MCTS-P+L: Our approach without neural pruning and (iii) MCTS+P-L: No access to library of concepts during continual learning and thereby lacks ability to use macro actions.
+We implement three variants of the MCTS search to perform a grounded plan search over the action space $\mathcal{A}$: (i) MCTS+P+L: Our approach as described in section 5.2, that uses the learnt concepts (L) from $\mathcal{L}$ in subsequent searches, along with pruning (P) $\mathcal{A}_{p}$ using $\pi_{\text{neural}}$, (ii) MCTS-P+L: Our approach without neural pruning and (iii) MCTS+P-L: No access to library of concepts during continual learning.
 This method greedily selects the action from $\mathcal{A}_{p}$ as given by $\pi_{\text{neural}}$.
-We provide more details about the 3 methods in Appendix  D.4.
+For details refer App.
+D.4.
 Metrics.
 We adopt the following metrics to evaluate our models: (i) Program Accuracy: A binary score obtained through human evaluation.
-(ii) Target Construction IoU: Intersection over Union (2D-IoU) between the bounding boxes (excluding depth).
-(iii) Target Construction Loss: Mean Squared Error (MSE) loss between bounding boxes (including depth).
-Note that both IoU and MSE are computed invariant to order between objects with same set of attributes.
+$1$ for constructing the structure fully, 00 otherwise.
+(ii) Target Construction IoU: Intersection over Union (2D-IoU) between bounding boxes.
+(iii) Target Construction Loss: Mean Squared Error (MSE) loss over the bounding boxes + depth of the center.
 8 Results
-Our experiments evaluate the following (i) How does our model perform as compared to baselines in terms of concept learning and execution ability (In-Distribution) ?
-(ii) How does our model perform in comparison to baselines in ability to generalize to concept instances larger than those seen during training (Out-of-Distribution) ?
-(iii) How robust and efficient is our learning pipeline ?
-(iv) How can the acquired concepts be used in complex reasoning and planning tasks ?
-6.1 Effectiveness of Concept Learning
-Tables  2 and  2 compares the the program accuracy [^2] Purely neural approach is marked as NA due to absence of program representations.
-and IoU/MSE metrics on attaining the final state of the proposed and the baseline approaches in the in-distribution setting.
-The proposed SPG approach significantly improves over other approaches.
-We make the following observations.
-First, for complex compositional structures, the accuracy of the pre-trained models is poor (zero), indicating their inability to reason over numerous and complex spatial relations inherent in structures (e.g., staircases, bridge, x-shaped etc.).
-Additionally, we note that program inference via an LLM (GPT4) shows improved performance learning simple structures (e.g., rows or columns) using the tokenized language description of the scene compared to a VLM (GPT-4V) provided with a visual input (reflecting in better IoU/MSE scores).
-However, we observe that for complex structures, the performance of VLM (GPT-4v) is better than an LLM (GPT) indicating the difficulty in complex spatial reasoning inherent in interpreting complex concepts.
-Purely neural approach that is completely grounded in the demonstration performs much better on complex structures as compared to pre-trained foundation models.
-However our approach still outperforms it and is more data efficient.
-Table 1: Program Accuracy (PA)
+Our experiments evaluate the following questions.
+Q1: How does our model perform when compared to baselines in terms of concept learning and execution ability (In-Distribution)?
+Q2: How does our model generalize to concept instances not seen (larger) during training (Out-of-Distribution)?
+Q3: How robust and efficient is our concept learning pipeline?
+Q4: How can the acquired concepts be used in for embodied instruction following tasks?
+Q1: Concept Learning Accuracy
+We compare the program accuracy (Table 1) and the IoU/MSE values (Table 2) of the final states attained by SPG and the baselines w.r.t.
+the gold states in the in-distribution setting and find that SPG significantly outperforms other approaches.
+Values for Purely neural approaches are marked NA because Neural Outputs are not physically grounded.
+We make the following observations: (i) For complex compositional structures, the accuracy of the pre-trained models is poor (zero), indicating their inability to reason over the numerous and complex spatial relations present in these structures.
+(ii) While program inference via the LLM is better than the VLM for learning simple structures, it is worse for complex structures.
+This indicates the inherent weakness of the textual descriptions of complex spatial relations present in complex structures.
+(iii) While the data-intensive purely neural approaches perform much better on complex structures when compared to the pre-trained foundation models, they are still weaker than SPG.
+Table 1: Program Accuracy
 Model Simple Complex
@@
 SD 0.49 $\pm$ 0.01 1.48 $\pm$ 0.24 0.46 $\pm$ 0.02 3.71 $\pm$ 1.53
-6.2 Generalization Performance
-Table  4 compares the generalization performance on Dataset III [^3] Full table with standard-errors reported in Appendix,  8.
-The proposed model outperforms other approaches.
-Further, we observe a relative decrease in 2D-IoU being 7.27% for simple and 5.74% for complex structures.
-In contrast Struct-Diffusion with grounder shows a relative decrease of 63.25% on simple structures and 74.72% on complex structures; highlighting their inability to generalize inductively.
-Pre-trained models on the other hand have much higher relative decrease in performance for complex structures (53.87% and 41.64%), which is attributed to their inability generate the correct program that can generalize inductively to unseen data.
-Table 9 in Appendix quantifies relative performance decrease values as compared to Dataset I.
-Table 3: Out-of-Distribution Performance
+Q2: Generalization Performance
+Table 3, compares the generalization performance on Dataset III for models trained on Dataset I (full table in Appendix,  9).
+We see that SPG outperforms other approaches.
+We further consider the relative decrease (R.D.) in performance (2D-IoU) on going from the in-distribution to the out-of-distribution (OOD) setting.
+We make the following observations: (i) SPG suffers a R.D.
+of 7.27% for simple and 5.74% for complex structures.
+(ii) In contrast, the SD+G baseline shows a large R.D.
+of 63.25% on simple structures and 74.72% on complex structures; highlighting the inability of Purely Neural Models to generalize inductively.
+(iii) Pre-trained models also have a large R.D.
+in perf.
+for complex structures (GPT-4 : 53.87% & GPT4V : 41.64%), which is attributed to their inability to generate the correct program that can generalize inductively to unseen data.
+Table 3: OOD Performance. R.Dis the relative decrease in IoU from Table 2. MSE is in 1e-3 units
 Model Simple Complex
-IoU MSE (1e-3) IoU MSE (1e-3)
-SPG(Ours) 0.89 0.43 0.80 1.49
-GPT-4V 0.58 13.2 0.29 10.9
-GPT-4 0.78 5.51 0.13 19.1
-SD+G 0.27 6.21 0.15 14.2
-SD 0.24 6.86 0.15 11.6
-Table 4: Perf. on Dataset II: Names Reversed
+IoU R.DMSE IoU R.DMSE
+SPG(Ours) 0.89 7.27 0.43 0.80 5.74 1.49
+GPT-4V 0.58 23.33 13.2 0.29 41.64 10.9
+GPT-4 0.78 12.61 5.51 0.13 53.87 19.1
+SD+G 0.27 63.25 6.21 0.15 74.72 14.2
+SD 0.24 51.84 6.86 0.15 67.67 11.6
+Q3: Robustness and Efficiency Analysis
+Reliance on pre-trained Knowledge vs.
+Demonstration.
+Next, we evaluate the degree to which concept learning relies on prior knowledge vs.
+the action sequences observed in demonstrations.
+We compare pre-trained models against our approach by learning programs on Dataset II7), which uses arbitrary names for concepts.
+This forces all models to rely on demonstration data because there is no real-world knowledge associated with the name of the concept, say "rewot" instead of "tower".
+Table  4 indicates the corresponding performances, with our approach outperforming others.
+For the IoU/MSE values along with standard errors refer to app.
+Table  10.
+The R.D.
+in performance (program accuracy w.r.t.
+Table  1) for simple structures for our approach (12%) is lower than GPT-4 (14%) & much lower than GPT-4V (30%).
+The poorer generalization of pre-trained models can be attributed to their over-reliance on prior knowledge and failure to effectively incorporate the data from demonstrations.
+In contrast, SPG better captures the semantics of a novel concept, especially ones whose knowledge may not be available for the LLMs/VLMs at training time.
+Table 4: Perf. on Dataset II with Reversed Names. Acc. is Prog. Accuracy, MSE in 1e-3 units
 Model Simple Complex
-IoU MSE (1e-3) IoU MSE (1e-3)
-SPG(Ours) 0.86 1.74 0.78 3.93
-GPT-4 0.78 3.16 0.00 22.73
-GPT-4V 0.71 3.92 0.09 21.29
-6.3 Robustness and Efficiency Analysis
-6.3.1 Reliance on pre-trained Knowledge vs. Demonstration.
-Next, we evaluate the degree to which concept learning relies on background knowledge vs.
-observed action sequences for its construction.
-We compare pre-trained models against our approach by learning programs on another training dataset differing only in terms of the names of the structures (names being reversed).
-Table  10 in the Appendix gives the corresponding program accuracies.
-Table  4 indicates the corresponding IoU and MSE, with our approach again showing the best performance [^4] For the IoU/MSE values along with standard errors refer to table  11 in Appendix..
-The relative decrease in performance for our approach (12%) is lower than GPT-4(14%) and much lower than GPT-4-V(30%) (in terms of program accuracy).
-The poorer generalization of pre-trained models can be attributed to over-reliance prior knowledge and failing to effectively use demonstrations.
-6.3.2 Analysis of MCTS Variants for Concept Learning
+Acc. IoU MSE Acc. IoU MSE
+SPG(Ours) 0.88 0.86 1.74 0.78 0.78 3.93
+GPT-4V 0.23 0.71 3.92 0.00 0.09 21.29
+GPT-4 0.67 0.78 3.16 0.00 0.00 22.73
+MCTS Variants for Concept Learning.
+Figure  5 compares the program accuracy for the three methods of plan search.
+For the MCTS-L+P method, the program accuracy is expected to be independent of expansion steps as it greedily chooses the action for which $\pi_{\text{neural}}$ gives the highest probability.
+For the MCTS+L based methods the accuracy increases beyond 0.6 with time, which demonstrates that having a composable library of concepts allows us to learn a much richer class of inductive concepts.
+MCTS+P+L saturates to a program accuracy of 0.933 in just 4000 expansion steps compared to MCTS+L-P taking 512000 expansion steps, which demonstrates a significant increase in learning efficiency via use of the neural pruner.
+For very low number of expansions steps (<40) accuracy of MCTS+L based methods is lower than MCTS-L as the former expends expansion steps on UCB exploration (instead of greedy actions).
 Figure 5: MCTS Variants. Num. of expansion steps in search (log scale) (X-axis) vs Program accuracy (Y-axis).
-Figure  4 compares the program accuracy for the three methods of plan search.
-For the third method the program accuracy is expected to be independent of expansion steps as it greedily chooses the action for which $\pi_{\text{neural}}$ gives the highest probability.
-For the MCTS+L based methods the accuracy increases beyond 0.6 with time, which demonstrates that having a composable library of concepts allows us to learn a much richer class of inductive concepts.
-MCTS+P+L saturates to a program accuracy of 0.933 in 4000 expansion steps, whereas MCTS-P+L takes 512000 expansion steps, which demonstrates a significant increase in learning efficiency via use of a neural prunner.
-For very low number of expansions steps (<40) accuracy of MCTS+L based methods is lower than MCTS-L as they spend some expansion steps on exploration based on UCB bound (rather than taking greedy choice).
-6.3.3 Effectiveness of Continual Learning of Visual Concepts
-Figure 5: Acquisition of new visual concepts. Likelihood of correct grounding of the referred object as training proceeds.
-Having a disentangled representations allows us to 1) intersperse learning of new visual attributes with learning of new inductive concepts 2) avoid catastrophic forgetting of already learnt attributes.
-For example, the model can learn the chocolate from an instruction “construct a tower using chocolate blocks of size 4”, even if it has not seen the color in the pretraining phase.
-Because of our modular architecture, we can learn the color as a new embedding in the space of visual attributes.
-Plot  5 demonstrates the benefit of having such disentangled representations.
-As the training proceeds the probability of selecting the chocolate block increases with time.
-Further, the likelihood of selecting a magenta colored object (in a different scene) is unchanged and the problem of catastrophic forgetting is avoided.
-Additional details for continual learning of visual concepts appear in Appendix  D.2.
-6.4 Application of Learnt Concepts for Embodied Instruction Following
-Complex instruction execution via LLM.
-We demonstrate our ability to build over the learnt program representations and perform complex guided robot manipulation tasks, in a simulated Franka Panda environment.
-E.g., consider instructing a robot to : “Construct a tower of green die having the same height as the existing tower of white die.” or “Construct a tower of total 6 blocks using alternating blue and red blocks”.
-For both the above tasks we prompt GPT-4 by providing it with the set of inductive concepts that are already learnt, the set of primitive actions, and some pre-defined helper functions by using Python import statements in a manner similar to [code-as-policies].
-GPT-4 generates an executable Python code in terms of these functions which on running generates the resultant trajectory.
-Top portion of figure  6 shows the execution for both these tasks.
-We provide details in Appendix  D.3.
-Grounding learnt concepts into visual input for plan synthesis.
-We further demonstrate that the concepts we have acquired can help us to perform goal conditioned planning.
-Bottom portion of figure  6 demonstrates the results of our approach for the task of constructing a staircase beginning from an adversarial and assistive initial state.
-Note, the planner that we learn is a grounded neuro-symbolic planner, as a PDDL based planner cannot be hand-coded easily ( D.5), and LLMs/VLMs are unable to perform such complex reasoning tasks.
-Relevant details are provided in the Appendix  D.5
-Figure 6: Application of learnt concepts.
-Top: Using LLM to generate the executable code for a novel tasks, given the concept definitions.
-Bottom: Integrating a neuro-symbolic planner over the concepts.
-In the bottom-left figure the planner optimally outputs the sequence of actions to replace the green cube.
-In the bottom-right, the planner is able to complete a staircase from an initially constructed row.
-7 Limitations
-Currently our methodology and evaluation setup assumes that each object is visible in the initial scene and does not handle partial observability.
-We assume all the objects are cubes of similar size in order to ensure dynamic stability of structures while stacking.
-During training we assume tracking and the ability to associate object with corresponding instances.
-We assume that the demonstration data given for learning is free of errors.
-8 Conclusions
-This paper introduces a novel approach for learning inductive spatial concepts as neuro-symbolic programs from language-annotated demonstrations.
-We observe that direct generation of a program via an LLM/VLM does not generalize to new concepts and purely neural approaches fail to learn the inherent inductive prior inherent in such concepts.
+Significance of MCTS in SPG.
+To assess the necessity and importance of MCTS, we carry out an ablation study where we replace it with an LLM planner during the planning stage, referred to as SPG-M+LMP.
+In the "plan" stage of our pipeline, GPT-4V is prompted to output a plan given the concept library and RGB keyframes from the demonstration.
+Our experiment shows that GPT-4V struggles to generate correct plans, particularly for complex structures like pyramids, arch_bridge and boundaries, resulting in significantly lower performance than SPG (Table 5).
+Additionally, some plans generated by GPT-4V are not physically grounded, leading to errors in both the planning and generalization stages, which compounds the inaccuracies.
+This demonstrates that combining symbolic search with LLMs offers a substantial advantage over using only LLMs.
+Table 5: Ablation studies. Ablations with SPG-M+LMP and GPT-4V+VR. MSE values are in 1e-3.
+Model Simple Complex
+Acc. IoU MSE Acc. IoU MSE
+SPG(Ours) 1.0 0.96 0.01 0.83 0.85 2.06
+SPG-M+LMP 0.55 0.68 11.1 0.16 0.19 20.0
+GPT-4V+VRF 0.66 0.75 6.8 0.16 0.46 12.0
+Ablating with Pre-trained Models + Visual Reward Filter In line with program synthesis techniques using LLMs (Li et al., 2022; Chen et al., 2021), we sample five programs from GPT-4V and rank them according to the visual reward obtained from their execution.
+Furthermore, we provide the ground-truth programs of the concepts that are needed to learn the given new concept, thus employing a form of teacher forcing in program generation.
+Even with these measures, it performs significantly worse than SPG, see Table 5 (GPT-4V+VRF).
+While this performance is better than that of GPT-4V, it still unable to generate correct programs, especially for complex structures.
+Q4: Incorporating Constraints into learnt concepts
+Instructions with Discrete Constraints:Consider the instruction $\mathcal{\lambda}_{1}$: “Construct a staircase of size 5 such that all blocks have the same color as the block to their left.
+No block should have the same color as the block on top of it.” This example highlights the interpretive & compositional strength of our pipeline.
+We model the constraints as a CSP problem & use the Z3 solver (De Moura et al., 2008) to resolve them.
+The resulting grounded program for $\mathcal{\lambda}_{1}$ is executed via the planner, producing the desired 3D structure.
+Fig 6 shows the executed output in the simulator, satisfying both the structural concept and the imposed constraints.
+Figure 6: Constructed staircase obeying all constraints in $\mathcal{\lambda}_{1}$, which are extracted via an LLM and solved via Z3 CSP-solver.
+For simpler instructions such as $\mathcal{\lambda}_{2}$: ”Construct a tower of green die having the same height as the existing tower of white die.” and $\mathcal{\lambda}_{3}:$”Construct a tower of total 6 blocks using alternating blue and red blocks.”, the LLM (GPT-4) is able to understand the structural requirements (through the library $\mathcal{L}$) as well as the (simpler) constraints.
+Therefore, the pipeline (with access to tower through $\mathcal{L}$ or via SPG framework), generates Python code in terms of these functions, which on execution generates the required action sequence as shown in Fig.
+7.
+(also see app.
+D.3).
+Figure 7: Using an LLM to solve the constraints in a novel task and generate executable code, given the concept definitions.
+Efficient Plan Synthesis for Grounded Goals: As detailed in sec.
+6 and App.
+D.5 , we demonstrate the ability to perform goal-conditioned planning using the acquired concepts.
+Fig 8 demonstrates the results of our approach for the tasks of constructing a staircase beginning from adversarial and assistive initial states.
+Figure 8: Integrating a neuro-symbolic planner over the concepts.
+Top: The planner is able to optimally replace the green cube from the adversarial initial state by unstacking and re-stacking the faulty tower.
+Bottom: The planner is able to complete a staircase from an initially constructed row by layering rows upon rows, a method of construction it has not seen while learning staircase.
+9 Conclusion
+This paper introduces a novel approach for learning inductive representation of grounded spatial concepts as neuro-symbolic programs via language-guided demonstrations.
 Our approach factors program learning as: Sketch: generating the high-level program signature via an LLM, Plan: searching for a grounded plan that maximises the total discounted reward with the respect to the demonstration, and Generalize: abstracting the grounded plan into an inductively generalize-able abstract plan via an LLM.
-Continual learning is achieved via learning of modular programs by giving preference to shorter programs (through composition of learnt ones) and allowing for learning of primitive neural concepts via gradient updates.
-Extensive evaluation demonstrates accurate program learning and stronger generalization in relation to baselines.
-Further, we show the grounding of learned concepts in visual data to facilitate reasoning and planning for embodied instruction following.
-[…]
+Continual learning is achieved via learning of modular programs by giving preference to shorter programs through composition of learnt ones.
+Extensive evaluation demonstrates accurate program learning and stronger generalization in relation to purely LLM based as well as purely neural baselines.
+Grounding of learned concepts in visual data facilitates reasoning and planning for embodied instruction following.
+Limitations include reliance on perfect demonstrations, assumption of full observability of all objects and experiments confined to simulation.
+Incorporating noisy demonstrations, reasoning with beliefs and interleaving planning and execution remains part of future work.
 Appendix A Additional Details on Technical Approach
-Figure  7 illustrates the pipeline for learning inductive concepts.
-Figure 7: SPG: Concept Learning Learning of grounded inductive programs is factored into three phases (a) Sketch: generating a sketch for the task intended in the instruction in the context of the environment (b) Plan: searching for the likely plan (guided by physical rewards) to construct the given assembly and (c) Generalize: generating a program representation to explain a class of demonstrations.
-Finally, our architecture allows learning of increasingly complex symbolic abstractions composed of simpler ones.
-Figure  8 illustrates the pipeline for online inference to realize to realize construction of novel structures.
+Figure  9 illustrates the pipeline for online inference to realize to realize construction of novel structures.
 Figure 9: SPG: Inference First the library of concept $\mathcal{L}$ is loaded with the corresponding set of learnt programs.
@@
 Standard Python constructs such as for loops, if else $\cdots$) as assumed in addition to the constructs defined here.
-A.2 Ground-Truth Inductive Concepts
+Table 6: Symbols and Semantics Signature and semantics for the primitive concepts and operations that are used in the construction of the programs used to express inductive spatial concepts.
+Function Signature Semantics
+filter (VisualConcept, ObjSet) $\rightarrow$ ObjSet Returns the objects that contain the VisualConcept
+move_head (Head, Dir) $\rightarrow$ Head Moves the head to the given direction (May or may not take input/return the head, based on a flag)
+assign_head a.k.a move_head(overloaded) (Head, ObjIdx) $\rightarrow$ Head Given the index/one-hot representation for an object, it moves the head to the position corresponding to that object
+keep_at_head (ObjSet, Head) $\rightarrow$ None Keeps the argmax of ObjSet at the head
+reset_head None $\rightarrow$ Head Sets the head to the top position of stack and pops this position from the stack as well
+store_head Head $\rightarrow$ None Pushes the current position of head into the stack
+Table 7: Symbolic representation. The table lists the type definitions used in the implementation of SPG programs.
+Defined Types Python Type Usage
+IntArg int Argument for the structures that defines the size (height, length etc)
+Obj torch.Tensor One-hot vector whose non-zero index represents the selected objects
+ObjSet torch.tensor Probability mask over the selected objects
+Dir string Primitive directions like left, right, front, top, etc
+ConceptName string name of the visual, action or inductive concept
+Head torch.Tensor Bounding box with depth. 3D cuboidal space.
+A.2 Curriculum Learning
+We follow a curriculum approach where the visual concepts are trained first from simpler linguistically-described demonstrations ($t_{0}$ in figure  10).
+This is followed by learning of action concepts through sequentially composed pick and place tasks.
+Its essential to use such long range sequential instructions in order to ensure that the semantics of action concepts are learnt for placement of objects at a height much above the tabletop ($t_{1}$ in figure  10).
+After the pre-training phase, the agent can continually learn new inductive concepts and visual attributes.
+($t_{2},t_{3}$ in figure  10)
+Figure 10: Continual learning through curriculum: Using simple pick and place demonstrations we learn visual attributes such as blue cube, yellow lego ($t_{0}$).
+Using long range instructions which are sequentially concatenated descriptions of pick and place tasks we train our action concepts such as left, right, top ($t_{1}$).
+After pre-training the agent can perform continual learning of concepts such as learning generalized representation for tower ($t_{2}$).
+Because of disentangled representation of neural and symbolic concepts, interspersed learning of new visual attributes such as chocolate color are also possible through few demonstrations of structure creation($t_{3}$).
+A.3 Details for Plan-Search and Generalization
+Modifications to the Simulation and Reward Back propagation Steps: Next, we outline the modifications in the simulation and the reward back propagation steps of the standard MCTS algorithm for our setting.
+During program search we assume access of intermediate scenes in the demonstration.
+This allows us to provide intermediate rewards that can guide the search well.
+We observed that making the following changes in simulation and back propagation step increased the efficiency of our search procedure.
+Fig 11 illustrates the possible states explored by MCTS and the reward calculation.
+  • Simulation: Rather than performing Monte Carlo simulations at each newly expanded leaf node (to estimate its value) we completely avoid these simulation steps.
+This was motivated by the fact that our reward is not completely sparse and the intermediate IoU rewards for each object we place allow us to guide the search effectively.
+  • Back propagation: We perform off policy Q-learning updates during back propagation similar to one indicated by (Khandelwal et al., 2016) :
+    $V(s_{t})=max_{a\in\mathcal{A}}Q(s_{t},a).$ (4)
+    $\tau(s_{t},a)=s_{t+1}$ (5)
+    $Q(s_{t},a)=r_{t}+\gamma V(s_{t+1})$ (6)
+Figure 11: A sample MCTS search tree outlining the states explored and the calculation of reward.
+Improving Modularity and Scalability of MCTS procedure: We present additional details on the MCTS procedure for searching for a plan conditioned on a program signature and guided by the demonstration.
+In order to incorporate the objective of searching for physically realizeable plans and facilitating generation via re-use of concepts, the following conceptual changes are incorporated in the standard MCTS procedure (Sutton & Barto, 2018).
+Modularity (MCTS + L): We want to allow learning of novel inductive concepts in terms of existing ones.
+This would ensure that the plan $H_{P}^{*}$ corresponding to a given demonstration is concise and can be easily generalized to the $H_{G}^{*}$.
+C.3 in appendix give example of two plans for the structure Pyramid one which is modular and can be successfully generalized by GPT-4, other for which GPT-4 fails in generalization due to lack of modularity.
+This can be seen as a form of regularization in terms of the length of concept description, by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2)
+$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))+\alpha\log|H|\right]$ (7)
+In order to allow modular learning of programs, for every inductive concept already stored in the library we define corresponding action instantiations which can be a potential candidate actions during our search.
+As an example, for the concept Tower we have one of the action instantiation as Make_Tower(3, objects) which would be the action of constructing desired tower.
+This can be visualized as a compound/macro-action which is composed of primitive actions keep_at_head(objects), move_head(‘top’).
+We define $\mathcal{A}_{c}$ as the space of such compound actions, and $\mathcal{A}_{p}$ as the space of primitive action/function consisting reset_head(), move_head(direction), keep_at_head(objects), store_head().
+Realization of equation (6) (increased preference of macro-actions over primitive ones) is done through discounted IoU rewards during our search (Make_Tower(3) would have a reward of 1+1+1, as compared to 1 + $\gamma^{2}$(1) + $\gamma^{4}$(1) for a sequence of 3 (keep_at_head(objects), move_head(’top’)0)).
+We refer to MCTS using concept instantiations from $\mathcal{L}$ as macro actions in search as MCTS+L.
+Scalablility (MCTS+P): as more concepts are added to the library $\mathcal{L}$ the action space of our search $\mathcal{A}$ (specifically $\mathcal{A}_{c}$, the space of compound concepts) increases, therefore we want to prune the search space effectively.
+For this during the pre-training phase we train a reactive policy $\pi_{neural}$ which given the current state, $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (part of the demonstration) would output one of the primitive action, $a_{t}^{*}\in\mathcal{A}_{p}$ where $\mathcal{A}_{p}$ is the primitive action space, i.e.
+$a_{t}^{*}=\pi_{neural}(a_{t}|\tilde{s_{t}},{s_{t+1}}),\;\;a_{t}\in\mathcal{A}_{p}$ Note that while expanding our search tree we only search among the space of compound actions $\mathcal{A}_{c}$ and the action $a_{t}^{*}$, thereby reducing the branching factor of search from $|\mathcal{A}_{c}\cup\mathcal{A}_{p}|$ to $|\mathcal{A}_{c}|+1$.
+We refer to MCTS using neural pruning as MCTS+P.
+Therefore our MCTS algorithm is modular through hierarchical composition of learnt concepts and efficient through pruning of action space and is referred to as MCTS+L+P.
+Generalization: GIven multiple equal length plans for a given demonstration, we seek to recover a plan one that can be easily generalized by the LLM.
+LABEL:lst:correct_gen shows a plan which could be correctly abstracted out into a generic program by GPT-4.
+Whereas  6 shows another plan with similar semantics, for which GPT-4 is unable to correctly find the generalized program (Note that row or column of size 1 is equivalent to keep_at_head).
+We tackle this problem in the following manner.
+  1. 1.
+    Rather than getting a single plan from the plan search we get the top k plans $\{H_{P,i}\}_{i=1}^{i=k}$.
+In order to get these top k plans we expand the complete tree (based on UCB criteria) starting from the root node corresponding to the initial state, till a predefined budget of expansions.
+Then we select the top k paths(potential plans) from the root node to all the leaf nodes (where the top k ones are those that give the highest accumulated IoU reward with respect to the given demonstration).
+  2. 2.
+    Later we abstract out each of these k plans into corresponding generalized programs, $\{H_{G,i}\}_{i=1}^{i=k}$ using GPT-4.
+We again run each of these programs on the given demonstration and then choose the one which gives the highest IoU reward (resolving ties based on predefined order).
+Note that some program $H_{G,i}$ upon execution may result in a plan $\tilde{H}_{P,i}$ different from the original plan ${H}_{P,i}$ using which it was generalized.
+This can be attributed to potential errors in GPT-4s program generalization process.
+A.4 Additional Details: Learning with Increasing Number of Demonstrations
+Given k demonstrations for a novel inductive concept, we independently find k task sketch $\{H_{S,i}^{*}\}_{i=1}^{i=k}$ and grounded plans $\{H_{P,i}^{*}\}_{i=1}^{i=k}$.
+During the generalization phase we give these k pair of task-sketch and corresponding plans to GPT-4 and ask into infer a single abstraction over them.
+C.5 in appendix gives a concrete example.
+Equation for generalize step (getting $H_{G}^{*}$ from $H_{P}^{*},H_{S}^{*}$) can be modified as follows.
+$H_{G}^{*}\leftarrow Generalize(H_{G}\,|\,{\{H_{P,i}^{*},\,H_{S,i}^{*}\}}_{i=1}^{i=k};\,{\theta_{G}}),\;\;H_{G}\in\mathcal{H}^{\mathcal{G}}$ (8)
+A.5 Detailed Experimental Methodology
+Input to the Method: The input consists of a language instruction and a human demonstration represented as a sequence of RGBD keyframes.
+Output/Aim of the Model: The goal is to learn a representation of the unknown concept in the instruction, assuming there is only one unknown concept.
+If the unknown concept is inductive (e.g., "tower"), the model learns a program definition $deftower()$ and stores it in the program library.
+If the unknown concept is a primitive concept, a concept embedding is learned through backpropagation.
+Evaluation: The learned model is evaluated based on the correctness of the program representation for inductive concepts and the correctness of object placements, measured through the Intersection over Union (IoU) metric (see Metrics, line 262).
+d.
+Examples: Suppose the current library contains the concepts “red”, “tower”.
+Given the instruction “construct a staircase of height 3 using red blocks,” the process is as follows:
+Parsing: The instruction is parsed into a sketch: Staircase(height=3, objects=filter(red, blocks)).
+Grounding: The “objects” parameter is grounded using the visual grounder, which identifies the indices of the red blocks, e.g., [1, 2, 3, 4, 5, 6]. That is, filter(red, blocks) = [1,2,3,4,5,6].
+Planning: The planning step uses the demonstrations (sequence of keyframes) to identify the sequence of actions that best explains the demonstrations.
+In this case, the plan might be: Tower(height=1, objects=[1,2,3,4,5,6]), move_head(right), Tower(height=2, objects=[2,3,4,5,6]), move_head(right), Tower(height=3, objects=[4,5,6]
+Generalization: The generalization step abstracts the plan obtained from three such demonstrations into a program. The resulting program would be:
+1def staircase(height, objects):
+2 for i in range(height):
+3 tower(height=i, objects)
+4 move_head(right)
+Note: Whenever an object is placed, the objects list is modified in place, and the index of the placed object is removed.
+primitive actions: The movement of any object is achieved by first determining the placement position by moving the head (an imaginary bounding box) in specific directions and then placing the object to be moved at the head.
+The head is implemented as a 3D bounding box defined by coordinates (x1, y1, x2, y2, d), where x1, y1, x2, and y2 are the 2D corners of the bounding box, and d is the depth at the center.
+The primitive action move_head(direction) shifts the bounding box in the required direction.
+The action keep_at_head(object_list) picks the first object in the list and places it at the center of the bounding box.
+Two other primitives, store_head and reset_head, are used to save the current position of the head, allowing the search to return to useful positions later if needed.
+Appendix B Additional details regarding datasets
+Figure  12 demonstrates the kind of inductive concepts for which we want to learn generic (i.e instance agnostic) representations.
+Figure 12: Illustration of the inductive concepts.
+Dataset for Pre-training: We use 5k examples of constructing twin-towers i.e.
+2 towers adjacent to each other, for learning semantics of move_head(dir), a basic set of visual attributes, reactive policy $\pi_{\text{neural}}$, and neural modules required for grounded planning.
+The twin towers allow us to learn various action semantics for all possible configurations of blocks in 3D-space (and not being limited to blocks placed directly on table top surface).
+Since we are not aware of the underlying semantics of tower during pre-training phase the corresponding natural language instruction consists of step by step pick and place actions.
+13 gives example demonstrations from this dataset.
+Figure 13: Example from pre-training dataset.
+Dataset for Inductive Structures: We learn a variety of structures which we have divided into Simple and Complex structures.
+A structure is considered complex if it can be expressed as an inductive composition of simpler structures.
+As an example, we can express a staircase to be a composition of towers of increasing height.
+The structures are listed in the Table 8.
+Figure  14 shows the hierarchical relationship among these structures in the form of a DAG (directed acyclic graph).
+F.3 gives the ground truth program representations for each structure.
+Table 8: Structure Types. Examples of simple and complex structures considered in this work for the robot to construct.
+Simple Structures Complex Structures
+Row, Column, Tower X (cross-shape), Staircase
+Inverted-Row, Inverted-Column Inverted-Staircase, Pyramid
+Diagonal-45, Diagonal-135 Arch-Bridge, Boundary
+Diagonal-225, Diagonal-315
+Figure 14: Hierarchy of the structures/programs. This diagram shows the hiearchical nature of the structures in our dataset. MH is abbrevation for move_head and KH is abbreviation for keep_at_head
+Appendix C Prompting Strategy and Examples
+This section gives various prompting examples for our approach and baselines, along with examples motivating particular design decisions in our approach.
+C.1 Prompt Example for Task Sketch Generation Stage (Sketch)
+In order to get a program representation (high level task sketch) of the given natural language instruction, we prompt GPT-4 with few shot examples in a manner similar to (Liang et al., 2023).
+Code segment  LABEL:lst:task_sketch_generation gives an example of getting the task sketch given the demonstration for constructing a staircase.
+We first import the available primitive operators and functions and also give examples in order to demonstrate the signature of the available primitives(line 1-8).
+Then we give incontext examples of how to parse various natural language instructions in a program representation(line 10-14).
+We append to this prompt the instruction for current task(line 16-17).
+1# importing the available functions
+2from visual_operators import filter
+3from inductive_operators import get_parameters, find_structure
+4
+5# function signature of the imported functions
+6filter(color, cube) # filter the objects that are cubes and color ..
+7get_parameters(structure) # parameters of the structure ...
+8find_structure(type, description) # finding structure of given type, description
+9
+10# examples:
+11# instruction: Find the tower with green cubes
+12find_structure(type = tower, description = filter(green, cube))
+13# instruction: Construct a tower of height 3 with yellow cubes
+14Tower(height = 3, objects = filter(yellow, cubes))
+15
+16# current task: Construct a staircase of 4 steps using cyan legos
+17
+18# (GPT-4s output)
+19Staircase(steps = 4, filter(cyan, legos))
+Listing 1: Task Sketch Generation Using GPT-4 (Sketch)
+C.2 Prompt Example for Generalizing a sequence of actions/plan to a general program (Generalize)
+Code segment  LABEL:lst:generalization give an example of getting the general Python program from the plan found using MCTS.
+We first provide a base prompt giving details to GPT-4 about the desired task (line 1-2).
+Then we give the input arguments and the corresponding output/plan for a given demonstration (line 3-5).
+We expect the GPT-4 to output the final Python program (line 7-10)staircase (line 6-8).
+1# Write a general python code which on the given input produces the desired output, do not output anything other than the function description.
+2
+3# input: n = 4, objects = ObjSet
+4# output: tower(1, ObjSet), move_head(‘right’), tower(2, ObjSet), move_head(‘right’), tower(3, ObjSet), move_head(‘right’), tower(4, ObjSet)
+5
+6# Program (GPT-4s output)
+7def staircase(n, objects):
+8 for i in range(n):
+9 tower(i+1, objects)
+10 move_head(’right’)
+Listing 2: Plan to Program using GPT-4 (Generalize)
+C.3 Benefit of Estimating Modular/Smaller Plans
+The below examples demonstrate the benefit of learning new inductive concepts in terms of already acquired inductive concepts (more modular representation).
+The program  LABEL:lst:pyramid_rows is obtained through generalization of plan $H_{p}$ that represents pyramid in terms of rows of decreasing size.
+The generated program generalizes to pyramid of different height.
+1# input: n = 3, objects = ObjSet
+2# output: row(5, ObjSet), move_head(’right’), move_head(’top’), row(3, ObjSet), move_head(’right’), move_head(’top’), row(1, ObjSet)
+3
+4# Program (GPT-4s output)
+5def pyramid(n, objects)
+6 for i in range(n, 0, -1):
+7 # Calculate the number of objects in the current row
+8 row_count = 2 * i - 1
+9
+10 # Perform actions for the current layer
+11 row(row_count, objects)
+12 if i > 1: # No need to move head after the last row
+13 move_head(’right’)
+14 move_head(’top’)
+Listing 3: Pyramid in terms of rows
+On the other hand program  LABEL:lst:py-prim which is obtained in an attempt to generalize the plan expressed in terms of primitives is incorrect.
+1# input: n = 3, objects = ObjSet
+2# output: store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’top’), move_head(’right’), store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’right’), move_head(’top), keep_at_head(ObjSet)
+3
+4# Program (GPT-4s output)
+5def pyramid(n, objects)
+6# Store initial position
+7 store_head()
+8
+9 # Initial sequence of placing and moving right
+10 for _ in range(n):
+11 keep_at_head(objects)
+12 move_head(’right’)
+13
+14 # Reset and move to top right position
+15 reset_head()
+16 move_head(’top’)
+17 move_head(’right’)
+18
+19 # Store again and repeat placing and moving right
+20 store_head()
+21 for _ in range(n):
+22 keep_at_head(objects)
+23 move_head(’right’)
+24
+25 # Reset and adjust final position
+26 reset_head()
+27 move_head(’right’)
+28 move_head(’top’)
+29 keep_at_head(objects)
+Listing 4: Pyramid in terms of primitive actions
+C.4 Comparing Same-length Plans in terms of Generalizability
+Below we illustrate plan to program generalization for 2 different plans which are correct and have same length, but LLMs find it difficult to generalize the second one as compared to first.
+(Note that row and column of size 1 are equivalent to keep_at_head)
+1# input: n = 3, objects = ObjSet
+2# output: keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet)
+3
+4
+5# Program
+6def tower(n, objects):
+7 for _ in range(n):
+8 keep_at_head(objects)
+9 move_head(’top’)
+Listing 5: Plan for tower that can be easily generalized (correct generalization)
+1# input: n = 3, objects = ObjSet
+2# output: row(1, ObjSet), move_head(’top’), keep_at_head(1, ObjSet), move_head(’top’), column(1, ObjSet)
+3
+4
+5# Program
+6def tower(n, objects):
+7 for i in range(1, n + 1):
+8 row(i, objects)
+9 move_head(’top’)
+10 keep_at_head(objects)
+11 move_head(’top’)
+12 if i < n:
+13 column(i, objects)
+14 move_head(’top’)
+Listing 6: Plan for tower that is difficult to generalize (Incorrect generalization)
+C.5 Generalizing via Multiple Demonstrations
+Given multiple demonstrations we independently find task sketch and corresponding grounded plans for each demonstration.
+These are further given to GPT-4 for generalization.
+Code segment  LABEL:lst:mul_plan_gen gives an example of getting a single Python program from multiple demonstrations.
+Note that we explicitly prompt the LLM that some of the grounded plans might be incorrect (which may lead to more robust generalization in case of noisy demonstrations).
+1
+2# Function Call: wor(height = 3, objects = ObjSet_1)
+3# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
+4# Function Call: wor(height = 3, objects = ObjSet_1)
+5# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
+6# Function Call: wor(height = 3, objects = ObjSet_1)
+7# Execution: column(size=1, obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
+8
+9#Write the function definition, which generalizes the above executions. Note that some of the executions can be partially wrong.
+10‘‘‘python
+11def wor(height, objects):
+12‘‘‘
+13
+14GPT-4s Output .....
+Listing 7: Generalizing through multiple plans
+C.6 Prompt Examples for Learning Programs using LLM/VLM Models
+Below we describe the prompting methodologies for learning programs through LLM/VLM models.
+Note that although the prompt examples described below are for the case of learning novel structure from 1 demonstration, we use 3 demonstration per novel structure in our main results (for both our approach and LLM/VLM baseline).
+LLM/GPT-4 Code segment LABEL:lst:llm_baseline_code depicts our prompting methodology given a demonstration for a new concept tower.
+For this baseline we aim to check demonstration following and spatial reasoning abilities of LLMs (GPT-4).
+We provide supervision of the intermediate scenes by using tokenized spatial relations between objects in the scene (Given in the form of Scene = [right(1, 0) …]).
+We further assume that only those objects that are required to perform the task are present in the scene (no distractor objects).
+For every structure (that needs to be learned at time t) we give LLM a prompt providing in-context example on how to generalize (line 19-35), the set of primitive operators (line 4) available and the set of structures learnt/present in library (till time t-1) (line 5-18).
+Finally we append to this prompt the expected declaration (arguments and keywords arguments) of the inductive concept that is to be learnt along with the spatial relations for each scene of the given demonstration (36-51).
+Note that we assume absence of distractor objects for this baseline.
+1# Consider a block world domain ......
+Given a structure creation task along with intermediate scnes complete a general Python function for it.
+The function should be in terms of primitive operators and already learnt structures that are present in the program library.
+Enclose the function within backtick (‘‘‘)
+2
+3primitive_operators = [keep_at_head, move_head ..]
+4# this would be our program library
+5learnt_structures = {
+6 "row": {
+7 "program_tree":
+8 ’’
+9 def row(size, objects):
+10 for i in range(size):
+11 keep_at_head(obj = objects)
+12 move_head(dir = right’)
+13 ’’,
+14 },
+15
+16 ......
+17}
+18# the example task
+19Example task:- Place all the objects to the right of each other.
+20Final state :- [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
+21Intermediate scenes :-
+22Scene 0 = []
+23Scene 1 = []
+24Scene 2 = [right(1, 0)]
+25Scene 3 = [right(1, 0), right(2, 1)]
+26Scene 4 = [right(1, 0), right(2, 1), right(3, 2)]
+27Scene 5 = [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
+28Python function :-
+29‘‘‘python
+30def placing_all_right(objects):
+31 for i in range(len(objects)):
+32 keep_at_head(objects) # select one object from the objects set and keep the head at this location
+33 move_head(dir = ’right’) # move the head to the right of the previous position
+34‘‘‘
+35# The current task for which program needs to be found
+36Current task:- Construct a tower of size 6.
+37Final state :- [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
+38Intermediate scenes :-
+39Scene 0 = []
+40Scene 1 = []
+41Scene 2 = [top(1, 0)]
+42Scene 3 = [top(1, 0), top(2, 1)]
+43Scene 4 = [top(1, 0), top(2, 1), top(3, 2)]
+44Scene 5 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3)]
+45Scene 6 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
+46Python function :-
+47‘‘‘python
+48def tower(size, objects):
+49??
+50‘‘‘
+Listing 8: Prompting Strategy for LLM baselines (GPT-4)
+VLM/GPT-4-V Unlike LLM, VLMs have the abilities to process the demonstration as a sequence of visual frames.
+Therefore rather than providing the symbolic spatial relations between every scene we instead directly provide all the intermediate scenes for the given demonstration.
+Further we also relax the assumption that there are no distractor objects.
+As shown in figure 15 We first give information about the set of primitive operators and the structures that we have already learnt (library of concepts).
+In order to visually ground the semantics of our primitive actions we give 3 example tasks (natural language instruction and intermediate scenes) that do not directly refer to any structure, along with corresponding sequence of actions taken (# Demonstration for visual grounding).
+We further provide another example (without scenes) demonstrating how to write generalizable Python function for a given task using our operators (# Example for generalization).
+Finally we give the natural language instruction and corresponding scenes for the current task along with signature of the program to be learnt (# Current task description).
+Figure 15: Prompt example for VLM-baseline.
+Figure shows the prompting strategy for GPT-4-V that includes the primitive actions, the example tasks for visual grounding, example of writing generalizable Python functions and the demonstration frames.
+Appendix D Supplementary Results
+Out-of-Distribution Performance:
+Table 9: Out-of-Distribution Performance (mean $\pm$ std-error)
+Model Simple Complex
+IoU MSE IoU MSE
+SPG(Ours) 0.892 $\pm$ 0.065 4.386e-4 $\pm$ 2.387e-4 0.804 $\pm$ 0.025 0.001 $\pm$ 5.391e-4
+GPT-4 0.776 $\pm$ 0.023 0.006 $\pm$ 0.001 0.131 $\pm$ 0.019 0.019 $\pm$ 1.498e-3
+GPT-4V 0.575 $\pm$ 0.026 0.013 $\pm$ 0.001 0.290 $\pm$ 0.016 0.011 $\pm$ 1.316e-3
+SD 0.236 $\pm$ 0.005 0.006 $\pm$ 7.495e-4 0.150 $\pm$ 0.011 0.011 $\pm$ 2.860e-3
+SD+G 0.273 $\pm$ 0.004 0.006 $\pm$ 6.180e-4 0.154 $\pm$ 0.010 0.014 $\pm$ 2.958e-3
+Performance Dataset II (i.e. name reversed evaluation): Table  4 gives the corresponding program accuracies, while Table  10 give the corresponding IoU/MSE metrics along with standard errors.
+Table 10: Performance on Dataset II(Names Reversed)
+Model Simple Complex
+IoU MSE (1e-3) IoU MSE (1e-3)
+SPG(Ours) 0.86 $\pm$ 0.03 1.74 $\pm$ 0.45 0.78 $\pm$ 0.02 3.93 $\pm$ 1.09
+GPT-4 0.78 $\pm$ 0.03 3.16 $\pm$ 0.51 0.00 $\pm$ 0.00 22.73 $\pm$ 1.48
+GPT-4V 0.71 $\pm$ 0.01 3.92 $\pm$ 0.45 0.09 $\pm$ 0.02 21.29 $\pm$ 1.59
+D.1 Qualitative Comparison between Purely-neural (Struct-Diff+Grounder) vs. Ours(SPG)
+Figure  16 gives compares the qualitative results for our approach against Struct-Diffusion with grounder on both in-distribution, Dataset I and out-of-distribution (larger size), Dataset III.
+In in-distribution setting the our method performs slightly better in terms of structure creation for both simple and complex structures, but the difference is not significant.
+However for out-of-distribution setting structures created by our approach are much better than those created by Struct-Diffusion+Grounder.
+Further for this setting structure creation by Struct-Diffusion seems to be much worse for complex structures than simple ones.
+Figure 16: Structure creation comparison between SPG(Ours), and Struct-Diff+Grounder
+D.2 Continual Learning of Neural Concepts
+Having a disentangled representations allows us to (i) intersperse learning of new visual attributes with learning of new inductive concepts (ii) avoid catastrophic forgetting of already learnt attributes.
+For example, the model can learn the chocolate from an instruction “construct a tower using chocolate blocks of size 4”, even if it has not seen the color in the pretraining phase.
+Because of our modular architecture, we can learn the color as a new embedding in the space of visual attributes.
+The plot in Fig.
+17 demonstrates the benefit of having such disentangled representations.
+As the training proceeds the probability of being able to select the chocolate blocks when required increases with time, while keeping the ability of selecting a magenta colored object (when required) remains the same.
+Figure 17: Disentanglement.
+The acquisition of new visual concepts.
+Plot shows an increase in the likelihood of correct grounding of an object referenced with a new neural concept (chocolate color) with training iterations.
+Given demonstration for the task “Construct a tower of height 4 using chocolate cubes”, we would like to learn the neural embedding for the unknown color chocolate (where we assume that tower has already been learnt and stored in the library $\mathcal{L}$).
+First the instruction is converted into corresponding plan sketch $\mathcal{H}_{S}$ = tower(4, Filter(chocolate, cubes)), which is passed to the visual grounder.
+The grounder detects the presence of an unknown attribute chocolate as an argument to filter, and randomly initializes a new neural embedding for it.
+Using this new embedding along with the already present embedding of cube and the visual features found through ResNet-34, the quasi-symbolic executor outputs a grounded task-sketch.
+The executor executes the grounded task-sketch by getting the semantics of the underlying function i.e.
+tower from the library $\mathcal{L}$.
+MSE+IoU loss computed over the final scene obtained and the expected final scene is backpropogated through the network.
+Note that during backpropogation all the neural modules (action semantics, visual attributes, ResNet-34) are frozen, except for the newly initialized embedding for chocolate.
+For the purpose of differentiable sampling during tower construction we use gumbel-softmax (Jang et al., 2017) with masking.
+Figure  18 illustrates our approach.
+Figure 18: Continual learning of visual primitives
+D.3 Details for Inference on Novel Tasks using an LLM
+Below we show the (Liang et al., 2023) inspired prompting methodology that we use to get the executable code corresponding to a language specified manipulation task.
+We initially begin by importing the helper functions, spatial direction, primitive functions, and learnt inductive concepts/structures (line 5-11).
+Then we give few examples for how to use and compose the various functions for different tasks (line 16-83).
+Finally we give the instruction of current task, and expect GPT-4 to output the corresponding executable code.
+1# Given a task you have to provide Python code for executing the task
+2
+3# importing available functions
+4
+5from spatial_directions import top, front, back, left, right
+6
+7from primitives import assign_head, move_head, keep_at_head
+8# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
+9
+10from helpers import find_size, filter
+11from structures import row, column, tower
+12
+13
+14#function signature of the imported functions
+15# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
+16filter(color, shape)
+17
+18# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
+19find_size(struct_name = str_name, objects = ObjSet)
+20
+21# assigns the head to the location of the object
+22assign_head(at_obj_loc)
+23
+24# moves the head in the given dir
+25move_head(dir)
+26
+27# keeps the object obj at the head
+28keep_at_head(obj)
+29
+30
+31#Examples:
+32#Instruction: Move the green block to the left of the red dice
+33assign_head(at_obj_loc = filter(red, dice))
+34move_head(left)
+35keep_at_head(obj = filter(green, cube))
+36
+37# Instruction: Find the size of the tower made of yellow legos
+38find_size(struct_name = tower, objects = filter(yellow, lego))
+39
+40#Instruction: Find the size of the row made of orange cubes
+41find_size(struct_name = row, objects = filter(orange, cube))
+42
+43# Instruction: Find the size of the column made of cyan cubes
+44find_size(struct_name = column, objects = filter(cyan, cube))
+45
+46# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
+47assign_head(at_obj_loc = filter(red, dice))
+48move_head(left)
+49keep_at_head(obj = filter(green, cube))
+50assign_head(at_obj_loc = filter(green, cube))
+51move_head(top)
+52keep_at_head(obj = filter(yellow, cube))
+53
+54# Instruction: Construct a row of green legos of length 3 to the right of the blue block
+55assign_head(at_obj_loc = filter(blue, block))
+56move_head(right)
+57row(length = 3, objects = filter(green, legos))
+58
+59
+60# Instruction: Construct a tower of size 3 using red cubes
+61tower(height = 3, objects = filter(red, cube))
+62
+63# Instruction: Construct a row of size 5 using blue legos
+64row(length = 5, objects = filter(blue, lego))
+65
+66# Instruction: Construct a column of size 6 using green die
+67column(length = 6, objects = filter(green, dice))
+68
+69# Instruction: Place 3 green blocks so that one block is to the right of the other
+70green_blocks = filter(green, block)
+71keep_at_head(green_blocks)
+72move_head(right)
+73keep_at_head(green_blocks)
+74move_head(right)
+75keep_at_head(green_blocks)
+76
+77# Instruction: Place 3 red legos on top of one another
+78red_legos = filter(red, lego)
+79keep_at_head(red_legos)
+80move_head(top)
+81keep_at_head(red_legos)
+82move_head(top)
+83keep_at_head(red_legos)
+84
+85
+86# CURRENT TASK
+87# Instruction: Construct tower of white cubes to the same height as
+88# existing tower of green die
+89
+90# GPT-4s output
+91#First, we have to find the height of the tower of green dice,
+92#then construct a tower of white cubes of the same size
+93
+94tower_size = find_size(struct_name = tower, objects = filter(green, die))
+95tower(height = tower_size , objects = filter(white, cube))
+Listing 9: Prompting method for the task of constructing tower of white cubes to the same height as existing tower of green die
+1# Given a task you have to provide Python code for executing the task
+2
+3# importing available functions
+4from spatial_directions import top, front, back, left, right
+5from primitives import assign_head, move_head, keep_at_head
+6# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
+7from helpers import find_size, filter
+8from structures import row, column, tower
+9#function signature of the imported functions
+10# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
+11filter(color, shape)
+12# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
+13find_size(struct_name = str_name, objects = ObjSet)
+14# assigns the head to the location of the object
+15assign_head(at_obj_loc)
+16# moves the head in the given dir
+17move_head(dir)
+18# keeps the object obj at the head
+19keep_at_head(obj)
+20
+21#Examples:
+22#Instruction: Move the green block to the left of the red dice
+23assign_head(at_obj_loc = filter(red, dice))
+24move_head(left)
+25keep_at_head(obj = filter(green, cube))
+26
+27# Instruction: Find the size of the tower made of yellow legos
+28find_size(struct_name = tower, objects = filter(yellow, lego))
+29
+30#Instruction: Find the size of the row made of orange cubes
+31find_size(struct_name = row, objects = filter(orange, cube))
+32
+33# Instruction: Find the size of the column made of cyan cubes
+34find_size(struct_name = column, objects = filter(cyan, cube))
+35
+36# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
+37assign_head(at_obj_loc = filter(red, dice))
+38move_head(left)
+39keep_at_head(obj = filter(green, cube))
+40assign_head(at_obj_loc = filter(green, cube))
+41move_head(top)
+42keep_at_head(obj = filter(yellow, cube))
+43
+44# Instruction: Construct a row of green legos of length 3 to the right of the blue block
+45assign_head(at_obj_loc = filter(blue, block))
+46move_head(right)
+47row(length = 3, objects = filter(green, legos))
+48
+49
+50# Instruction: Construct a tower of size 3 using red cubes
+51tower(height = 3, objects = filter(red, cube))
+52
+53# Instruction: Construct a row of size 5 using blue legos
+54row(length = 5, objects = filter(blue, lego))
+55
+56# Instruction: Construct a column of size 6 using green die
+57column(length = 6, objects = filter(green, dice))
+58
+59# Instruction: Place 3 green blocks so that one block is to the right of the other
+60green_blocks = filter(green, block)
+61keep_at_head(green_blocks)
+62move_head(right)
+63keep_at_head(green_blocks)
+64move_head(right)
+65keep_at_head(green_blocks)
+66
+67# Instruction: Place 3 red legos on top of one another
+68red_legos = filter(red, lego)
+69keep_at_head(red_legos)
+70move_head(top)
+71keep_at_head(red_legos)
+72move_head(top)
+73keep_at_head(red_legos)
+74
+75
+76# CURRENT TASK
+77# Instruction: Construct a tower of height 6 using red and blue blocks that are alternating
+78
+79# GPT-4s output
+80# Python code:
+81
+82# Define the red and blue blocks
+83red_blocks = filter(’red’, ’block’)
+84blue_blocks = filter(’blue’, ’block’)
+85
+86# Start at the bottom and alternate building the tower
+87for i in range(6):
+88 if i%2 == 0: # if the stack position is even
+89 keep_at_head(red_blocks)
+90 else: # if the stack position is odd
+91 keep_at_head(blue_blocks)
+92 if i != 5: # if not at the top of the tower
+93 move_head(top)
+Listing 10: Prompting method for the task of constructing tower of alternating red and blue cubes
+To find the size of a given structure in the given scene we define the function find_size, which takes the name of structure, all the objects in the initial scene, mask of the objects (a distribution over the objects based on the attributes), and the initial state.
+(we assume that this function has access to the semantics of all the concepts learnt so far, through a transition function).
+Algorithm  LABEL:alg:find-size gives the pseudocode for the function find_structure.
+Below we provide a brief explanation for it.
+  1. 1.
+    First we assign our head to every block in the available blocks (line 6)
+  2. 2.
+    Then we begin constructing/visualizing the corresponding structure from that block beginning with a size of 1. (line 7)
+  3. 3.
+    For each structure created/visualized we compare the blocks moved for the structure creation with corresponding blocks originally present in the scene, and perform a matching between these blocks and a subset of the blocks originally present (line 11-26).
+  4. 4.
+    If we are able to find a mapping for each moved block, such that each mapped pair has an IoU greater than a threshold, we increase the next potential size to test by 1 (line 26-27).
+  5. 5.
+    The final size is the size corresponding to 2nd last iteration, before termination (line 28).
+  6. 6.
+    We return the maximum of all the possible structures that are found (line 34)
+D.4 Details on MCTS Variants for Plan Search
+Here we provide the details for 3 different plan search methods, that search over the space $\mathcal{A}=\mathcal{A}_{c}\cup\mathcal{A}_{p}$
+  • MCTS+L+P: This is the approach that we describe in section 4.2.
+For every concept say Tower $\in\mathcal{L}$ we have a corresponding set of macro action say Make_Tower(3, objects) (L).
+Further we use a neural pruner $\pi_{\text{neural}}$ that outputs a primitive action $a_{p}^{*}$ (given current state and next expected state).
+We only consider the actions $\mathcal{A}_{c}\cup\{a_{p}^{*}\}$ during our search from the given state.
+This helps to reduce the effective branching factor and allows to search longer length plans within the same computational budget.
+  • MCTS+L-P: Here we do not use the reactive policy, therefore the branching factor for every node becomes $\mathcal{A}$ = $\mathcal{A}_{c}\cup\mathcal{A}_{p}$.
+  • MCTS-L+P: We only search among the space of primitive actions i.e $\mathcal{A}_{p}$.
+Given a state $\tilde{s}_{t}$ and corresponding next expected state $s_{t+1}$ we greedily pick the action $a_{p}^{*}=\pi_{\text{neural}}(\tilde{s}_{t},s_{t+1})$.
+This method is much more faster than the previous 2 methods as there is no explicit search.
+However the corresponding policy is trained only to output an action $a\in\mathcal{A}_{p}$ and lacks the ability to output modular plans composed of macro actions such as Make_Tower(3, objects) $\in\mathcal{A}_{c}$ (the action space $\mathcal{A}_{c}$ is increasing with time and the architecture of network needs to be changed accordingly).
+As a result the plans found are not modular and difficult to generalize.
+Further, the reactive policy is not trained to output reset_head(), store_head() as additional annotated data is required in order to train a classifier over them.
+This further decreases the space of grounded plans (and therefore corresponding generic programs) such a policy can represent.
+Training a reactive policy that can handle actions such as reset_head() and an action space $\mathcal{A}_{c}$ that grows with time is part of future work.
+D.5 Goal-conditioned Planning with Learnt Concepts
+Why is it difficult to hand encode a PDDL for our domain?
+Most of the PDDL description of blocks world assume actions involving only the spatial relation onTop, which limits their applicability to describing different structures like row that need spatial relations like onRight.
+Further a single action might lead to varied effects/post-conditions based on the initial state.
+19 gives 2 example of the same action moveOnTop(A, B) which would end up giving adding different number of spatial relations.
+Figure 19: Difficult to encode post-conditions. Illustration of a domain where encoding a PDDL for direct planning is challenging.
+Approach Overview.
+Given an instruction $\Lambda$ = “Construct a staircase of magenta die having 3 steps”, we first convert it into corresponding grounded task sketch $H_{S}^{*}$ = staircase(3, [3, 2, 1, $\cdots$]).
+Executing the corresponding program of staircase (by getting the semantics from the library $\mathcal{L}$) on the desired objects we get the expected final scene $S_{f}^{\prime}$ in bounding box space.
+The initial scene $S_{i}$ and expected final scene $S_{f}^{\prime}$ are converted into scene graph $SG_{i}^{\prime}$ and $SG_{f}^{\prime}$ (described in  19).
+The relations between the task relevant objects in $SG_{f}^{\prime}$ act as propositions/relations for goal check and the initial scene graph act as the initial state.
+Then a neuro-symbolic planner is used to obtain the optimal plan from the start state to a state that satisfies the goal.
+Below we also detail different aspects of the approach.
+Scene-graph Extraction.
+LABEL:scene-graph-algo gives the algorithm used for generating scene graph from a given scene (set of bounding boxes).
+Suppose we need to check whether there exists a relation of the form $(i,j,direction)$ i.e.
+block $i$ is in the direction $direction$ of block $j$, in a given scene.
+We first initialize the head at the position/bounding-box of block $j$ (line 7).
+Then we move the head in direction $direction$ (line 9).
+We claim that the relationship would exists if bounding box for block $i$ has IoU $>$ 0.75 with the predicted_head (line 10-12).
+Pre-conditions. We define the following 2 preconditions (and learn their grounding) in order to ensure that the generated plans are physically possible.
+  1. 1.
+    is-clear(blk, dir): We need to check whether a block blk has some free space in direction dir.
+For this we simply move the head in the direction dir with respect to the block blk.
+If the resulting position of head has 0 overlap with bounding boxes of all the other objects the predicate is True otherwise False.
+  2. 2.
+    will-not-be-floating(pred_loc): We need to check whether the resultant/predicted location of an object on taking an action would be dynamically stable or not.
+The location would be stable if either it is on top of some already placed object or it is on the table surface.
+The former can be checked through the resultant scene graph itself (that is obtained by applying the algorithm  LABEL:scene-graph-algo), while for the later we train an on-table classifier.
+This would take as input a bounding box and predict whether the box is on the table or not.
+For training this classifier we use the dataset of pretraining phase.
+The blockwise positive and negative sample annotation can be done automatically by giving GPT-4 the corresponding scene graph and then querying which objects are on the table and which aren’t.
+20 gives an example.
+(Though we have not taken this approach for the complete dataset of 5k samples due to high cost).
+Figure 20: Method to get annotation for training on-table classifier.
+Actions. We define the following two types actions
+  1. 1.
+    place-random(blk): To place a block at a random free position on the table.
+For this we train a generative model (VAE) which learns the underlying distribution of bounding boxes for all the blocks that lie on the table.
+Given a scene we would sample position (bounding box) from this until we get a position that is not overlapping with the existing blocks in the scene.
+For training the VAE we assume that for every demonstration in the pretraining data, the first scene has all the objects randomly placed on the table (we could have also used the positive examples used for training on-table classifier).
+  2. 2.
+    move(rel, blk1, blk2): This action corresponds to moving the blk1 in the direction rel of blk2, resulting in the addition of a relation (rel, blk1, blk2) in the set of spatial relations.
+This action is defined as a sequential composition of the actions assign_head(blk2), move_head(rel), keep_at_head(blk1) (blk1 is a one hot tensor for the corresponding block).
+Techniques and heuristic for efficient planning.
+Since the action space for the planner could be $o(n^{2})$, where n is the number of objects we adopt the following techniques to make planning scalable/efficient:
+  1. 1.
+    Heuristic - We define the heuristic value h(s) for a state s, as the number of relations that are present in the goal but are absent in the scene graph corresponding to the state s.
+Even though this heuristic is not admissible (as it may over overestimate the cost to goal), it was found to work optimally in most of the cases.
+  2. 2.
+    Greedy-pruning - We assume that all the actions resulting in states with higher or same heuristic value would be of the form place-random(blk).
+This means among the actions of the form move(rel, blk1, blk2) we only select those that lead to states with decreased heuristic value.
+  3. 3.
+    Relevant-object-set - Suppose O is the set of objects that are part of atleast one of the predicate in goal.
+We define O’ as the transitive closure of O with respect to the relation $Related$ in the initial state $s_{i}$, where $SG(s_{i})$ is scene graph for the initial state
+    $Related(a,b,s_{i})\iff\exists dir((dir,a,b)\in SG(s_{i})\vee(dir,b,a)\in SG(s_{i}))\;$ (9)
+    We assume O’ is the relevant set of object for completing the task and actions that move any other object should not be taken.
+Appendix E Broader Impact
+This work creates foundational knowledge in understanding human-like spatial abstractions.
+This work contributes towards the development of explainable and interpret-able learning architectures that may eventually contribute towards the development of embodied agents collaborating with and assisting humans in performing tasks.
+No negative impact of this work is envisioned.
+Appendix F Hyperparameters, Architecture details and Ground Truth Concepts
+F.1 Architecture for neural modules
+Action Simulator:
+1import torch.nn as nn
+2
+3class ActionSimulatorNetwork(nn.Module):
+4 def __init__(self, bbox_mode, hidden_size = 256):
+5 super(ActionSimulatorNetwork, self).__init__()
+6 self.bbox_mode = bbox_mode
+7 self.hidden_size = hidden_size
+8
+9 self.action_semantics_encoder = nn.Sequential(
+10 nn.Linear(5, hidden_size),
+11 nn.ReLU(),
+12 nn.Linear(hidden_size, hidden_size),
+13 nn.ReLU()
+14 )
+15 self.argument_encoder = nn.Sequential(
+16 nn.Linear(5, hidden_size),
+17 nn.ReLU(),
+18 nn.Linear(hidden_size, hidden_size),
+19 nn.ReLU()
+20 )
+21 self.decoder = nn.Sequential(
+22 nn.Linear(hidden_size, hidden_size),
+23 nn.ReLU(),
+24 nn.Linear(hidden_size, 5),
+25 nn.Tanh()
+26 )
+Listing 11: Action Simulator Network in PyTorch
+Reactive Policy($\pi_{\text{neural}}$):
+1import torch.nn as nn
+2
+3class NeuralSearch(nn.Module):
+4 def __init__(self, action_space=6):
+5 super(NeuralSearch, self).__init__()
+6 self.action_space = action_space
+7 self.fc1 = nn.Linear(10, 256)
+8 # self.bn1 = nn.BatchNorm1d(256)
+9 self.bn1 = nn.Identity()
+10 self.fc2 = nn.Linear(256, 256)
+11 # self.bn2 = nn.BatchNorm1d(256)
+12 self.bn2 = nn.Identity()
+13 self.fc3 = nn.Linear(256, 256)
+14 # self.bn3 = nn.BatchNorm1d(256)
+15 self.bn3 = nn.Identity()
+16 self.fc4 = nn.Linear(256, action_space)
+Listing 12: Neural Search in PyTorch
+Random Position predictor (for grounding of place-random(blk)):
+1import torch.nn as nn
+2
+3class VAE(nn.Module):
+4 def __init__(self, input_dim, latent_dim):
+5 super(VAE, self).__init__()
+6 self.input_dim = input_dim
+7 self.latent_dim = latent_dim
+8 # Encoder
+9 self.fc1 = nn.Linear(input_dim, 512)
+10 self.bn1 = nn.BatchNorm1d(512)
+11 self.fc2 = nn.Linear(512, 512)
+12 self.bn2 = nn.BatchNorm1d(512)
+13 self.fc3 = nn.Linear(512, 512)
+14 self.bn3 = nn.BatchNorm1d(512)
+15 self.fc4 = nn.Linear(512, 512)
+16 self.bn4 = nn.BatchNorm1d(512)
+17 self.fc51 = nn.Linear(512, latent_dim) # Mean of the latent space
+18 self.fc52 = nn.Linear(512, latent_dim) # Log-variance of the latent space (log-var for numerical stability)
+19
+20 # Decoder
+21 self.fc5 = nn.Linear(latent_dim, 512)
+22 self.bn5 = nn.BatchNorm1d(512)
+23 self.fc6 = nn.Linear(512, 512)
+24 self.bn6 = nn.BatchNorm1d(512)
+25 self.fc7 = nn.Linear(512, 512)
+26 self.bn7 = nn.BatchNorm1d(512)
+27 self.fc8 = nn.Linear(512, 512)
+28 self.bn8 = nn.BatchNorm1d(512)
+29 self.fc9 = nn.Linear(512, input_dim)
+Listing 13: VAE in PyTorch
+On-table classifier (for grounding of will-not-be-floating(pred_loc):
+1import torch.nn as nn
+2
+3class TableClassifier(nn.Module):
+4 def __init__(self):
+5 super(TableClassifier, self).__init__()
+6 self.fc1 = nn.Linear(5, 16)
+7 self.bn1 = nn.BatchNorm1d(16)
+8 self.fc2 = nn.Linear(16, 16)
+9 self.bn2 = nn.BatchNorm1d(16)
+10 self.fc3 = nn.Linear(16, 16)
+11 self.bn3 = nn.BatchNorm1d(16)
+12 self.fc4 = nn.Linear(16, 1)
+13 self.bn4 = nn.BatchNorm1d(1)
+14 self.sigmoid = nn.Sigmoid()
+Listing 14: Table Classifier in PyTorch
+F.2 Hyperparameters used in experiment
+As indicated in  A.3 for the purpose of generalization through multiple candidate plans (from 1 demonstration) we chose the top-k plans (as measured by overall IoU achieved).
+The k chosen for all our experiments involving MCTS was 5.
+(The performance of our best approach was found to be the same for k=5 to 20).
+For every plan we obtain 3 programs from GPT-4 by re-prompting it 3 times with the same input prompt (with temperature > 0).
+From the pool of these 3*k programs we chose the one with highest IoU reward by running each of them on the given demonstration.
+The discount factor kept for our search is $\gamma$ = 0.95, and unless explicitly specified the number of expansions steps used = 5000.
+F.3 Ground-Truth Inductive Concepts
 1######
@@
 136######
-Listing 1: Definition of inductive concepts
-Table 5: Symbols and Semantics Signature and semantics for the primitive concepts and operations that are used in the construction of the programs used to express inductive spatial concepts.
-Function Signature Semantics
-filter (VisualConcept, ObjSet) $\rightarrow$ ObjSet Returns the objects that contain the VisualConcept
-move_head (Head, Dir) $\rightarrow$ Head Moves the head to the given direction (May or may not take input/return the head, based on a flag)
-assign_head a.k.a move_head(overloaded) (Head, ObjIdx) $\rightarrow$ Head Given the index/one-hot representation for an object, it moves the head to the position corresponding to that object
-keep_at_head (ObjSet, Head) $\rightarrow$ None Keeps the argmax of ObjSet at the head
-reset_head None $\rightarrow$ Head Sets the head to the top position of stack and pops this position from the stack as well
-store_head Head $\rightarrow$ None Pushes the current position of head into the stack
-Table 6: Symbolic representation. The table lists the type definitions used in the implementation of SPG programs.
-Defined Types Python Type Usage
-IntArg int Argument for the structures that defines the size (height, length etc)
-Obj torch.Tensor One-hot vector whose non-zero index represents the selected objects
-ObjSet torch.tensor Probability mask over the selected objects
-Dir string Primitive directions like left, right, front, top, etc
-ConceptName string name of the visual, action or inductive concept
-Head torch.Tensor Bounding box with depth. 3D cuboidal space.
-A.3 Curriculum Learning
-We follow a curriculum approach where the visual concepts are trained first from simpler linguistically-described demonstrations ($t_{0}$ in figure  9).
-This is followed by learning of action concepts through sequentially composed pick and place tasks.
-Its essential to use such long range sequential instructions in order to ensure that the semantics of action concepts are learnt for placement of objects at a height much above the tabletop ($t_{1}$ in figure  9).
-After the pre-training phase, the agent can continually learn new inductive concepts and visual attributes.
-($t_{2},t_{3}$ in figure  9)
-Figure 9: Continual learning through curriculum: Using simple pick and place demonstrations we learn visual attributes such as blue cube, yellow lego ($t_{0}$).
-Using long range instructions which are sequentially concatenated descriptions of pick and place tasks we train our action concepts such as left, right, top ($t_{1}$).
-After pre-training the agent can perform continual learning of concepts such as learning generalized representation for tower ($t_{2}$).
-Because of disentangled representation of neural and symbolic concepts, interspersed learning of new visual attributes such as chocolate color are also possible through few demonstrations of structure creation($t_{3}$).
-A.4 Details for Plan-Search and Generalization
-Plan-Search: We present additional details on the MCTS procedure for searching for a plan conditioned on a program signature and guided by the demonstration.
-In order to incorporate the objective of searching for physically realizeable plans and facilitating generation via re-use of concepts, the following conceptual changes are incorporated in the standard MCTS procedure [sutton2018reinforcement].
-Modularity (MCTS + L): We want to allow learning of novel inductive concepts in terms of existing ones.
-This would ensure that the plan $H_{P}^{*}$ corresponding to a given demonstration is concise and can be easily generalized to the $H_{G}^{*}$.
-C.3 in appendix give example of two plans for the structure Pyramid one which is modular and can be successfully generalized by GPT-4, other for which GPT-4 fails in generalization due to lack of modularity.
-This can be seen as a form of regularization in terms of the length of concept description, by making the prior $P(H)\propto|H|^{-\alpha}$ (where $\alpha$ > 0) in equation (2)
-$H^{*}=\arg\min_{H\in\mathcal{H}}\left[\mathtt{Loss}(\{S_{1}..S_{g}\},H(\Lambda))+\alpha\log|H|\right]$ (4)
-In order to allow modular learning of programs, for every inductive concept already stored in the library we define corresponding action instantiations which can be a potential candidate actions during our search.
-As an example, for the concept Tower we have one of the action instantiation as Make_Tower(3, objects) which would be the action of constructing desired tower.
-This can be visualized as a compound/macro-action which is composed of primitive actions keep_at_head(objects), move_head(‘top’).
-We define $\mathcal{A}_{c}$ as the space of such compound actions, and $\mathcal{A}_{p}$ as the space of primitive action/function consisting reset_head(), move_head(direction), keep_at_head(objects), store_head().
-Realization of equation (6) (increased preference of macro-actions over primitive ones) is done through discounted IoU rewards during our search (Make_Tower(3) would have a reward of 1+1+1, as compared to 1 + $\gamma^{2}$(1) + $\gamma^{4}$(1) for a sequence of 3 (keep_at_head(objects), move_head(’top’)0)).
-We refer to MCTS using concept instantiations from $\mathcal{L}$ as macro actions in search as MCTS+L.
-Scalablility (MCTS+P): as more concepts are added to the library $\mathcal{L}$ the action space of our search $\mathcal{A}$ (specifically $\mathcal{A}_{c}$, the space of compound concepts) increases, therefore we want to prune the search space effectively.
-For this during the pre-training phase we train a reactive policy $\pi_{neural}$ which given the current state, $\tilde{s_{t}}$ and the next expected state ${s_{t+1}}$ (part of the demonstration) would output one of the primitive action, $a_{t}^{*}\in\mathcal{A}_{p}$ where $\mathcal{A}_{p}$ is the primitive action space, i.e.
-$a_{t}^{*}=\pi_{neural}(a_{t}|\tilde{s_{t}},{s_{t+1}}),\;\;a_{t}\in\mathcal{A}_{p}$ Note that while expanding our search tree we only search among the space of compound actions $\mathcal{A}_{c}$ and the action $a_{t}^{*}$, thereby reducing the branching factor of search from $|\mathcal{A}_{c}\cup\mathcal{A}_{p}|$ to $|\mathcal{A}_{c}|+1$.
-We refer to MCTS using neural pruning as MCTS+P.
-Therefore our MCTS algorithm is modular through hierarchical composition of learnt concepts and efficient through pruning of action space and is referred to as MCTS+L+P.
-Modifications to the Simulation and Reward Back propagation Steps: Next, we outline the modifications in the simulation and the reward back propagation steps of the standard MCTS algorithm for our setting.
-During program search we assume access of intermediate scenes in the demonstration.
-This allows us to provide intermediate rewards that can guide the search well.
-We observed that making the following changes in simulation and back propagation step increased the efficiency of our search procedure.
-  • Simulation: Rather than performing Monte Carlo simulations at each newly expanded leaf node (to estimate its value) we completely avoid these simulation steps.
-This was motivated by the fact that our reward is not completely sparse and the intermediate IoU rewards for each object we place allow us to guide the search effectively.
-  • Back propagation: We perform off policy Q-learning updates during back propagation similar to one indicated by [complexmctsbackup] :
-    $V(s_{t})=max_{a\in\mathcal{A}}Q(s_{t},a).$ (5)
-    $\tau(s_{t},a)=s_{t+1}$ (6)
-    $Q(s_{t},a)=r_{t}+\gamma V(s_{t+1})$ (7)
-Generalization: GIven multiple equal length plans for a given demonstration, we seek to recover a plan one that can be easily generalized by the LLM.
-6 shows a plan which could be correctly abstracted out into a generic program by GPT-4.
-Whereas  7 shows another plan with similar semantics, for which GPT-4 is unable to correctly find the generalized program (Note that row or column of size 1 is equivalent to keep_at_head).
-We tackle this problem in the following manner.
-  1. 1.
-    Rather than getting a single plan from the plan search we get the top k plans $\{H_{P,i}\}_{i=1}^{i=k}$.
-In order to get these top k plans we expand the complete tree (based on UCB criteria) starting from the root node corresponding to the initial state, till a predefined budget of expansions.
-Then we select the top k paths(potential plans) from the root node to all the leaf nodes (where the top k ones are those that give the highest accumulated IoU reward with respect to the given demonstration).
-  2. 2.
-    Later we abstract out each of these k plans into corresponding generalized programs, $\{H_{G,i}\}_{i=1}^{i=k}$ using GPT-4.
-We again run each of these programs on the given demonstration and then choose the one which gives the highest IoU reward (resolving ties based on predefined order).
-Note that some program $H_{G,i}$ upon execution may result in a plan $\tilde{H}_{P,i}$ different from the original plan ${H}_{P,i}$ using which it was generalized.
-This can be attributed to potential errors in GPT-4s program generalization process.
-A.5 Additional Details: Learning with Increasing Number of Demonstrations
-Given k demonstrations for a novel inductive concept, we independently find k task sketch $\{H_{S,i}^{*}\}_{i=1}^{i=k}$ and grounded plans $\{H_{P,i}^{*}\}_{i=1}^{i=k}$.
-During the generalization phase we give these k pair of task-sketch and corresponding plans to GPT-4 and ask into infer a single abstraction over them.
-C.5 in appendix gives a concrete example.
-Equation for generalize step (getting $H_{G}^{*}$ from $H_{P}^{*},H_{S}^{*}$) can be modified as follows.
-$H_{G}^{*}\leftarrow Generalize(H_{G}\,|\,{\{H_{P,i}^{*},\,H_{S,i}^{*}\}}_{i=1}^{i=k};\,{\theta_{G}}),\;\;H_{G}\in\mathcal{H}^{\mathcal{G}}$ (8)
-Appendix B Additional details regarding datasets
-Figure  10 demonstrates the kind of inductive concepts for which we want to learn generic (i.e instance agnostic) representations.
-Figure 10: Illustration of the inductive concepts.
-Dataset for Pre-training: We use 5k examples of constructing twin-towers i.e.
-2 towers adjacent to each other, for learning semantics of move_head(dir), a basic set of visual attributes, reactive policy $\pi_{\text{neural}}$, and neural modules required for grounded planning.
-The twin towers allow us to learn various action semantics for all possible configurations of blocks in 3D-space (and not being limited to blocks placed directly on table top surface).
-Since we are not aware of the underlying semantics of tower during pre-training phase the corresponding natural language instruction consists of step by step pick and place actions.
-11 gives example demonstrations from this dataset.
-Figure 11: Example from pre-training dataset.
-Dataset for Inductive Structures: We learn a variety of structures which we have divided into Simple and Complex structures.
-A structure is considered complex if it can be expressed as an inductive composition of simpler structures.
-As an example, we can express a staircase to be a composition of towers of increasing height.
-The structures are listed in the Table 7.
-Figure  12 shows the hierarchical relationship among these structures in the form of a DAG (directed acyclic graph).
-A.2 gives the ground truth program representations for each structure.
-Table 7: Structure Types. Examples of simple and complex structures considered in this work for the robot to construct.
-Simple Structures Complex Structures
-Row, Column, Tower X (cross-shape), Staircase
-Inverted-Row, Inverted-Column Inverted-Staircase, Pyramid
-Diagonal-45, Diagonal-135 Arch-Bridge, Boundary
-Diagonal-225, Diagonal-315
-Figure 12: Hierarchy of the structures/programs. This diagram shows the hiearchical nature of the structures in our dataset. MH is abbrevation for move_head and KH is abbreviation for keep_at_head
-Appendix C Prompting Strategy and Examples
-This section gives various prompting examples for our approach and baselines, along with examples motivating particular design decisions in our approach.
-C.1 Prompt Example for Task Sketch Generation Stage (Sketch)
-In order to get a program representation (high level task sketch) of the given natural language instruction, we prompt GPT-4 with few shot examples in a manner similar to [code-as-policies].
-Code segment  2 gives an example of getting the task sketch given the demonstration for constructing a staircase.
-We first import the available primitive operators and functions and also give examples in order to demonstrate the signature of the available primitives(line 1-8).
-Then we give incontext examples of how to parse various natural language instructions in a program representation(line 10-14).
-We append to this prompt the instruction for current task(line 16-17).
-1# importing the available functions
-2from visual_operators import filter
-3from inductive_operators import get_parameters, find_structure
-4
-5# function signature of the imported functions
-6filter(color, cube) # filter the objects that are cubes and color ..
-7get_parameters(structure) # parameters of the structure ...
-8find_structure(type, description) # finding structure of given type, description
-9
-10# examples:
-11# instruction: Find the tower with green cubes
-12find_structure(type = tower, description = filter(green, cube))
-13# instruction: Construct a tower of height 3 with yellow cubes
-14Tower(height = 3, objects = filter(yellow, cubes))
-15
-16# current task: Construct a staircase of 4 steps using cyan legos
-17
-18# (GPT-4s output)
-19Staircase(steps = 4, filter(cyan, legos))
-Listing 2: Task Sketch Generation Using GPT-4 (Sketch)
-C.2 Prompt Example for Generalizing a sequence of actions/plan to a general program (Generalize)
-Code segment  3 give an example of getting the general Python program from the plan found using MCTS.
-We first provide a base prompt giving details to GPT-4 about the desired task (line 1-2).
-Then we give the input arguments and the corresponding output/plan for a given demonstration (line 3-5).
-We expect the GPT-4 to output the final Python program (line 7-10)staircase (line 6-8).
-1# Write a general python code which on the given input produces the desired output, do not output anything other than the function description.
-2
-3# input: n = 4, objects = ObjSet
-4# output: tower(1, ObjSet), move_head(‘right’), tower(2, ObjSet), move_head(‘right’), tower(3, ObjSet), move_head(‘right’), tower(4, ObjSet)
-5
-6# Program (GPT-4s output)
-7def staircase(n, objects):
-8 for i in range(n):
-9 tower(i+1, objects)
-10 move_head(’right’)
-Listing 3: Plan to Program using GPT-4 (Generalize)
-C.3 Benefit of Estimating Modular/Smaller Plans
-The below examples demonstrate the benefit of learning new inductive concepts in terms of already acquired inductive concepts (more modular representation).
-The program  4 is obtained through generalization of plan $H_{p}$ that represents pyramid in terms of rows of decreasing size.
-The generated program generalizes to pyramid of different height.
-1# input: n = 3, objects = ObjSet
-2# output: row(5, ObjSet), move_head(’right’), move_head(’top’), row(3, ObjSet), move_head(’right’), move_head(’top’), row(1, ObjSet)
-3
-4# Program (GPT-4s output)
-5def pyramid(n, objects)
-6 for i in range(n, 0, -1):
-7 # Calculate the number of objects in the current row
-8 row_count = 2 * i - 1
-9
-10 # Perform actions for the current layer
-11 row(row_count, objects)
-12 if i > 1: # No need to move head after the last row
-13 move_head(’right’)
-14 move_head(’top’)
-Listing 4: Pyramid in terms of rows
-On the other hand program  5 which is obtained in an attempt to generalize the plan expressed in terms of primitives is incorrect.
-1# input: n = 3, objects = ObjSet
-2# output: store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’top’), move_head(’right’), store_head(), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), move_head(’right’), keep_at_head(ObjSet), reset_head(), move_head(’right’), move_head(’top), keep_at_head(ObjSet)
-3
-4# Program (GPT-4s output)
-5def pyramid(n, objects)
-6# Store initial position
-7 store_head()
-8
-9 # Initial sequence of placing and moving right
-10 for _ in range(n):
-11 keep_at_head(objects)
-12 move_head(’right’)
-13
-14 # Reset and move to top right position
-15 reset_head()
-16 move_head(’top’)
-17 move_head(’right’)
-18
-19 # Store again and repeat placing and moving right
-20 store_head()
-21 for _ in range(n):
-22 keep_at_head(objects)
-23 move_head(’right’)
-24
-25 # Reset and adjust final position
-26 reset_head()
-27 move_head(’right’)
-28 move_head(’top’)
-29 keep_at_head(objects)
-Listing 5: Pyramid in terms of primitive actions
-C.4 Comparing Same-length Plans in terms of Generalizability
-Below we illustrate plan to program generalization for 2 different plans which are correct and have same length, but LLMs find it difficult to generalize the second one as compared to first.
-(Note that row and column of size 1 are equivalent to keep_at_head)
-1# input: n = 3, objects = ObjSet
-2# output: keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet), move_head(’top’), keep_at_head(ObjSet)
-3
-4
-5# Program
-6def tower(n, objects):
-7 for _ in range(n):
-8 keep_at_head(objects)
-9 move_head(’top’)
-Listing 6: Plan for tower that can be easily generalized (correct generalization)
-1# input: n = 3, objects = ObjSet
-2# output: row(1, ObjSet), move_head(’top’), keep_at_head(1, ObjSet), move_head(’top’), column(1, ObjSet)
-3
-4
-5# Program
-6def tower(n, objects):
-7 for i in range(1, n + 1):
-8 row(i, objects)
-9 move_head(’top’)
-10 keep_at_head(objects)
-11 move_head(’top’)
-12 if i < n:
-13 column(i, objects)
-14 move_head(’top’)
-Listing 7: Plan for tower that is difficult to generalize (Incorrect generalization)
-C.5 Generalizing via Multiple Demonstrations
-Given multiple demonstrations we independently find task sketch and corresponding grounded plans for each demonstration.
-These are further given to GPT-4 for generalization.
-Code segment  8 gives an example of getting a single Python program from multiple demonstrations.
-Note that we explicitly prompt the LLM that some of the grounded plans might be incorrect (which may lead to more robust generalization in case of noisy demonstrations).
-1
-2# Function Call: wor(height = 3, objects = ObjSet_1)
-3# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
-4# Function Call: wor(height = 3, objects = ObjSet_1)
-5# Execution: keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
-6# Function Call: wor(height = 3, objects = ObjSet_1)
-7# Execution: column(size=1, obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1), move_head(dir = right), keep_at_head(obj = ObjSet_1),
-8
-9#Write the function definition, which generalizes the above executions. Note that some of the executions can be partially wrong.
-10‘‘‘python
-11def wor(height, objects):
-12‘‘‘
-13
-14GPT-4s Output .....
-Listing 8: Generalizing through multiple plans
-C.6 Prompt Examples for Learning Programs using LLM/VLM Models
-Below we describe the prompting methodologies for learning programs through LLM/VLM models.
-Note that although the prompt examples described below are for the case of learning novel structure from 1 demonstration, we use 3 demonstration per novel structure in our main results (for both our approach and LLM/VLM baseline).
-LLM/GPT-4 Code segment 9 depicts our prompting methodology given a demonstration for a new concept tower.
-For this baseline we aim to check demonstration following and spatial reasoning abilities of LLMs (GPT-4).
-We provide supervision of the intermediate scenes by using tokenized spatial relations between objects in the scene (Given in the form of Scene = [right(1, 0) …]).
-We further assume that only those objects that are required to perform the task are present in the scene (no distractor objects).
-For every structure (that needs to be learned at time t) we give LLM a prompt providing in-context example on how to generalize (line 19-35), the set of primitive operators (line 4) available and the set of structures learnt/present in library (till time t-1) (line 5-18).
-Finally we append to this prompt the expected declaration (arguments and keywords arguments) of the inductive concept that is to be learnt along with the spatial relations for each scene of the given demonstration (36-51).
-Note that we assume absence of distractor objects for this baseline.
-1# Consider a block world domain ......
-Given a structure creation task along with intermediate scnes complete a general Python function for it.
-The function should be in terms of primitive operators and already learnt structures that are present in the program library.
-Enclose the function within backtick (‘‘‘)
-2
-3primitive_operators = [keep_at_head, move_head ..]
-4# this would be our program library
-5learnt_structures = {
-6 "row": {
-7 "program_tree":
-8 ’’
-9 def row(size, objects):
-10 for i in range(size):
-11 keep_at_head(obj = objects)
-12 move_head(dir = right’)
-13 ’’,
-14 },
-15
-16 ......
-17}
-18# the example task
-19Example task:- Place all the objects to the right of each other.
-20Final state :- [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
-21Intermediate scenes :-
-22Scene 0 = []
-23Scene 1 = []
-24Scene 2 = [right(1, 0)]
-25Scene 3 = [right(1, 0), right(2, 1)]
-26Scene 4 = [right(1, 0), right(2, 1), right(3, 2)]
-27Scene 5 = [right(1, 0), right(2, 1), right(3, 2), right(4, 3)]
-28Python function :-
-29‘‘‘python
-30def placing_all_right(objects):
-31 for i in range(len(objects)):
-32 keep_at_head(objects) # select one object from the objects set and keep the head at this location
-33 move_head(dir = ’right’) # move the head to the right of the previous position
-34‘‘‘
-35# The current task for which program needs to be found
-36Current task:- Construct a tower of size 6.
-37Final state :- [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
-38Intermediate scenes :-
-39Scene 0 = []
-40Scene 1 = []
-41Scene 2 = [top(1, 0)]
-42Scene 3 = [top(1, 0), top(2, 1)]
-43Scene 4 = [top(1, 0), top(2, 1), top(3, 2)]
-44Scene 5 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3)]
-45Scene 6 = [top(1, 0), top(2, 1), top(3, 2), top(4, 3), top(5, 4)]
-46Python function :-
-47‘‘‘python
-48def tower(size, objects):
-49??
-50‘‘‘
-Listing 9: Prompting Strategy for LLM baselines (GPT-4)
-VLM/GPT-4-V Unlike LLM, VLMs have the abilities to process the demonstration as a sequence of visual frames.
-Therefore rather than providing the symbolic spatial relations between every scene we instead directly provide all the intermediate scenes for the given demonstration.
-Further we also relax the assumption that there are no distractor objects.
-As shown in figure 13 We first give information about the set of primitive operators and the structures that we have already learnt (library of concepts).
-In order to visually ground the semantics of our primitive actions we give 3 example tasks (natural language instruction and intermediate scenes) that do not directly refer to any structure, along with corresponding sequence of actions taken (# Demonstration for visual grounding).
-We further provide another example (without scenes) demonstrating how to write generalizable Python function for a given task using our operators (# Example for generalization).
-Finally we give the natural language instruction and corresponding scenes for the current task along with signature of the program to be learnt (# Current task description).
-Figure 13: Prompt example for VLM-baseline.
-Figure shows the prompting strategy for GPT-4-V that includes the primitive actions, the example tasks for visual grounding, example of writing generalizable Python functions and the demonstration frames.
-Appendix D Supplementary Results
-Out-of-Distribution Performance:
-Table 8: Out-of-Distribution Performance (mean $\pm$ std-error)
-Model Simple Complex
-IoU MSE IoU MSE
-SPG(Ours) 0.892 $\pm$ 0.065 4.386e-4 $\pm$ 2.387e-4 0.804 $\pm$ 0.025 0.001 $\pm$ 5.391e-4
-GPT-4 0.776 $\pm$ 0.023 0.006 $\pm$ 0.001 0.131 $\pm$ 0.019 0.019 $\pm$ 1.498e-3
-GPT-4V 0.575 $\pm$ 0.026 0.013 $\pm$ 0.001 0.290 $\pm$ 0.016 0.011 $\pm$ 1.316e-3
-SD 0.236 $\pm$ 0.005 0.006 $\pm$ 7.495e-4 0.150 $\pm$ 0.011 0.011 $\pm$ 2.860e-3
-SD+G 0.273 $\pm$ 0.004 0.006 $\pm$ 6.180e-4 0.154 $\pm$ 0.010 0.014 $\pm$ 2.958e-3
-Relative decrease in Out-of-Distribution performance:
-Table 9: Relative Decrease in IoU
-Model Simple (%) Complex (%)
-SPG(Ours) 7.27 5.74
-GPT-4 12.61 53.87
-GPT-4V 23.33 41.64
-SD 51.84 67.67
-SD+G 63.25 74.72
-Performance Dataset II (i.e. name reversed evaluation): Table  10 gives the corresponding program accuracies, while Table  11 give the corresponding IoU/MSE metrics along with standard errors.
-Table 10: PA (Reversed Names)
-Model Simple Complex
-SPG(Ours) 0.88 0.78
-GPT-4 0.67 0.00
-GPT-4V 0.23 0.00
-Table 11: Performance on Dataset II(Names Reversed)
-Model Simple Complex
-IoU MSE (1e-3) IoU MSE (1e-3)
-SPG(Ours) 0.86 $\pm$ 0.03 1.74 $\pm$ 0.45 0.78 $\pm$ 0.02 3.93 $\pm$ 1.09
-GPT-4 0.78 $\pm$ 0.03 3.16 $\pm$ 0.51 0.00 $\pm$ 0.00 22.73 $\pm$ 1.48
-GPT-4V 0.71 $\pm$ 0.01 3.92 $\pm$ 0.45 0.09 $\pm$ 0.02 21.29 $\pm$ 1.59
-D.1 Qualitative Comparison between Purely-neural (Struct-Diff+Grounder) vs. Ours(SPG)
-Figure  14 gives compares the qualitative results for our approach against Struct-Diffusion with grounder on both in-distribution, Dataset I and out-of-distribution (larger size), Dataset III.
-In in-distribution setting the our method performs slightly better in terms of structure creation for both simple and complex structures, but the difference is not significant.
-However for out-of-distribution setting structures created by our approach are much better than those created by Struct-Diffusion+Grounder.
-Further for this setting structure creation by Struct-Diffusion seems to be much worse for complex structures than simple ones.
-Figure 14: Structure creation comparison between SPG(Ours), and Struct-Diff+Grounder
-D.2 Continual Learning of Neural Concepts
-Given demonstration for the task “Construct a tower of height 4 using chocolate cubes”, we would like to learn the neural embedding for the unknown color chocolate (where we assume that tower has already been learnt and stored in the library $\mathcal{L}$).
-First the instruction is converted into corresponding plan sketch $\mathcal{H}_{S}$ = tower(4, Filter(chocolate, cubes)), which is passed to the visual grounder.
-The grounder detects the presence of an unknown attribute chocolate as an argument to filter, and randomly initializes a new neural embedding for it.
-Using this new embedding along with the already present embedding of cube and the visual features found through ResNet-34, the quasi-symbolic executor outputs a grounded task-sketch.
-The executor executes the grounded task-sketch by getting the semantics of the underlying function i.e.
-tower from the library $\mathcal{L}$.
-MSE+IoU loss computed over the final scene obtained and the expected final scene is backpropogated through the network.
-Note that during backpropogation all the neural modules (action semantics, visual attributes, ResNet-34) are frozen, except for the newly initialized embedding for chocolate.
-For the purpose of differentiable sampling during tower construction we use gumbel-softmax [gumbel] with masking.
-Figure  15 illustrates our approach.
-Figure 15: Continual learning of visual primitives
-D.3 Details for Inference on Novel Tasks using an LLM
-Below we show the [code-as-policies] inspired prompting methodology that we use to get the executable code corresponding to a language specified manipulation task.
-We initially begin by importing the helper functions, spatial direction, primitive functions, and learnt inductive concepts/structures (line 5-11).
-Then we give few examples for how to use and compose the various functions for different tasks (line 16-83).
-Finally we give the instruction of current task, and expect GPT-4 to output the corresponding executable code.
-1# Given a task you have to provide Python code for executing the task
-2
-3# importing available functions
-4
-5from spatial_directions import top, front, back, left, right
-6
-7from primitives import assign_head, move_head, keep_at_head
-8# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
-9
-10from helpers import find_size, filter
-11from structures import row, column, tower
-12
-13
-14#function signature of the imported functions
-15# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
-16filter(color, shape)
-17
-18# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
-19find_size(struct_name = str_name, objects = ObjSet)
-20
-21# assigns the head to the location of the object
-22assign_head(at_obj_loc)
-23
-24# moves the head in the given dir
-25move_head(dir)
-26
-27# keeps the object obj at the head
-28keep_at_head(obj)
-29
-30
-31#Examples:
-32#Instruction: Move the green block to the left of the red dice
-33assign_head(at_obj_loc = filter(red, dice))
-34move_head(left)
-35keep_at_head(obj = filter(green, cube))
-36
-37# Instruction: Find the size of the tower made of yellow legos
-38find_size(struct_name = tower, objects = filter(yellow, lego))
-39
-40#Instruction: Find the size of the row made of orange cubes
-41find_size(struct_name = row, objects = filter(orange, cube))
-42
-43# Instruction: Find the size of the column made of cyan cubes
-44find_size(struct_name = column, objects = filter(cyan, cube))
-45
-46# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
-47assign_head(at_obj_loc = filter(red, dice))
-48move_head(left)
-49keep_at_head(obj = filter(green, cube))
-50assign_head(at_obj_loc = filter(green, cube))
-51move_head(top)
-52keep_at_head(obj = filter(yellow, cube))
-53
-54# Instruction: Construct a row of green legos of length 3 to the right of the blue block
-55assign_head(at_obj_loc = filter(blue, block))
-56move_head(right)
-57row(length = 3, objects = filter(green, legos))
-58
-59
-60# Instruction: Construct a tower of size 3 using red cubes
-61tower(height = 3, objects = filter(red, cube))
-62
-63# Instruction: Construct a row of size 5 using blue legos
-64row(length = 5, objects = filter(blue, lego))
-65
-66# Instruction: Construct a column of size 6 using green die
-67column(length = 6, objects = filter(green, dice))
-68
-69# Instruction: Place 3 green blocks so that one block is to the right of the other
-70green_blocks = filter(green, block)
-71keep_at_head(green_blocks)
-72move_head(right)
-73keep_at_head(green_blocks)
-74move_head(right)
-75keep_at_head(green_blocks)
-76
-77# Instruction: Place 3 red legos on top of one another
-78red_legos = filter(red, lego)
-79keep_at_head(red_legos)
-80move_head(top)
-81keep_at_head(red_legos)
-82move_head(top)
-83keep_at_head(red_legos)
-84
-85
-86# CURRENT TASK
-87# Instruction: Construct tower of white cubes to the same height as
-88# existing tower of green die
-89
-90# GPT-4s output
-91#First, we have to find the height of the tower of green dice,
-92#then construct a tower of white cubes of the same size
-93
-94tower_size = find_size(struct_name = tower, objects = filter(green, die))
-95tower(height = tower_size , objects = filter(white, cube))
-Listing 10: Prompting method for the task of constructing tower of white cubes to the same height as existing tower of green die
-1# Given a task you have to provide Python code for executing the task
-2
-3# importing available functions
-4from spatial_directions import top, front, back, left, right
-5from primitives import assign_head, move_head, keep_at_head
-6# HEAD is a imaginary pointer keeping track of the current spatial location in consideration
-7from helpers import find_size, filter
-8from structures import row, column, tower
-9#function signature of the imported functions
-10# finds all the objects with the given color and shape, returns a mask denoting the probability of object selection
-11filter(color, shape)
-12# finds the size of the structure struct_name that is formed with objects of the given type, returns the size of the structure (whose type is integer), arguments for this should be provided as kwargs
-13find_size(struct_name = str_name, objects = ObjSet)
-14# assigns the head to the location of the object
-15assign_head(at_obj_loc)
-16# moves the head in the given dir
-17move_head(dir)
-18# keeps the object obj at the head
-19keep_at_head(obj)
-20
-21#Examples:
-22#Instruction: Move the green block to the left of the red dice
-23assign_head(at_obj_loc = filter(red, dice))
-24move_head(left)
-25keep_at_head(obj = filter(green, cube))
-26
-27# Instruction: Find the size of the tower made of yellow legos
-28find_size(struct_name = tower, objects = filter(yellow, lego))
-29
-30#Instruction: Find the size of the row made of orange cubes
-31find_size(struct_name = row, objects = filter(orange, cube))
-32
-33# Instruction: Find the size of the column made of cyan cubes
-34find_size(struct_name = column, objects = filter(cyan, cube))
-35
-36# Instruction: Move the green block to the left of the red dice and the yellow block to the top of the green block
-37assign_head(at_obj_loc = filter(red, dice))
-38move_head(left)
-39keep_at_head(obj = filter(green, cube))
-40assign_head(at_obj_loc = filter(green, cube))
-41move_head(top)
-42keep_at_head(obj = filter(yellow, cube))
-43
-44# Instruction: Construct a row of green legos of length 3 to the right of the blue block
-45assign_head(at_obj_loc = filter(blue, block))
-46move_head(right)
-47row(length = 3, objects = filter(green, legos))
-48
-49
-50# Instruction: Construct a tower of size 3 using red cubes
-51tower(height = 3, objects = filter(red, cube))
-52
-53# Instruction: Construct a row of size 5 using blue legos
-54row(length = 5, objects = filter(blue, lego))
-55
-56# Instruction: Construct a column of size 6 using green die
-57column(length = 6, objects = filter(green, dice))
-58
-59# Instruction: Place 3 green blocks so that one block is to the right of the other
-60green_blocks = filter(green, block)
-61keep_at_head(green_blocks)
-62move_head(right)
-63keep_at_head(green_blocks)
-64move_head(right)
-65keep_at_head(green_blocks)
-66
-67# Instruction: Place 3 red legos on top of one another
-68red_legos = filter(red, lego)
-69keep_at_head(red_legos)
-70move_head(top)
-71keep_at_head(red_legos)
-72move_head(top)
-73keep_at_head(red_legos)
-74
-75
-76# CURRENT TASK
-77# Instruction: Construct a tower of height 6 using red and blue blocks that are alternating
-78
-79# GPT-4s output
-80# Python code:
-81
-82# Define the red and blue blocks
-83red_blocks = filter(’red’, ’block’)
-84blue_blocks = filter(’blue’, ’block’)
-85
-86# Start at the bottom and alternate building the tower
-87for i in range(6):
-88 if i%2 == 0: # if the stack position is even
-89 keep_at_head(red_blocks)
-90 else: # if the stack position is odd
-91 keep_at_head(blue_blocks)
-92 if i != 5: # if not at the top of the tower
-93 move_head(top)
-Listing 11: Prompting method for the task of constructing tower of alternating red and blue cubes
-To find the size of a given structure in the given scene we define the function find_size, which takes the name of structure, all the objects in the initial scene, mask of the objects (a distribution over the objects based on the attributes), and the initial state.
-(we assume that this function has access to the semantics of all the concepts learnt so far, through a transition function).
-Algorithm  1 gives the pseudocode for the function find_structure.
-Below we provide a brief explanation for it.
-  1. 1.
-    First we assign our head to every block in the available blocks (line 6)
-  2. 2.
-    Then we begin constructing/visualizing the corresponding structure from that block beginning with a size of 1. (line 7)
-  3. 3.
-    For each structure created/visualized we compare the blocks moved for the structure creation with corresponding blocks originally present in the scene, and perform a matching between these blocks and a subset of the blocks originally present (line 11-26).
-  4. 4.
-    If we are able to find a mapping for each moved block, such that each mapped pair has an IoU greater than a threshold, we increase the next potential size to test by 1 (line 26-27).
-  5. 5.
-    The final size is the size corresponding to 2nd last iteration, before termination (line 28).
-  6. 6.
-    We return the maximum of all the possible structures that are found (line 34)
-Algorithm 1 Find Size
-1:$name$: structure name, $objs$: object list, $state$: initial state, $mask$: object mask
-2:Size of the maximum sized structure found
-3:$found\leftarrow[]$
-4:for each $cand$ in $objs$ do
-5:     $size\leftarrow 1$
-6:     $curr\leftarrow state.copy()$
-7:     while true do
-8:         $new\_state\leftarrow assign\_head(cand)$
-9:         $vis\_state,num\_mov,rew\leftarrow transition($
-10:          $new\_state,name,[(^{\prime}size^{\prime},size),(^{\prime}objects^{\prime},mask)])$
-11:         $topk\leftarrow torch.topk(mask,num\_mov)$
-12:         $possible\leftarrow true$
-13:         $matches\leftarrow[]$
-14:         for each $idx$ in $topk$ do
-15:              $possible\leftarrow$ (idx   in  objs)
-16:              $match\_ok\leftarrow false$
-17:              for $m\_idx,obj$ in enumerate(state.state) do
-18:                  $iou\leftarrow iou2d(vis\_state.state[idx],$
-19:                   $state.state[m\_idx])$
-20:                  $match\_ok\leftarrow(iou>0.75)and$
-21:                   $(m\_idx\,in\,objs)$
-22:                  if $match\_ok$ then
-23:                       $matches.append(m\_idx)$
-24:                       break
-25:                  end if
-26:              end for
-27:              if not $match\_ok$ then
-28:                  break
-29:              end if
-30:         end for
-31:         if $possible$ then
-32:              $size\leftarrow size+1$
-33:         else
-34:              $found.append((size-1,matches))$
-35:              break
-36:         end if
-37:     end while
-38:end for
-39:return max(size for size, match in $found$)
-D.4 Details on MCTS Variants for Plan Search
-Here we provide the details for 3 different plan search methods, that search over the space $\mathcal{A}=\mathcal{A}_{c}\cup\mathcal{A}_{p}$
-  • MCTS+L+P: This is the approach that we describe in section 4.2.
-For every concept say Tower $\in\mathcal{L}$ we have a corresponding set of macro action say Make_Tower(3, objects) (L).
-Further we use a neural pruner $\pi_{\text{neural}}$ that outputs a primitive action $a_{p}^{*}$ (given current state and next expected state).
-We only consider the actions $\mathcal{A}_{c}\cup\{a_{p}^{*}\}$ during our search from the given state.
-This helps to reduce the effective branching factor and allows to search longer length plans within the same computational budget.
-  • MCTS+L-P: Here we do not use the reactive policy, therefore the branching factor for every node becomes $\mathcal{A}$ = $\mathcal{A}_{c}\cup\mathcal{A}_{p}$.
-  • MCTS-L+P: We only search among the space of primitive actions i.e $\mathcal{A}_{p}$.
-Given a state $\tilde{s}_{t}$ and corresponding next expected state $s_{t+1}$ we greedily pick the action $a_{p}^{*}=\pi_{\text{neural}}(\tilde{s}_{t},s_{t+1})$.
-This method is much more faster than the previous 2 methods as there is no explicit search.
-However the corresponding policy is trained only to output an action $a\in\mathcal{A}_{p}$ and lacks the ability to output modular plans composed of macro actions such as Make_Tower(3, objects) $\in\mathcal{A}_{c}$ (the action space $\mathcal{A}_{c}$ is increasing with time and the architecture of network needs to be changed accordingly).
-As a result the plans found are not modular and difficult to generalize.
-Further, the reactive policy is not trained to output reset_head(), store_head() as additional annotated data is required in order to train a classifier over them.
-This further decreases the space of grounded plans (and therefore corresponding generic programs) such a policy can represent.
-Training a reactive policy that can handle actions such as reset_head() and an action space $\mathcal{A}_{c}$ that grows with time is part of future work.
-D.5 Goal-conditioned Planning with Learnt Concepts
-Why is it difficult to hand encode a PDDL for our domain?
-Most of the PDDL description of blocks world assume actions involving only the spatial relation onTop, which limits their applicability to describing different structures like row that need spatial relations like onRight.
-Further a single action might lead to varied effects/post-conditions based on the initial state.
-16 gives 2 example of the same action moveOnTop(A, B) which would end up giving adding different number of spatial relations.
-Figure 16: Difficult to encode post-conditions. Illustration of a domain where encoding a PDDL for direct planning is challenging.
-Approach Overview.
-Given an instruction $\Lambda$ = “Construct a staircase of magenta die having 3 steps”, we first convert it into corresponding grounded task sketch $H_{S}^{*}$ = staircase(3, [3, 2, 1, $\cdots$]).
-Executing the corresponding program of staircase (by getting the semantics from the library $\mathcal{L}$) on the desired objects we get the expected final scene $S_{f}^{\prime}$ in bounding box space.
-The initial scene $S_{i}$ and expected final scene $S_{f}^{\prime}$ are converted into scene graph $SG_{i}^{\prime}$ and $SG_{f}^{\prime}$ (described in  16).
-The relations between the task relevant objects in $SG_{f}^{\prime}$ act as propositions/relations for goal check and the initial scene graph act as the initial state.
-Then a neuro-symbolic planner is used to obtain the optimal plan from the start state to a state that satisfies the goal.
-Below we also detail different aspects of the approach.
-Scene-graph Extraction.
-2 gives the algorithm used for generating scene graph from a given scene (set of bounding boxes).
-Suppose we need to check whether there exists a relation of the form $(i,j,direction)$ i.e.
-block $i$ is in the direction $direction$ of block $j$, in a given scene.
-We first initialize the head at the position/bounding-box of block $j$ (line 7).
-Then we move the head in direction $direction$ (line 9).
-We claim that the relationship would exists if bounding box for block $i$ has IoU $>$ 0.75 with the predicted_head (line 10-12).
-Algorithm 2 Get Scene Graph
-1:procedure get_scene_graph($bboxes$) // bboxes describe the corresponding scene for which scene graph is to be found
-2:     $spatial\_relations\leftarrow[]$
-[…]
-4:     for $i,bbox$ in $enumerate(bboxes)$ do
-5:         for $j,other\_bbox$ in $enumerate(bboxes)$ do
-6:              $final\_head\leftarrow bboxes[i]$
-7:              $init\_head\leftarrow bboxes[j]$
-8:              for $direction$ in $directions$ do
-9:                  $pred\_head\leftarrow\text{move\_head}(direction,init\_head)$
-10:                  $iou\_score\leftarrow IoU(final\_head,pred\_head)$
-11:                  if $iou\_score>0.75$ then
-12:                       $spatial\_relations.append((i,j,direction))$
-13:                  end if
-14:              end for
-15:         end for
-16:     end for
-17:     return $spatial\_relations$
-18:end procedure
-Pre-conditions. We define the following 2 preconditions (and learn their grounding) in order to ensure that the generated plans are physically possible.
-  1. 1.
-    is-clear(blk, dir): We need to check whether a block blk has some free space in direction dir.
-For this we simply move the head in the direction dir with respect to the block blk.
-If the resulting position of head has 0 overlap with bounding boxes of all the other objects the predicate is True otherwise False.
-  2. 2.
-    will-not-be-floating(pred_loc): We need to check whether the resultant/predicted location of an object on taking an action would be dynamically stable or not.
-The location would be stable if either it is on top of some already placed object or it is on the table surface.
-The former can be checked through the resultant scene graph itself (that is obtained by applying the algorithm  2), while for the later we train an on-table classifier.
-This would take as input a bounding box and predict whether the box is on the table or not.
-For training this classifier we use the dataset of pretraining phase.
-The blockwise positive and negative sample annotation can be done automatically by giving GPT-4 the corresponding scene graph and then querying which objects are on the table and which aren’t.
-17 gives an example.
-(Though we have not taken this approach for the complete dataset of 5k samples due to high cost).
-Figure 17: Method to get annotation for training on-table classifier.
-Actions. We define the following two types actions
-  1. 1.
-    place-random(blk): To place a block at a random free position on the table.
-For this we train a generative model (VAE) which learns the underlying distribution of bounding boxes for all the blocks that lie on the table.
-Given a scene we would sample position (bounding box) from this until we get a position that is not overlapping with the existing blocks in the scene.
-For training the VAE we assume that for every demonstration in the pretraining data, the first scene has all the objects randomly placed on the table (we could have also used the positive examples used for training on-table classifier).
-  2. 2.
-    move(rel, blk1, blk2): This action corresponds to moving the blk1 in the direction rel of blk2, resulting in the addition of a relation (rel, blk1, blk2) in the set of spatial relations.
-This action is defined as a sequential composition of the actions assign_head(blk2), move_head(rel), keep_at_head(blk1) (blk1 is a one hot tensor for the corresponding block).
-Techniques and heuristic for efficient planning.
-Since the action space for the planner could be $o(n^{2})$, where n is the number of objects we adopt the following techniques to make planning scalable/efficient:
-  1. 1.
-    Heuristic - We define the heuristic value h(s) for a state s, as the number of relations that are present in the goal but are absent in the scene graph corresponding to the state s.
-Even though this heuristic is not admissible (as it may over overestimate the cost to goal), it was found to work optimally in most of the cases.
-  2. 2.
-    Greedy-pruning - We assume that all the actions resulting in states with higher or same heuristic value would be of the form place-random(blk).
-This means among the actions of the form move(rel, blk1, blk2) we only select those that lead to states with decreased heuristic value.
-  3. 3.
-    Relevant-object-set - Suppose O is the set of objects that are part of atleast one of the predicate in goal.
-We define O’ as the transitive closure of O with respect to the relation $Related$ in the initial state $s_{i}$, where $SG(s_{i})$ is scene graph for the initial state
-    $Related(a,b,s_{i})\iff\exists dir((dir,a,b)\in SG(s_{i})\vee(dir,b,a)\in SG(s_{i}))\;$ (9)
-    We assume O’ is the relevant set of object for completing the task and actions that move any other object should not be taken.
-Appendix E Broader Impact
-This work creates foundational knowledge in understanding human-like spatial abstractions.
-This work contributes towards the development of explainable and interpret-able learning architectures that may eventually contribute towards the development of embodied agents collaborating with and assisting humans in performing tasks.
-No negative impact of this work is envisioned.
-Appendix F Details regardign Compute Requirements
+Listing 15: Definition of inductive concepts
+Appendix G Computational Requirements: Details
 All our experiments were run on a server with the following machine specifications.
@@
 Time taken for our approach during inference is less than 2 minutes per dataset.
-Appendix G Hyperparameters and architecture details
-G.1 Architecture for neural modules
-Action Simulator:
-1import torch.nn as nn
-2
-3class ActionSimulatorNetwork(nn.Module):
-4 def __init__(self, bbox_mode, hidden_size = 256):
-5 super(ActionSimulatorNetwork, self).__init__()
-6 self.bbox_mode = bbox_mode
-7 self.hidden_size = hidden_size
-8
-9 self.action_semantics_encoder = nn.Sequential(
-10 nn.Linear(5, hidden_size),
-11 nn.ReLU(),
-12 nn.Linear(hidden_size, hidden_size),
-13 nn.ReLU()
-14 )
-15 self.argument_encoder = nn.Sequential(
-16 nn.Linear(5, hidden_size),
-17 nn.ReLU(),
-18 nn.Linear(hidden_size, hidden_size),
-19 nn.ReLU()
-20 )
-21 self.decoder = nn.Sequential(
-22 nn.Linear(hidden_size, hidden_size),
-23 nn.ReLU(),
-24 nn.Linear(hidden_size, 5),
-25 nn.Tanh()
-26 )
-Listing 12: Action Simulator Network in PyTorch
-Reactive Policy($\pi_{\text{neural}}$):
-1import torch.nn as nn
-2
-3class NeuralSearch(nn.Module):
-4 def __init__(self, action_space=6):
-5 super(NeuralSearch, self).__init__()
-6 self.action_space = action_space
-7 self.fc1 = nn.Linear(10, 256)
-8 # self.bn1 = nn.BatchNorm1d(256)
-9 self.bn1 = nn.Identity()
-10 self.fc2 = nn.Linear(256, 256)
-11 # self.bn2 = nn.BatchNorm1d(256)
-12 self.bn2 = nn.Identity()
-13 self.fc3 = nn.Linear(256, 256)
-14 # self.bn3 = nn.BatchNorm1d(256)
-15 self.bn3 = nn.Identity()
-16 self.fc4 = nn.Linear(256, action_space)
-Listing 13: Neural Search in PyTorch
-Random Position predictor (for grounding of place-random(blk)):
-1import torch.nn as nn
-2
-3class VAE(nn.Module):
-4 def __init__(self, input_dim, latent_dim):
-5 super(VAE, self).__init__()
-6 self.input_dim = input_dim
-7 self.latent_dim = latent_dim
-8 # Encoder
-9 self.fc1 = nn.Linear(input_dim, 512)
-10 self.bn1 = nn.BatchNorm1d(512)
-11 self.fc2 = nn.Linear(512, 512)
-12 self.bn2 = nn.BatchNorm1d(512)
-13 self.fc3 = nn.Linear(512, 512)
-14 self.bn3 = nn.BatchNorm1d(512)
-15 self.fc4 = nn.Linear(512, 512)
-16 self.bn4 = nn.BatchNorm1d(512)
-17 self.fc51 = nn.Linear(512, latent_dim) # Mean of the latent space
-18 self.fc52 = nn.Linear(512, latent_dim) # Log-variance of the latent space (log-var for numerical stability)
-19
-20 # Decoder
-21 self.fc5 = nn.Linear(latent_dim, 512)
-22 self.bn5 = nn.BatchNorm1d(512)
-23 self.fc6 = nn.Linear(512, 512)
-24 self.bn6 = nn.BatchNorm1d(512)
-25 self.fc7 = nn.Linear(512, 512)
-26 self.bn7 = nn.BatchNorm1d(512)
-27 self.fc8 = nn.Linear(512, 512)
-28 self.bn8 = nn.BatchNorm1d(512)
-29 self.fc9 = nn.Linear(512, input_dim)
-Listing 14: VAE in PyTorch
-On-table classifier (for grounding of will-not-be-floating(pred_loc):
-1import torch.nn as nn
-2
-3class TableClassifier(nn.Module):
-4 def __init__(self):
-5 super(TableClassifier, self).__init__()
-6 self.fc1 = nn.Linear(5, 16)
-7 self.bn1 = nn.BatchNorm1d(16)
-8 self.fc2 = nn.Linear(16, 16)
-9 self.bn2 = nn.BatchNorm1d(16)
-10 self.fc3 = nn.Linear(16, 16)
-11 self.bn3 = nn.BatchNorm1d(16)
-12 self.fc4 = nn.Linear(16, 1)
-13 self.bn4 = nn.BatchNorm1d(1)
-14 self.sigmoid = nn.Sigmoid()
-Listing 15: Table Classifier in PyTorch
-G.2 Hyperparameters used in experiment
-As indicated in  A.4 for the purpose of generalization through multiple candidate plans (from 1 demonstration) we chose the top-k plans (as measured by overall IoU achieved).
-The k chosen for all our experiments involving MCTS was 5.
-(The performance of our best approach was found to be the same for k=5 to 20).
-For every plan we obtain 3 programs from GPT-4 by re-prompting it 3 times with the same input prompt (with temperature > 0).
-From the pool of these 3*k programs we chose the one with highest IoU reward by running each of them on the given demonstration.
-The discount factor kept for our search is $\gamma$ = 0.95, and unless explicitly specified the number of expansions steps used = 5000.